### v1

# MSACL: Multi-Step Actor-Critic Learning with Lyapunov Certificates for Exponentially Stabilizing Control

Abstract
Achieving provable stability guarantees in model-free reinforcement learning (RL) for control systems remains a fundamental challenge, particularly in balancing exploration efficiency with rigorous safety requirements. This article introduces MSACL, a novel framework that seamlessly integrates exponential stability theory with maximum entropy reinforcement learning through multi-step Lyapunov certificate learning. Unlike conventional approaches that rely on complex reward engineering or single-step constraints, MSACL leverages off-policy multi-step trajectory data to learn Lyapunov certificates that strictly satisfy theoretical stability conditions. By introducing the concept of Exponential Stability Labels (ESL) and a $\lambda$-weighted aggregation mechanism, the framework effectively balances the bias-variance trade-off inherent in multi-step learning. Policy optimization is subsequently guided by a stability-aware advantage function, which ensures that the learned policy $\pi$ promotes the rapid descent of the Lyapunov certificate. We rigorously evaluate the proposed method across six benchmark environments, encompassing four stabilization tasks and two high-dimensional nonlinear tracking scenarios. The experimental results demonstrate consistent superiority over baseline RL algorithms and state-of-the-art Lyapunov-based RL algorithms. Remarkably, MSACL achieves both exponential stability and rapid convergence under simple, intuitive reward structures, while exhibiting significant robustness to environmental uncertainties and exceptional generalization to unseen reference trajectories. Our sensitivity analysis reveals that the multi-step horizon $n$ is a critical factor in stability learning, with $n=20$ established as a robust default across diverse dynamical systems. By establishing a rigorous connection between Lyapunov stability theory and off-policy actor-critic frameworks, MSACL provides a foundational approach for developing verifiably safe learning-based controllers in safety-critical applications. To facilitate reproducibility and further research, the source code and benchmark environments of MSACL will be made publicly available soon.
I Introduction
Reinforcement learning (RL) has emerged as a powerful paradigm for controlling complex, nonlinear dynamical systems without requiring analytical models [1]. Despite its empirical success in high-dimensional tasks [2, 3], the widespread deployment of RL in safety-critical applications, including autonomous driving and human-robot interaction, remains constrained by the lack of formal reliability guarantees [4, 5]. In these domains, stability is not merely a desirable metric but an imperative constraint; failure to maintain the system within a safe operating region can lead to catastrophic consequences. While approaches like constrained Markov Decision Processes (cMDPs) [6, 7] and Hamilton-Jacobi reachability analysis [8, 9] offer partial solutions, certificate-based methods have established themselves as the gold standard for verifying learning agents. By leveraging rigorous control-theoretic tools, specifically Lyapunov functions for asymptotic stability [10, 11] and control barrier functions (CBFs) for set invariance [12, 13], these methods provide concise, data-driven proofs that guarantee the system will converge to an equilibrium or remain strictly within safety boundaries, offering a level of assurance that standard reward-driven learning cannot match.
The synthesis of stability and safety certificates has undergone a paradigm shift from computationally expensive analytical methods, such as sum-of-squares (SOS) optimization with its inherent restriction to polynomial dynamics [14], to scalable neural approaches. Around 2018, foundational work on neural certificates emerged, marking a turning point in bridging black-box learning with verifiable control. Early efforts by Berkenkamp et al. [15] employed Gaussian Processes (GP) to learn Lyapunov functions for safe exploration, while Richards et al. [16] and Chow et al. [17] developed theoretical frameworks for synthesizing Lyapunov functions to estimate regions of attraction. These studies established that Lyapunov certificates as well as barrier-based counterparts provide a rigorous mathematical foundation for analyzing the stability and safety of nonlinear systems. Building on this verification-centric perspective, recent research has advanced toward the joint synthesis of policies and certificates: Chang et al. [18], for instance, showed that a neural certificate can act as a “critic” or supervision signal during training, directly guiding policy learning to satisfy formal stability conditions. This evolution underscores the dual role of learned certificates, which not only enable post-hoc safety verification but also actively shape the learning process to ensure stability and safety by design, offering guarantees that standard trial-and-error reinforcement learning cannot provide.
To enable effective learning in complex continuous spaces, we build our approach upon the framework of Maximum Entropy Reinforcement Learning (MERL). Algorithms such as Soft Actor-Critic (SAC) [19] and its distributional variants like DSAC [20, 21] maximize not just the expected return but also the entropy of the policy. This entropy regularization encourages robust exploration and prevents the policy from converging prematurely to suboptimal deterministic behaviors. Given its superior sample efficiency and robustness to disturbances, the maximum entropy framework serves as an ideal substrate for developing advanced, stability-aware control algorithms.
However, standard model-free RL algorithms, including DDPG [22], TRPO [23], PPO [24], TD3 [25], D4PG [26], SAC, and DSAC, face significant limitations when applied to stabilization or tracking tasks. First, they lack intrinsic stability guarantees; since policies trained via these methods optimize solely for scalar reward maximization, stability is essentially an incidental byproduct that is achieved only when the agent implicitly discovers stable behaviors during the exploration process, rather than being structurally enforced. Second, the efficacy of these algorithms relies heavily on exhaustive reward engineering, requiring practitioners to laboriously calibrate coefficients for state deviation, control effort, and auxiliary terms to elicit stable behavior. Our proposed framework, MSACL (Multi-Step Actor-Critic Learning with Lyapunov Certificates), directly addresses these shortcomings. By adopting a formulation rooted in quadratic performance functionals such as LQR-like costs, MSACL eliminates the need for heuristic reward shaping. Most importantly, it guarantees stability by design, ensuring that the synthesized controller is not only high-performing but also mathematically certifiable.
While Lyapunov-based RL algorithms have been proposed to mitigate the two issues above, existing methods exhibit distinct shortcomings. The Policy Optimization with Lyapunov Certificates (POLYC) [27] algorithm offers theoretical guarantees but relies on on-policy learning, which results in prohibitive sample complexity for physical systems. Conversely, the Lyapunov-based Actor-Critic (LAC) [28] algorithm introduces an off-policy approach but fundamentally misaligns the objective: it minimizes a stability risk rather than a comprehensive control cost. This often leads to aggressive controllers that stabilize the system at the cost of excessive control effort, rendering them impractical for real-world actuators. Moreover, POLYC and LAC typically certify only asymptotic stability, a relatively weaker stability notion that ensures convergence but lacks the guaranteed exponential rate of decay. In practice, this guarantees eventual convergence but fails to impose a minimum convergence rate, often resulting in sluggish transient responses unacceptable for high-performance tasks. To overcome these limitations, MSACL is built upon an off-policy framework [29, 30], significantly improving data efficiency by leveraging historical agent-environment interaction data. Unlike LAC, MSACL optimizes a comprehensive objective covering both state regulation and control effort while enforcing exponential stability, which guarantees faster convergence rates and rigorous stability margins.
Another critical innovation of our approach is the utilization of multi-step data. Existing RL and Lyapunov-based methods typically rely on single-step transitions to enforce learning constraints. This pointwise reliance is susceptible to significant discretization errors when approximating the continuous-time Lie derivative of the Lyapunov function, particularly in the presence of local noise. To mitigate this and capture long-horizon dynamics, we introduce a multi-step learning mechanism. By training on $n$-step trajectories, we effectively bridge the gap between temporal difference learning and rigorous stability verification. We integrate this multi-step data with exponential stability theory to form a robust constraint mechanism. This allows the agent to learn a Lyapunov function that certifies exponential convergence over trajectory segments, rather than just instantaneous descent, thereby reconciling rigorous stability requirements with the need for flexible exploration.
In summary, this article presents a comprehensive RL framework for model-free exponentially stabilizing control. The main contributions are summarized as follows:
  1. (i)
    Exponential Stability Guarantee via Multi-Step Learning: We introduce the MSACL framework, which integrates exponential stability conditions into Lyapunov certificate learning using a multi-step data utilization mechanism. By employing Exponential Stability Labels (ESL), our approach enforces that the system state converges to the equilibrium exponentially, providing stronger performance guarantees than conventional asymptotic convergence methods.
  2. (ii)
    Sample-Efficient Off-Policy Framework: Developed within the maximum entropy RL paradigm, our framework enables efficient off-policy learning. By incorporating trajectory-wise importance sampling to correct for distribution shifts, MSACL allows for effective reuse of historical interaction data, significantly reducing sample complexity compared to existing on-policy Lyapunov-based RL algorithms.
  3. (iii)
    Comprehensive Benchmarking System: We developed a diverse suite of six nonlinear dynamical environments, spanning four stabilization tasks (e.g., VanderPol) and two high-dimensional tracking tasks (e.g., Quadrotor Trajectory Tracking), to rigorously evaluate and provide a standard for Lyapunov-based RL algorithms.
  4. (iv)
    Robustness and Generalization Verification: Our methodology demonstrates significant resilience to stochastic process noise and parametric uncertainties. Furthermore, in tracking scenarios, the learned neural policies exhibit robust generalization capabilities, maintaining high-precision performance even when navigating geometrically distinct, unseen reference trajectories.
  5. (v)
    Bias-Variance Trade-off Analysis: We provide a systematic analysis of the multi-step horizon parameter $n$ and its impact on stability learning. Through the introduction of $\lambda$-weighted aggregation, we effectively manage the trade-off between estimation bias (associated with small $n$) and sampling variance (associated with large $n$), establishing as a robust heuristic for diverse control applications.
The remainder of this article is organized as follows. Section II formulates the stabilization and tracking problems and establishes the theoretical foundations, including the criteria for exponential stability and the maximum entropy reinforcement learning framework. Section III details the methodology for learning the Lyapunov certificate, specifically describing the multi-step data collection strategy and the loss functions designed to enforce formal stability conditions. Section IV presents the learning process for the neural control policy, deriving the optimization objectives for the soft Q-networks, the policy network, and the temperature coefficient. Section V reports the experimental results, providing a comprehensive comparison with baseline algorithms across six benchmark environments, followed by an evaluation of robustness, generalization capabilities, and parameter sensitivity. Finally, Section VI provides a discussion on the underlying design philosophy and concludes the article.
II Preliminaries
In this section, we establish the theoretical foundations for the proposed MSACL framework. We first define the stabilization and tracking problems for unknown discrete-time systems. Subsequently, the concept of exponential stability is introduced along with a Lyapunov-based certificate lemma, providing the criteria for convergence guarantees. Finally, we present the maximum entropy reinforcement learning framework, bridging Lyapunov-based control with model-free actor-critic optimization.
II-A Problem Statement
In this article, we study the stabilization and tracking problems for general discrete-time deterministic control systems:
$\mathbf{x}_{t+1}=f(\mathbf{x}_{t},\mathbf{u}_{t}),$ (1)
where $\mathbf{x}_{t}\in\mathcal{X}\subseteq\mathbb{R}^{n}$ denotes the state at time $t$, $\mathbf{u}_{t}\in\mathcal{U}\subseteq\mathbb{R}^{m}$ represents the control input, $\mathcal{X}$ and $\mathcal{U}$ denote the sets of admissible states and control inputs, respectively, and $f:\mathcal{X}\times\mathcal{U}\to\mathcal{X}$ is the flow map. We assume that $f$ is unknown in our setting, thus focusing on model-free control. It is worth noting that the discrete-time framework presented in this article can be similarly extended to the continuous-time case; interested readers may refer to [10] and [27] for further details.
In the context of control systems, our objective is to find a state feedback policy $\pi:\mathcal{X}\to\mathcal{U}$ such that the control input $\mathbf{u}_{t}=\pi(\mathbf{x}_{t})$ renders the closed-loop system $\mathbf{x}_{t+1}=f(\mathbf{x}_{t},\pi(\mathbf{x}_{t}))$ to satisfy desired properties (e.g., stability, safety). For clarity, we consider the scenario where the state $\mathbf{x}_{t}$ is fully observable, and the policy $\pi$ is a time-invariant mapping. Specifically, we focus on tasks including
  • Stabilization: The objective is to find a policy $\pi$ such that the state ultimately converges to a target state $\mathbf{x}_{g}\in\mathcal{X}$. Without loss of generality, we assume $\mathbf{x}_{g}$ to be the origin, i.e., $\mathbf{x}_{g}=\mathbf{0}$.
  • Tracking: The goal is to find a policy $\pi$ such that the state $\mathbf{x}_{t}$ eventually converges to a reference signal $\mathbf{x}_{t}^{\text{ref}}$ (which may represent a desired trajectory, velocity, or attitude). We define the tracking error as $\mathbf{e}_{t}=\mathbf{x}_{t}-\mathbf{x}_{t}^{\text{ref}}$, thereby transforming the tracking task into a stabilization task where $\mathbf{e}_{t}$ must converge to zero.
In brief, the core problem addressed in this article is formulated as follows:
Problem 1
For the control system (1), find a policy $\pi$ such that: (i) in stabilization tasks, $\mathbf{x}_{t}\to\mathbf{0}$; (ii) in tracking tasks, $\mathbf{e}_{t}\to\mathbf{0}$. Moreover, the policy should be optimized to achieve the fastest possible convergence rate (e.g., exponential stability).
II-B Exponential Stability and Lyapunov Certificates
Definition 1
For the control system (1), the equilibrium point $\mathbf{x}_{g}=\mathbf{0}$ is said to be exponentially stable if there exist constants $C>0$ and $0<\eta<1$ such that
$\|\mathbf{x}_{t}-\mathbf{x}_{g}\|\leq C\eta^{t}\|\mathbf{x}_{0}-\mathbf{x}_{g}\|,\quad\forall\mathbf{x}_{0}\in\mathcal{X}_{0},$ (2)
where $\mathbf{x}_{0}\in\mathcal{X}_{0}\subseteq\mathcal{X}$ denotes the initial state and $t\in\mathbb{Z}_{\geq 0}$.
To guarantee that a closed-loop system is exponentially stable with respect to $\mathbf{x}_{g}$, we introduce the following lemma:
Lemma 1
For the control system (1), if there exists a continuous function $V:\mathcal{X}\mapsto\mathbb{R}$ and positive constants $\alpha_{1},\alpha_{2}>0$ and $0<\alpha_{3}<1$ satisfying:
$\alpha_{1}\|\mathbf{x}_{t}\|^{2}\leq V(\mathbf{x}_{t})$ $\leq\alpha_{2}\|\mathbf{x}_{t}\|^{2},$ (3)
$V(\mathbf{x}_{t+1})-V(\mathbf{x}_{t})$ $\leq-\alpha_{3}V(\mathbf{x}_{t}),$
for all $t\in\mathbb{Z}_{\geq 0}$, where $\mathbf{x}_{t+1}=f(\mathbf{x}_{t},\mathbf{u}_{t})$, then the origin $\mathbf{x}_{g}=\mathbf{0}$ is exponentially stable. The function $V$ is referred to as an exponentially stabilizing control Lyapunov function (ESCLF) [10, 31].
Proof: From the condition $V(\mathbf{x}_{t+1})\leq(1-\alpha_{3})V(\mathbf{x}_{t})$, we can iteratively derive the multi-step relationship:
$V(\mathbf{x}_{t+1})$ $\leq(1-\alpha_{3})V(\mathbf{x}_{t}),$ (4)
$V(\mathbf{x}_{t+2})$ $\leq(1-\alpha_{3})V(\mathbf{x}_{t+1})\leq(1-\alpha_{3})^{2}V(\mathbf{x}_{t}),$
$\vdots$
$V(\mathbf{x}_{t+n})$ $\leq(1-\alpha_{3})V(\mathbf{x}_{t+n-1})\leq\dots\leq(1-\alpha_{3})^{n}V(\mathbf{x}_{t}).$
By incorporating the lower and upper quadratic bounds $\alpha_{1}\|\mathbf{x}_{t}\|^{2}\leq V(\mathbf{x}_{t})\leq\alpha_{2}\|\mathbf{x}_{t}\|^{2}$, it follows that:
$\|\mathbf{x}_{t+n}\|\leq\sqrt{\frac{\alpha_{2}}{\alpha_{1}}}(1-\alpha_{3})^{n/2}\|\mathbf{x}_{t}\|.$ (5)
Specifically, by setting the initial time to $t=0$ and denoting the current time step as $n=t\in\mathbb{Z}_{\geq 0}$, we obtain:
$\|\mathbf{x}_{t}\|\leq\sqrt{\frac{\alpha_{2}}{\alpha_{1}}}(1-\alpha_{3})^{t/2}\|\mathbf{x}_{0}\|.$
Defining $C=\sqrt{\alpha_{2}/\alpha_{1}}$ and $\eta=\sqrt{1-\alpha_{3}}$, the result satisfies Definition 1, thereby proving that the control system (1) is exponentially stable with respect to $\mathbf{x}_{g}=\mathbf{0}$.
II-C Standard RL and Maximum Entropy RL
In the reinforcement learning framework, the control system (1) is interpreted as a discrete-time deterministic environment, where $\mathcal{X}$ and $\mathcal{U}$ represent continuous state and action spaces, respectively. The application of control inputs to the system (1) corresponds to an agent interacting with this environment. Specifically, at each time step $t$, the agent observes the state $\mathbf{x}_{t}$ and executes an action $\mathbf{u}_{t}$ according to a policy $\pi$. Given a reward function $r:\mathcal{X}\times\mathcal{U}\to\mathbb{R}$, which is typically designed to penalize deviations from the target state $\mathbf{x}_{g}$ or reference $\mathbf{x}_{t}^{\text{ref}}$ in accordance with Problem 1, the environment returns a scalar reward $r_{t}:=r(\mathbf{x}_{t},\mathbf{u}_{t})$ and transitions to the next state $\mathbf{x}_{t+1}$. Consequently, the interaction sequence is denoted by the transition tuple $(\mathbf{x}_{t},\mathbf{u}_{t},r_{t},\mathbf{x}_{t+1})$.
While classical control often employs deterministic laws, this article utilizes stochastic policies to facilitate exploration and robustness within the maximum entropy RL (MERL) framework. A stochastic policy is defined as a mapping $\pi:\mathcal{X}\to\mathcal{P}(\mathcal{U})$, where $\mathbf{u}_{t}\sim\pi(\cdot|\mathbf{x}_{t})$ is sampled from a probability distribution. We denote the state and state-action marginals induced by $\pi$ as $\rho_{\pi}(\mathbf{x}_{t})$ and $\rho_{\pi}(\mathbf{x}_{t},\mathbf{u}_{t})$, respectively. Unlike standard RL, which maximizes the expected cumulative discounted reward, MERL optimizes an entropy-augmented objective:
$J_{\pi}=\underset{(\mathbf{x}_{i\geq t},\mathbf{u}_{i\geq t})\sim\rho_{\pi}}{\mathbb{E}}\left[\sum_{i=t}^{\infty}\gamma^{i-t}\left[r_{i}+\alpha\mathcal{H}(\pi(\cdot|\mathbf{x}_{i}))\right]\right],$ (6)
where $\gamma\in[0,1)$ is the discount factor, $\alpha>0$ is the temperature coefficient governing the trade-off between reward and exploration, and $\mathcal{H}(\pi(\cdot|\mathbf{x}))=\mathbb{E}_{\mathbf{u}\sim\pi(\cdot|\mathbf{x})}[-\log\pi(\mathbf{u}|\mathbf{x})]$ is the policy entropy.
The corresponding soft Q-function, $Q^{\pi}(\mathbf{x}_{t},\mathbf{u}_{t})$, represents the expected soft return, which comprises both rewards and entropy terms, starting from $(\mathbf{x}_{t},\mathbf{u}_{t})$ and following $\pi$ thereafter. The soft Bellman operator $\mathcal{T}^{\pi}$ is defined as:
$\mathcal{T}^{\pi}Q^{\pi}(\mathbf{x}_{t},\mathbf{u}_{t})\triangleq r_{t}+\gamma\underset{\mathbf{u}_{t+1}\sim\pi}{\mathbb{E}}\Big[$ $Q^{\pi}(\mathbf{x}_{t+1},\mathbf{u}_{t+1})$ (7)
$-\alpha\log\pi(\mathbf{u}_{t+1}|\mathbf{x}_{t+1})\Big],$
where $\mathbf{x}_{t+1}=f(\mathbf{x}_{t},\mathbf{u}_{t})$. In the soft policy iteration (SPI) algorithm [19], the policy is optimized by alternating between soft policy evaluation, where $Q^{\pi}$ is updated via iterative application of $\mathcal{T}^{\pi}$, and soft policy improvement, which is equivalent to finding a policy $\pi$ that maximizes the soft Q-function augmented with an entropy term for $\mathbf{x}_{t}$:
$\pi_{\rm new}=\arg\max_{\pi}\underset{\mathbf{x}_{t}\sim\rho_{\pi},\mathbf{u}_{t}\sim\pi}{\mathbb{E}}\left[Q^{\pi_{\rm old}}(\mathbf{x}_{t},\mathbf{u}_{t})-\alpha\log\pi(\mathbf{u}_{t}|\mathbf{x}_{t})\right].$ (8)
Although SPI converges to the optimal maximum entropy policy $\pi^{*}$ in tabular settings, practical control problems involve continuous spaces requiring function approximation. Thus, we parameterize the soft Q-function $Q_{\theta}(\mathbf{x}_{t},\mathbf{u}_{t})$ and the policy $\pi_{\phi}(\mathbf{u}_{t}|\mathbf{x}_{t})$ using neural networks with parameters $\theta$ and $\phi$. We assume $\pi_{\phi}$ follows a diagonal Gaussian distribution with mean $\mu_{\phi}(\mathbf{x}_{t})$ and standard deviation $\sigma_{\phi}(\mathbf{x}_{t})$. Actions are sampled using the reparameterization trick: $\mathbf{u}_{t}=g_{\phi}(\mathbf{x}_{t};\epsilon_{t})$, where $\epsilon_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{m})$, ensuring that the objective remains differentiable with respect to $\phi$.
III Learning Lyapunov Certificate
In this section, we develop the learning mechanism for the Lyapunov certificate within a model-free framework. We first introduce a sliding-window data collection strategy to satisfy multi-step trajectory requirements. Subsequently, we design a comprehensive loss function for the Lyapunov network, incorporating boundedness constraints and a multi-step stability loss. To enhance training efficiency and robustness, we employ trajectory-wise importance sampling for off-policy correction, a labeling mechanism for positive and negative samples, and a $\lambda$-weighted aggregation to balance the bias-variance trade-off.
III-A Data Collection
Our algorithm is developed within the maximum entropy RL framework, where data is acquired through agent-environment interactions with the system (1). Unlike the standard SAC algorithm [19] that samples single-step transitions, we leverage the multi-step Lyapunov condition presented in (4), which necessitates data sequences spanning $n$ consecutive steps. We define a single-step interaction tuple as $\mathbf{d}_{t}=(\mathbf{x}_{t},\mathbf{u}_{t},r_{t},\pi_{\phi}(\mathbf{u}_{t}|\mathbf{x}_{t}),\mathbf{x}_{t+1})$, where $\pi_{\phi}(\mathbf{u}_{t}|\mathbf{x}_{t})$ represents the probability density of sampling action $\mathbf{u}_{t}$ at state $\mathbf{x}_{t}$. The collected multi-step data consists of sequences $\{\mathbf{d}_{t},\mathbf{d}_{t+1},\dots,\mathbf{d}_{t+n-1}\}$, which are stored in the experience replay buffer $\mathcal{D}$.
Data storage commences only after $n$ consecutive steps have been accumulated. Specifically, during the initial n1n-1 steps of any rollout, no data is committed to $\mathcal{D}$. Once the rollout length reaches or exceeds $n$, the collection process begins. As illustrated in Fig. 1, we employ a double-ended queue (deque) to temporarily buffer interaction data. During each agent-environment interaction, new data enters the deque. Once the deque reaches a length of $n$, the complete $n$-step sequence is transferred to $\mathcal{D}$, and the oldest data point is removed to maintain the deque size, enabling a continuous sliding window for data collection.
Figure 1: Schematic of the $n$-step sliding window data collection process. A double-ended queue (deque) of size $n$ is utilized to buffer consecutive interaction transitions $\mathbf{d}_{t}$, ensuring that complete $n$-step trajectories are available for the verification of the multi-step Lyapunov certificate.
III-B Loss Function Design of Lyapunov Network
Similar to the soft Q-network and policy network, we introduce a Lyapunov network parameterized by $\psi$, denoted as $V_{\psi}$, to estimate the Lyapunov certificate. In designing the loss function for the Lyapunov certificate, we strictly adhere to the conditions established in Lemma 1, which encompass the boundedness and multi-step stability requirements.
(1) Boundedness Loss: Based on the first condition in Lemma 1, we construct the boundedness loss function to ensure the Lyapunov certificate is quadratically constrained:
$\mathcal{L}_{\text{Bnd}}(\psi)=\frac{1}{Nn}$ $\sum_{i=1}^{N}\sum_{k=0}^{n-1}\Big[\max\left(\alpha_{1}\|\mathbf{x}_{t+k}^{i}\|^{2}-V_{\psi}(\mathbf{x}_{t+k}^{i}),0\right)$ (9)
$+\max\left(V_{\psi}(\mathbf{x}_{t+k}^{i})-\alpha_{2}\|\mathbf{x}_{t+k}^{i}\|^{2},0\right)\Big],$
where $N$ denotes the number of multi-step trajectories in a batch, $i$ denotes the $i$-th multi-step sample in the batch, and $n$ represents the number of steps within each trajectory. This loss term explicitly constrains $V_{\psi}$ to satisfy the boundedness condition defined in (3).
(2) Multi-step Stability Loss: Consider a sampled multi-step trajectory $\{\mathbf{d}_{t},\dots,\mathbf{d}_{t+n-1}\}$ generated by a previous policy $\pi_{\phi_{\text{old}}}$. Since our goal is to train the Lyapunov certificate $V_{\psi}$ to be compatible with the current policy $\pi_{\phi_{\text{new}}}$ for subsequent policy updates, we must account for the off-policy nature of the data. We introduce trajectory-wise importance sampling to correct the distribution shift. For an $n$-step trajectory, the importance sampling ratio for the segment from $t$ to $t+k$ is defined as ($k=1,2,\dots,n-1$):
$\text{IS}_{\text{clip},k}=\prod_{j=t}^{t+k-1}\min\left(\frac{\pi_{\phi}(\mathbf{u}_{j}|\mathbf{x}_{j})}{\pi_{\phi_{\text{old}}}(\mathbf{u}_{j}|\mathbf{x}_{j})},1\right).$ (10)
Note that we apply a clipping operation to the importance sampling ratios to prevent variance explosion and enhance training stability.
Furthermore, we address the second condition in (3), namely the stability requirement. While (4) represents the ideal theoretical decay, empirical trajectories may not strictly satisfy the exponential stability bound in (5) during early stages of training. To bridge the gap between the stability condition and the stochastic samples, we introduce the concepts of positive and negative samples.
Definition 2
Within an $n$-step sample, for a state $\mathbf{x}_{t+k}$, if it satisfies the exponential decay condition:
$\|\mathbf{x}_{t+k}\|\leq\sqrt{\frac{\alpha_{2}}{\alpha_{1}}}(1-\alpha_{3})^{k/2}\|\mathbf{x}_{t}\|,$
then $\mathbf{d}_{t+k}$ is classified as a positive sample; otherwise, it is classified as a negative sample.
Accordingly, we define the Exponential Stability Label (ESL) as:
$\text{ESL}_{k}=\begin{cases}1,&\text{if }\|\mathbf{x}_{t+k}\|\leq\sqrt{\frac{\alpha_{2}}{\alpha_{1}}}(1-\alpha_{3})^{k/2}\|\mathbf{x}_{t}\|\\ -1,&\text{otherwise}.\end{cases}$
This labeling mechanism allows the Lyapunov network to distinguish between transitions that contribute to system stability and those that violate the certificate requirements.
For trajectories composed of multi-step interaction data, we introduce a weighting factor $\lambda\in(0,1)$, inspired by the (truncated) $\lambda$-return in reinforcement learning [33], to balance the trade-off between bias and variance. Specifically, in the stability condition $V_{\psi}(\mathbf{x}_{t+k})\leq(1-\alpha_{3})^{k}V_{\psi}(\mathbf{x}_{t})$ from (4), the parameter $k$ determines the temporal horizon of the system’s evolution. When $k$ is small, the short trajectory horizon provides a limited view of the system’s long-term dynamics under policy $\pi_{\phi}$. Given that the Lyapunov network is randomly initialized and potentially inaccurate in the early stages of training, relying on short-horizon data can introduce significant estimation bias.
Conversely, a larger $k$ captures more extensive information regarding the trajectory evolution (extended horizon), thereby providing a stronger and more accurate learning signal that reduces bias. However, since each transition is sampled from the environment, every step introduces stochasticity; consequently, longer trajectories accumulate higher cumulative variance. Therefore, a small $k$ leads to high bias but low variance, while a large $k$ results in low bias but high variance. To optimize this trade-off, we employ the weighting factor $\lambda$ to aggregate multi-step information.
By integrating the positive/negative sample labels with the weighting factor $\lambda$, we construct the Lyapunov stability loss to enforce a rigorous correspondence between empirical samples and the stability condition (4):
$\mathcal{L}_{\text{Stab}}(\psi)=\frac{1}{N}\sum_{i=1}^{N}\left[\sum_{k=1}^{n-1}\frac{\lambda^{k-1}\mathcal{L}_{\text{diff},k}^{i}(\psi)}{\sum_{j=1}^{n-1}\lambda^{j-1}}\right],$ (11)
and the step-wise difference loss $\mathcal{L}_{\text{diff},k}^{i}(\psi)$ is defined as:
$\mathcal{L}_{\text{diff},k}^{i}(\psi)=$ $\text{IS}_{\text{clip},k}^{i}\cdot\max\Big(0,$
$\text{ESL}_{k}^{i}\cdot\left(V_{\psi}(\mathbf{x}_{t+k}^{i})-(1-\alpha_{3})^{k}V_{\psi}(\mathbf{x}_{t}^{i})\right)\Big).$
(3) Combined Lyapunov Loss: By integrating the boundedness and multi-step stability losses, the total loss function for optimizing the Lyapunov network is formulated as:
$\mathcal{L}_{\text{Lya}}(\psi)=\omega_{\text{Bnd}}\mathcal{L}_{\text{Bnd}}(\psi)+\omega_{\text{Stab}}\mathcal{L}_{\text{Stab}}(\psi),$ (12)
where $\omega_{\text{Bnd}}$ and $\omega_{\text{Stab}}$ are positive weighting coefficients.
IV Learning Policy with Exponential Stability Constraints
This section presents the optimization framework for the proposed MSACL algorithm. We first establish the soft Q-network updates using multi-step Bellman residuals to enhance data efficiency. We then introduce a novel policy loss function that incorporates a stability advantage derived from the Lyapunov certificate, utilizing a PPO-style clipping mechanism to ensure robust convergence toward exponentially stable behavior. Finally, we describe the automated temperature adjustment and synthesize the complete actor-critic learning procedure.
IV-A Loss Function Design of Soft Q Networks
Based on the soft Bellman operator (7), we define the sampled Bellman residual for the $i$-th transition within a multi-step sequence for the soft Q-network $Q_{\theta}$ as:
$\mathcal{B}_{t}^{i}(\theta)=Q_{\theta}(\mathbf{x}_{t}^{i},\mathbf{u}_{t}^{i})-\Big[$ $r_{t}^{i}+\gamma Q_{\bar{\theta}}(\mathbf{x}_{t+1}^{i},\mathbf{u}_{t+1}^{i})$ (13)
$-\alpha\log\pi_{\phi}(\mathbf{u}_{t+1}^{i}|\mathbf{x}_{t+1}^{i})\Big],$
where $Q_{\bar{\theta}}$ denotes the target network parameterized by $\bar{\theta}$, which is updated via polyak averaging to stabilize the learning process. Utilizing the collected multi-step sequences, the loss function for the soft Q-network is defined as the mean squared Bellman error (MSBE) across the entire trajectory:
$\mathcal{L}_{\text{SoftQ}}(\theta)=\frac{1}{Nn}\sum_{i=1}^{N}\sum_{k=0}^{n-1}(\mathcal{B}_{t+k}^{i}(\theta))^{2}.$ (14)
While standard implementations of maximum entropy RL algorithms, such as SAC [19], typically set the upper limit of the inner summation in (14) to zero (i.e., $n=1$) to utilize only single-step transitions, our formulation leverages the complete $n$-step sequence for parameter updates. Theoretically, minimizing the MSBE requires state-action pairs $(\mathbf{x}_{t},\mathbf{u}_{t})$ to be sampled from a distribution that ensures an unbiased stochastic gradient estimate. Specifically, the objective is given by:
$\mathcal{J}_{\text{SoftQ}}(\theta)=$ $\underset{\mathbf{x}_{t}\sim\rho_{\pi},\mathbf{u}_{t}\sim\pi}{\mathbb{E}}\Big[Q_{\theta}(\mathbf{x}_{t},\mathbf{u}_{t})-\Big(r_{t}+\gamma\underset{\mathbf{x}_{t+1}\sim\rho_{\pi},\mathbf{u}_{t+1}\sim\pi}{\mathbb{E}}$
$\big(Q_{\bar{\theta}}(\mathbf{x}_{t+1},\mathbf{u}_{t+1})-\alpha\log\pi_{\phi}(\mathbf{u}_{t+1}|\mathbf{x}_{t+1})\big)\Big)\Big]^{2}.$
Strictly speaking, the transitions within a multi-step sequence are temporally correlated, and utilizing the full sequence simultaneously deviates from the independent and identically distributed (i.i.d.) sampling requirement commonly assumed in stochastic optimization. However, this approach represents a deliberate trade-off between mathematical rigor and sample efficiency. By incorporating all transitions within each collected sequence into the loss function, we significantly enhance the data utilization rate and the information density of each training batch. Experimental evidence suggests that as long as the sequence length $n$ remains moderate, the bias introduced by these short-term correlations does not undermine the stability of the Q-function learning. In practice, the benefits of improved sample efficiency and accelerated convergence often outweigh the theoretical concerns regarding strict uniform sampling, ultimately contributing to more robust control performance.
Remark 1
In the practical implementation of Algorithm 1, we utilize clipped double-Q learning with two networks, $Q_{\theta_{1}}$ and $Q_{\theta_{2}}$, to mitigate the overestimation bias of soft Q-values, as motivated by the TD3 framework [25]. Furthermore, we incorporate the “delayed policy update” technique to ensure that the policy is updated against a more stable value function, thereby facilitating smoother convergence of the actor-critic process.
IV-B Loss Function Design of Policy Network
This subsection details the optimization process for the policy network, which integrates the maximum entropy objective with a Lyapunov-based stability constraint. In the standard SAC framework [19], policy optimization aims to maximize the expected soft Q-value. For the $i$-th multi-step sequence, the SAC-component of the loss function is expressed as:
$\mathcal{L}_{\pi,\text{SAC}}^{i}(\phi)=\frac{1}{n}\sum_{k=0}^{n-1}\Big[$ $Q_{\theta}(\mathbf{x}_{t+k}^{i},g_{\phi}(\mathbf{x}_{t+k}^{i};\epsilon_{t+k}^{i}))$
$-\alpha\log\pi_{\phi}(g_{\phi}(\mathbf{x}_{t+k}^{i};\epsilon_{t+k}^{i})|\mathbf{x}_{t+k}^{i})\Big].$
To ensure the learned policy satisfies the exponential stability requirements, we utilize the learned Lyapunov certificate $V_{\psi}$ to guide the policy updates. Inspired by the Generalized Advantage Estimation (GAE) in PPO [24], we define a stability advantage function based on $V_{\psi}$, which serves as a metric for convergence performance:
$A_{t,k}=(1-\alpha_{3})^{k}V_{\psi}(\mathbf{x}_{t})-V_{\psi}(\mathbf{x}_{t+k}).$ (15)
A positive stability advantage ($A_{t,k}\geq 0$) indicates that the system trajectory from $\mathbf{x}_{t}$ to $\mathbf{x}_{t+k}$ strictly adheres to the exponential decay condition. Maximizing this advantage encourages the policy to select actions that accelerate the decay of $V_{\psi}$, thereby directly increasing the state convergence rate toward the equilibrium $\mathbf{x}_{g}$. Conversely, $A_{t,k}<0$ signifies a violation of the stability certificate, representing a disadvantage that the policy must learn to avoid. This stability-driven optimization paradigm aligns with recent developments in stable model-free control, such as D-learning [34, 35]. To robustly handle the bias-variance trade-off across varying horizons, we apply the weighting factor $\lambda$ to obtain the aggregated stability advantage:
$A_{t,\lambda}=\sum_{k=1}^{n-1}\frac{\lambda^{k-1}A_{t,k}}{\sum_{j=1}^{n-1}\lambda^{j-1}}.$ (16)
For a multi-step sample $\{\mathbf{d}_{t},\dots,\mathbf{d}_{t+n-1}\}$, we define the policy importance sampling ratio based on the initial transition as $\rho^{i}(\phi)=\frac{\pi_{\phi}(\mathbf{u}_{t}^{i}|\mathbf{x}_{t}^{i})}{\pi_{\phi_{\text{old}}}(\mathbf{u}_{t}^{i}|\mathbf{x}_{t}^{i})}$. Incorporating the clipping mechanism to prevent excessively large policy updates, the final policy loss function is formulated as:
$\mathcal{L}_{\pi}(\phi)=-\frac{1}{N}\sum_{i=1}^{N}\Big[$ $\mathcal{L}_{\pi,\text{SAC}}^{i}(\phi)+\min\Big(\rho^{i}(\phi)A_{t,\lambda}^{i},$ (17)
$\text{clip}(\rho^{i}(\phi),1-\varepsilon,1+\varepsilon)A_{t,\lambda}^{i}\Big)\Big].$
Remark 2
It is important to distinguish between the two importance sampling (IS) strategies employed in this article. In (11), sequence-level IS (from $t$ to t+k1t+k-1) is used to train the Lyapunov certificate $V_{\psi}$. While sequence-level IS typically introduces higher variance, it provides the low-bias correction necessary for the Lyapunov network to accurately reflect the stability of the current policy $\pi_{\phi}$ using off-policy data. Since the policy parameters are fixed during the Lyapunov update phase, this variance remains manageable via the truncation in (10). However, when optimizing the policy $\pi_{\phi}$ itself, sequence-level IS would introduce prohibitive variance into the gradient estimates, often leading to training divergence. Consequently, following the design principle of PPO, we utilize only the first-step IS ratio $\rho(\phi)$ in (17). This approach effectively reduces the variance of the policy gradient, ensuring stable optimization while still benefiting from the multi-step stability information embedded within the weighted advantage $A_{t,\lambda}$.
IV-C Automated Temperature Adjustment
In maximum entropy RL, the temperature coefficient $\alpha$ in (6) is critical for balancing exploration and exploitation. Following the automated entropy adjustment mechanism introduced in [19], we update $\alpha$ by minimizing the following loss function:
$\mathcal{L}_{\text{Ent}}(\alpha)=\frac{1}{Nn}\sum_{i=1}^{N}\sum_{k=0}^{n-1}\left[-\alpha\log\pi_{\phi}(\mathbf{u}_{t+k}^{i}|\mathbf{x}_{t+k}^{i})-\alpha\bar{\mathcal{H}}\right],$ (18)
where $\bar{\mathcal{H}}$ denotes the target entropy, typically set as $\bar{\mathcal{H}}=-m$, with $m$ representing the dimension of the action space.
By integrating the loss functions for the soft Q-network $Q_{\theta}$, the Lyapunov network $V_{\psi}$, the policy network $\pi_{\phi}$, and the temperature coefficient $\alpha$, we iteratively update the parameters within an actor-critic framework using the Adam optimizer [37]. The complete algorithmic procedure is summarized in Algorithm 1, and the overall flowchart is illustrated in Fig. 2.
Figure 2: Flowchart of the proposed MSACL algorithm.
Algorithm 1 Multi-Step Actor-Critic Learning with Lyapunov Certificates (MSACL)
Double soft Q-networks $Q_{\theta_{1}},Q_{\theta_{2}}$, Lyapunov network $V_{\psi}$, policy network $\pi_{\phi}$, temperature coefficient $\alpha$
Target soft Q-network parameters $\bar{\theta}_{i}\leftarrow\theta_{i},i=1,2$
Replay buffer $\mathcal{D}$, double-ended queue (deque) $\mathcal{DQ}$ with length $n$, delay frequency $d$
for each iteration do
  for each sampling step do
   Select action $\mathbf{u}_{t}\sim\pi_{\phi}(\cdot|\mathbf{x}_{t})$
   Obtain reward $r_{t}$ and observe next state $\mathbf{x}_{t+1}$
   Store tuple $\mathbf{d}_{t}=(\mathbf{x}_{t},\mathbf{u}_{t},r_{t},\pi_{\phi}(\mathbf{u}_{t}|\mathbf{x}_{t}),\mathbf{x}_{t+1})$ in $\mathcal{DQ}$
   if length($\mathcal{DQ}$) =n=n then
     Store $\{\mathbf{d}_{i}\}_{i=t}^{t+n-1}=\{\mathbf{d}_{t},\dots,\mathbf{d}_{t+n-1}\}$ in $\mathcal{D}$
   end if
  end for
  for each update step do
   Sample NN multi-step sequences $\{\mathbf{d}_{i}\}_{i=t}^{t+n-1}$ from $\mathcal{D}$
   Update $V_{\psi}$ via $\psi\leftarrow\psi-\beta_{\psi}\nabla_{\psi}\mathcal{L}_{\text{Lya}}(\psi)$
   Update $Q_{\theta_{i}}$ via $\theta_{i}\leftarrow\theta_{i}-\beta_{\theta}\nabla_{\theta_{i}}\mathcal{L}_{\text{SoftQ}}(\theta_{i})$ 11footnotemark: 1
   if $\left(\text{iteration count}\mod d\right)=0$ then 22footnotemark: 2
     Update $\pi_{\phi}$ via $\phi\leftarrow\phi-\beta_{\phi}\nabla_{\phi}\mathcal{L}_{\pi}(\phi)$
     Update $\alpha$ via $\alpha\leftarrow\alpha-\beta_{\alpha}\nabla_{\alpha}\mathcal{L}_{\text{Ent}}(\alpha)$
   end if
   Update target parameters: $\bar{\theta}_{i}\leftarrow\tau\theta_{i}+(1-\tau)\bar{\theta}_{i}$
  end for
end for
Remark 3
In the flowchart of Algorithm 1, we incorporate two implementation refinements. Specifically, regarding footnote 1: for the soft Q-loss function (14), we replace the target value $Q_{\bar{\theta}}$ with $\min\{Q_{\bar{\theta}_{1}},Q_{\bar{\theta}_{2}}\}$ to mitigate overestimation bias, as noted in Remark 1. Regarding footnote 2: we implement delayed updates for both the policy network $\pi_{\phi}$ and the temperature coefficient $\alpha$. However, within each delayed update cycle, we perform $d$ consecutive updates for these parameters as a “delayed update compensation” mechanism to ensure adequate policy improvement. These techniques are adapted from the cleanrl framework [38], and our experiments demonstrate that integrating these refinements significantly enhances overall performance and training stability.
V Experiments
In this section, we conduct a comprehensive experimental evaluation of the MSACL algorithm across six benchmark environments, encompassing both stabilization and high-dimensional nonlinear tracking tasks. We compare our approach against baseline model-free RL algorithms and Lyapunov-based RL algorithms using rigorous performance and stability metrics. The analysis further includes the visualization of learned Lyapunov certificates, extensive robustness and generalization testing under parametric uncertainties and stochastic noise, and a sensitivity study investigating the impact of the multi-step horizon $n$ on the learning process.
V-A Benchmark Environments
Figure 3: Benchmark environments. (a) Uncontrolled VanderPol (Phase Portrait), (b) Pendulum, (c) Planar DuctedFan, (d) Two-link planar robot, (e) SingleCarTracking, (f) QuadrotorTracking.
In this subsection, we present the benchmark environments utilized to evaluate the experimental performance of the proposed and baseline algorithms, encompassing both stabilization and tracking tasks. We categorize these environments by their primary control objectives. For stabilization tasks, the environments include: Controlled VanderPol (hereafter referred to as VanderPol), Pendulum, Planar DuctedFan (DuctedFan), and Two-link planar robot (Two-link). For tracking tasks, the benchmark comprises SingleCarTracking and QuadrotorTracking. As illustrated in Figure 3, comprehensive specifications and dynamical parameters for all environments are provided in Appendix VI-B.
In stabilization tasks, the control objective is to drive the system state to the equilibrium point, i.e., $\mathbf{x}_{g}=\mathbf{0}\in\mathbb{R}^{n}$. In tracking tasks, the goal is to ensure the system state follows a time-varying reference trajectory $\mathbf{x}_{t}^{\text{ref}}$, which may represent desired coordinates, velocities, or attitudes. Across all tasks, we employ a unified reward structure:
$r_{t}=\underbrace{-\left(\mathbf{x}_{t}^{\top}\mathbf{Q_{r}}\mathbf{x}_{t}+\mathbf{u}_{t}^{\top}\mathbf{R_{r}}\mathbf{u}_{t}\right)}_{\text{Penalty}}+\underbrace{r_{\mathrm{approach}}}_{\text{Encouragement}}$ (19)
where the first term is a penalty term inspired by the performance functionals commonly utilized in optimal control [39], with $\mathbf{Q_{r}}$ and $\mathbf{R_{r}}$ being positive definite weight matrices. This term penalizes deviations from the equilibrium and excessive control effort at each time step, effectively minimizing the cumulative cost $\sum_{t=0}^{\infty}\left[\mathbf{x}_{t}^{\top}\mathbf{Q_{r}}\mathbf{x}_{t}+\mathbf{u}_{t}^{\top}\mathbf{R_{r}}\mathbf{u}_{t}\right]$. For tracking tasks, the state $\mathbf{x}_{t}$ in the penalty term is replaced by the tracking error $\mathbf{e}_{t}=\mathbf{x}_{t}-\mathbf{x}_{t}^{\text{ref}}$. The second term provides sparse encouragement, defined as:
$r_{\mathrm{approach}}=\begin{cases}r_{\delta},&\text{if }\|\mathbf{x}_{t}\|_{\infty}\leq\delta\\ 0,&\text{otherwise}\end{cases}$
Specifically, when the state approaches the equilibrium (i.e., its infinity norm falls below a small threshold $\delta$, such as 0.01), the environment provides a constant positive reward $r_{\delta}$ (e.g., +1) to further incentivize convergence or minimize the steady-state tracking error. This reward design is intentionally parsimonious and intuitive. Subsequent experiments demonstrate that under this straightforward scheme, our algorithm successfully achieves the control objectives across all benchmark environments.
These six benchmark environments are constructed following the OpenAI Gym framework [40]. Notably, our implementation of the Pendulum environment is significantly more rigorous than the standard OpenAI Gym version. Specifically, we expand the initial state space to $\pi,\pi]\times[-10,10]$ and the action range to [5,5][-5,5], while introducing boundary-based termination conditions and extending the maximum episode length to 1000 steps. These modifications increase the control difficulty, intentionally designed to highlight the performance disparities between different algorithms. Detailed specifications for this and other environments are provided in Appendix VI-B. During each agent-environment interaction, the environment returns a standardized 5-tuple $(\mathbf{x}_{t},\mathbf{u}_{t},r_{t},\pi_{\phi}(\mathbf{u}_{t}|\mathbf{x}_{t}),\mathbf{x}_{t+1})$, where the state transition from $\mathbf{x}_{t}$ to $\mathbf{x}_{t+1}$ is updated using the standard explicit Euler integration method.
V-B Baseline Algorithms
To evaluate the effectiveness of the proposed MSACL algorithm, we compare it against classic model-free RL algorithms, including SAC [19] and PPO [24], as well as state-of-the-art Lyapunov-based RL algorithms, specifically LAC [28] and POLYC [27]. While SAC and PPO have demonstrated superior performance across various benchmarks, they are evaluated here on stabilization and tracking tasks to provide a performance baseline. LAC and POLYC incorporate Lyapunov-based constraints into standard RL frameworks to provide stability guarantees; these approaches serve as primary inspirations for our method.
In the LAC algorithm, the authors replace the standard reward signal with a quadratic cost function, which is defined across environments as:
$c_{t}=\mathbf{x}_{t}^{\top}\mathbf{Q_{c}}\mathbf{x}_{t},$
where $\mathbf{Q_{c}}$ is a positive definite matrix. Notably, this formulation considers only state-related penalties, whereas our reward definition $r_{t}$ explicitly accounts for both state deviations and control effort. By comparing against these baselines, we aim to demonstrate the advantages of MSACL in terms of stability metrics and convergence speed.
To ensure a fair and consistent comparison, all algorithms are implemented within the General Optimal control Problem Solver (GOPS) framework [41], an open-source PyTorch-based RL library. Within this unified architecture, the actor and critic networks utilize identical structures consisting of multi-layer perceptrons (MLPs) with two hidden layers of 256 units each and ReLU activation functions. All actor networks employ stochastic policies parameterized by diagonal Gaussian distributions. Regarding the critic architectures: LAC utilizes only a Lyapunov network, while POLYC incorporates both a value function and a Lyapunov certificate. For fairness, SAC and LAC also implement clipped double Q-learning and delayed policy updates as discussed in Remark 1. Parameter updates for all algorithms are performed using the Adam optimizer. The core hyperparameter configurations are summarized in Table I, with full specifications available in our code repository, which will be publicly released.
TABLE I: Hyperparameters in the algorithms
Hyperparameters Value
Shared
   Actor learning rate $3\text{e}-4$
   Critic learning rate $1\text{e}-3$
   Discount factor ($\gamma$) 0.99
   Policy delay update interval ($d$) 2
   Target smoothing coefficient ($\tau$) 0.05
   Reward/Cost scale 100
   Optimizer Adam (default $\beta_{1}$ and $\beta_{2}$)
Maximum-entropy framework
   Learning rate of $\alpha$ $1\text{e}-3$
   Target entropy ($\overline{\mathcal{H}}$) $\overline{\mathcal{H}}=-\text{dim}(\mathcal{U})=-m$
Off-policy
   Replay buffer warm size $5\times 10^{3}$
   Replay buffer size $1\times 10^{6}$
   Samples collected per iteration 20
   Replay batch size 256
On-policy
   Sample batch size 1600
   Replay batch size 1600
   GAE factor 0.95
MSACL
   $\alpha_{1}$ 1
   $\alpha_{2}$ 2
   $\alpha_{3}$ 0.15 (0.125 is also acceptable)
   $\omega_{\text{Bnd}}$ in (12) 1
   $\omega_{\text{Stab}}$ in (12) 10
   $\varepsilon$ in (17) 0.1
V-C Performance Evaluation and Stability Analysis
In this subsection, we present the training outcomes, evaluate stability metrics under the learned policies, and visualize the obtained Lyapunov certificates. To ensure statistical significance, each algorithm was trained using five random seeds per benchmark. The training progress is illustrated in Fig. 4, utilizing the following performance metrics:
  • AMCR (Average Mean Cumulative Reward): The sample mean of the Mean Cumulative Reward (MCR) computed across multiple independent training trials. The MCR is defined as the total undiscounted reward accumulated throughout an episode, normalized by the episode length to provide a consistent per-step measure of policy performance.
  • AMCC (Average Mean Cumulative Cost): The sample mean of the Mean Cumulative Cost (MCC) computed across multiple independent training trials. The MCC is defined as the total undiscounted cost accumulated throughout an episode, normalized by the episode length to quantify the average per-step magnitude of state deviations and control effort.
(a) VanderPol
(b) Pendulum
(c) DuctedFan
(d) Two-link
(e) SingleCarTracking
(f) QuadrotorTracking
MSACL (Ours)            SAC [19]            LAC [28]            PPO [24]            POLYC [27]
Figure 4: Training performance across six benchmark environments. The results show the average MCR and MCC over five independent runs. Solid lines and shaded regions represent the mean and one standard deviation, respectively. The horizontal axis indicates the number of training steps.
The training curves demonstrate that MSACL outperforms or matches all baseline algorithms across all environments. Notably, the AMCR for MSACL remains positive and stable during the late training phase. Given our reward design, a positive MCR indicates that the system state has successfully entered the target $\delta$-neighborhood. Higher reward values further imply faster convergence, as the agent spends a greater proportion of the episode receiving the positive reward $r_{\delta}$.
For subsequent comparative analysis, we select the optimal policy model for each algorithm, defined as the specific model that achieved the highest MCR among the five independent training runs. These optimal policies are evaluated from 100 fixed random initial states across all benchmarks. To rigorously assess stability and convergence, we utilize the following metrics across four target radii ($0.2,0.1,0.05,\text{and }0.01$):
  • Reach Rate (RR): The percentage of trajectories that enter the specified radius.
  • Average Reach Step (ARS): The mean number of steps required to first enter the radius.
  • Average Hold Step (AHS): The mean number of steps the state remains within the radius after the first entry.
The evaluation results are summarized in Table II.
TABLE II: Experimental Results of Baseline Algorithms Across Benchmark Environments
Benchmark Algorithm AMCR AMCC RR / ARS / AHS at different radii
0.2 0.1 0.05 0.01
VanderPol MSACL 33.2 $\pm$ 64.5 31.0 $\pm$ 29.4 1.0 / 57.6 / 943.2 1.0 / 58.4 / 942.5 1.0 / 58.8 / 942.1 1.0 / 59.0 / 941.9
SAC 32.3 $\pm$ 64.4 31.1 $\pm$ 29.5 1.0 / 56.2 / 944.6 1.0 / 58.9 / 942.0 1.0 / 61.0 / 939.9 1.0 / 67.6 / 933.3
LAC 21.2 $\pm$ 65.1 31.1 $\pm$ 29.7 1.0 / 73.1 / 927.8 1.0 / 93.4 / 907.5 1.0 / 112.0 / 888.9 1.0 / 168.6 / 832.3
PPO 27.1 $\pm$ 64.1 30.8 $\pm$ 29.4 1.0 / 65.6 / 935.3 1.0 / 77.8 / 923.1 1.0 / 87.6 / 913.4 1.0 / 110.8 / 890.1
POLYC 32.3 $\pm$ 64.5 30.8 $\pm$ 29.4 1.0 / 59.3 / 941.6 1.0 / 61.6 / 939.3 1.0 / 63.2 / 937.8 1.0 / 66.4 / 934.5
Pendulum MSACL 35.4 $\pm$ 360.1 30.8 $\pm$ 171.7 0.98 / 35.3 / 965.6 0.98 / 40.3 / 960.6 0.98 / 41.2 / 959.7 0.98 / 42.5 / 958.4
SAC -544.4 $\pm$ 1073.4 382.1 $\pm$ 640.0 0.74 / 34.0 / 966.8 0.74 / 36.0 / 964.9 0.74 / 37.0 / 963.9 0.74 / 41.6 / 959.3
LAC -392.6 $\pm$ 899.9 269.2 $\pm$ 503.4 0.78 / 37.6 / 963.3 0.78 / 51.0 / 950.0 0.78 / 64.6 / 936.4 0.78 / 111.6 / 889.4
PPO -6443.4 $\pm$ 4827.7 5843.8 $\pm$ 4912.2 0.0 / – / – 0.0 / – / – 0.0 / – / – 0.0 / – / –
POLYC -2576.9 $\pm$ 4404.2 2379.5 $\pm$ 3986.4 0.69 / 27.8 / 972.4 0.69 / 33.6 / 967.2 0.69 / 37.8 / 963.1 0.69 / 45.4 / 955.5
Ducted Fan MSACL 88.6 $\pm$ 3.5 2.1 $\pm$ 1.3 1.0 / 35.3 / 965.3 1.0 / 52.7 / 948.0 1.0 / 59.6 / 941.3 1.0 / 68.9 / 932.0
SAC -12.6 $\pm$ 433.8 81.1 $\pm$ 347.0 0.95 / 32.3 / 968.1 0.95 / 46.3 / 954.6 0.95 / 55.4 / 945.5 0.95 / 88.6 / 910.9
LAC 84.2 $\pm$ 4.0 2.0 $\pm$ 1.3 1.0 / 40.6 / 959.8 1.0 / 58.7 / 941.8 1.0 / 70.1 / 930.0 1.0 / 102.0 / 898.8
PPO -8.5 $\pm$ 372.7 73.0 $\pm$ 285.4 0.94 / 40.6 / 958.9 0.94 / 59.4 / 941.3 0.94 / 76.6 / 924.1 0.94 / 113.4 / 887.4
POLYC 82.3 $\pm$ 4.3 2.4 $\pm$ 1.4 1.0 / 44.8 / 955.3 1.0 / 68.7 / 932.0 1.0 / 87.4 / 913.1 1.0 / 130.5 / 870.4
Two-link MSACL 89.0 $\pm$ 7.7 0.4 $\pm$ 0.3 1.0 / 14.5 / 986.3 1.0 / 22.0 / 978.9 1.0 / 27.9 / 973.0 1.0 / 41.5 / 957.2
SAC -159.6 $\pm$ 362.5 179.4 $\pm$ 257.3 0.67 / 12.4 / 988.4 0.67 / 18.4 / 982.4 0.67 / 24.3 / 976.6 0.67 / 36.9 / 964.0
LAC 82.5 $\pm$ 8.2 0.3 $\pm$ 0.2 1.0 / 14.4 / 986.5 1.0 / 28.7 / 972.2 1.0 / 45.9 / 955.0 1.0 / 92.6 / 908.3
PPO -676.9 $\pm$ 863.7 472.7 $\pm$ 566.2 0.48 / 12.4 / 988.5 0.48 / 18.1 / 982.6 0.48 / 24.2 / 976.7 0.48 / 46.9 / 954.0
POLYC -172.0 $\pm$ 335.2 190.6 $\pm$ 246.6 0.60 / 14.8 / 939.0 0.60 / 22.8 / 978.1 0.60 / 28.9 / 972.0 0.60 / 40.2 / 960.7
SingleCarTracking MSACL 84.7 $\pm$ 32.7 2.9 $\pm$ 14.1 0.99 / 37.6 / 961.6 0.99 / 55.3 / 940.7 0.99 / 76.5 / 924.1 0.99 / 96.4 / 904.5
SAC 10.3 $\pm$ 164.9 21.9 $\pm$ 50.3 0.85 / 36.5 / 962.9 0.85 / 54.3 / 946.6 0.85 / 81.4 / 919.5 0.85 / 193.3 / 807.6
LAC -350.5 $\pm$ 196.8 94.5 $\pm$ 47.4 0.17 / 14.2 / 986.7 0.17 / 51.2 / 949.6 0.17 / 107.5 / 891.0 0.17 / 221.1 / 763.4
PPO 10.2 $\pm$ 188.6 21.5 $\pm$ 49.2 0.85 / 38.6 / 961.8 0.85 / 61.2 / 937.6 0.85 / 80.6 / 919.1 0.85 / 102.7 / 894.4
POLYC 40.1 $\pm$ 152.8 14.7 $\pm$ 39.9 0.90 / 34.9 / 962.0 0.90 / 58.0 / 939.8 0.90 / 67.6 / 932.3 0.90 / 81.8 / 912.3
QuadrotorTracking MSACL 838.8 $\pm$ 1.5 14.5 $\pm$ 2.1 1.0 / 33.1 / 967.8 1.0 / 34.0 / 966.9 1.0 / 36.9 / 964.0 1.0 / 49.7 / 619.8
SAC -17263.8 $\pm$ 1380.0 17246.4 $\pm$ 1379.1 0.0 / – / – 0.0 / – / – 0.0 / – / – 0.0 / – / –
LAC -51.1 $\pm$ 1.2 32.7 $\pm$ 1.1 1.0 / 471.9 / 20.6 0.0 / – / – 0.0 / – / – 0.0 / – / –
PPO -12836.6 $\pm$ 4042.3 12571.3 $\pm$ 4022.4 0.0 / – / – 0.0 / – / – 0.0 / – / – 0.0 / – / –
POLYC -17528.1 $\pm$ 2162.6 17318.7 $\pm$ 2155.6 0.0 / – / – 0.0 / – / – 0.0 / – / – 0.0 / – / –
  • Note: The values for AMCR and AMCC represent the mean performance of the optimal policy (defined as the policy achieving the highest evaluation MCR) averaged over 100 independent trials, with $\pm$ denoting one standard deviation. Stability metrics (RR, ARS, and AHS) are evaluated across four distinct radii, where “–” indicates that the system failed to enter the specified radius within the episode duration. Bold entries highlight the superior results: for performance metrics, the maximum AMCR and minimum AMCC are bolded; for stability metrics, the highest RR is prioritized for bolding, with the shortest ARS and longest AHS selected in the event of identical RR values.
These results highlight the superiority of MSACL. Across all six environments, MSACL consistently achieves the highest AMCR. In terms of AMCC, MSACL attains the absolute minimum in the Pendulum, SingleTrackCar, and QuadrotorTracking tasks, while remaining highly competitive in others. Furthermore, MSACL achieves the highest RR across all radii. Most significantly, in the high-dimensional nonlinear QuadrotorTracking task, MSACL is the only method to achieve a 100% reach rate at the $0.01$ precision level, whereas baselines fail to meet this stringent requirement. These results underscore the robustness and precision of MSACL in complex control applications.
Finally, we visualize the Lyapunov networks $V_{\psi}$ corresponding to the optimal policies. The learned Lyapunov certificates and their associated contour plots are shown in Fig. 5. For high-dimensional systems, the plots represent the relationship between $V_{\psi}$ and two specific state components while setting all other components to zero.
Remark 4
Since MSACL is built upon the SAC framework, SAC serves as the primary baseline for our ablation study. As shown in Fig. 4 and Table II, the integration of the Lyapunov certificate significantly enhances both training performance and stability metrics compared to the standard SAC algorithm. This validates that embedding Lyapunov-based constraints into maximum entropy RL effectively yields neural policies with superior stability guarantees for deterministic control systems.
Figure 5: Visualization of learned Lyapunov certificates and their corresponding contour plots across various benchmark environments: (a) VanderPol; (b) Pendulum; (c)–(e) DuctedFan (projections in the $(x,y)$-plane with $\theta=0$, $(y,\theta)$-plane with $x=0$, and $(x,\theta)$-plane with $y=0$); (f) Two-link; (g) SingleCarTracking; (h)–(j) QuadrotorTracking (projections in the $(e_{x},e_{y})$-plane with $e_{z}=0$, $(e_{y},e_{z})$-plane with $e_{x}=0$, and $(e_{x},e_{z})$-plane with $e_{y}=0$).
V-D Evaluation of Robustness and Generalization
It is well known that neural policies, characterized by high-dimensional parameter spaces, are susceptible to overfitting within specific training environments. This can lead to degraded robustness against stochastic noise and parametric variations, as well as limited generalization to unseen scenarios. In this subsection, we evaluate the robustness and generalization capabilities of the policies trained via MSACL to demonstrate their applicability in real-world environments with inherent uncertainties.
For stabilization tasks (VanderPol, Pendulum, DuctedFan, and Two-link), we introduce process noise $\bm{\epsilon}_{t}\sim\mathcal{N}(\mathbf{0},\bm{\sigma}_{\text{env}})$ to the system dynamics, resulting in the perturbed form:
$\mathbf{x}_{t+1}=f(\mathbf{x}_{t},\mathbf{u}_{t})+\bm{\epsilon}_{t}.$
Additionally, we modify key physical parameters of the systems to assess whether the MSACL-trained policies can maintain stability under environmental variations.
For tracking tasks (SingleCarTracking and QuadrotorTracking), we evaluate robustness by injecting dynamical noise and assess generalization by requiring the policies to track reference trajectories $\mathbf{x}_{t}^{\text{ref}}$ that differ significantly from those encountered during training. The specific parameter modifications and noise magnitudes utilized for these evaluations are detailed in Table III.
TABLE III: System Parameters and Noise Magnitudes
Benchmarks System Parameters Noise Magnitude ($\sigma_{\mathrm{env}}$)
VanderPol $\mu=0.5,1.5$ $0.5,0.8$
Pendulum $L=0.5,1.0$ $0.5,1.0$
DuctedFan $m=5,12$ $0.1,0.3$
Two-link $l_{1}=0.75,l_{2}=1.5$ $0.1,0.2$
SingleCarTracking $0.1,0.3$
QuadrotorTracking $0.01,0.02$
The robustness testing results for the stabilization tasks are illustrated in Fig. 6. Despite the introduction of parametric perturbations and process noise, all system trajectories originating from the specified initial states successfully converge to the neighborhood of the equilibrium $\mathbf{x}_{g}=\mathbf{0}$.
The results for the tracking tasks are presented in Fig. 7 and Fig. 8. It is evident that the trained neural policies effectively mitigate noise interference within a reasonable range. In the SingleCarTracking environment, while the policy was trained on straight-line references, we evaluate its generalization using circular and sinusoidal trajectories. Experimental validation demonstrates excellent tracking performance; as shown in the position error plot in Fig. 7e, the tracking error $e_{\mathrm{xy}}=\sqrt{e_{x}^{2}+e_{y}^{2}}$ remains within 0.1 during the steady-state phase.
For the QuadrotorTracking task, the policy was trained using a horizontal helical trajectory. During generalization testing, we evaluate the policy against two unseen trajectories: a vertical helix and a Lissajous curve, as shown in Fig. 8c and Fig. 8d. Notably, the MSACL-trained policy demonstrates superior generalization, with the tracking error $e_{\mathrm{xyz}}=\sqrt{e_{x}^{2}+e_{y}^{2}+e_{z}^{2}}$ stabilizing around 0.02 in the later stages.
Figure 6: System trajectories under the neural policies in stabilization tasks subjected to environmental parametric uncertainties and process noise: (a) VanderPol with state variables $\mathbf{x}=[x,\dot{x}]^{\top}$; (b) Pendulum with angular position $\theta$ and angular velocity $\dot{\theta}$; (c) DuctedFan with coordinates $x,y$ and orientation $\theta$; (d) Two-link with joint angles $q_{1}$ and $q_{2}$. All trajectories exhibit robust convergence to the neighborhood of the equilibrium $\mathbf{x}_{g}=\mathbf{0}$ despite external disturbances.
Figure 7: Robustness (a)–(b) and generalization (c)–(e) evaluation for the SingleCarTracking task: (a) Car trajectories versus the straight-line reference $\mathbf{x}_{t}^{\text{ref}}=\{[t,0]\in\mathbb{R}^{2}\mid t\in[0,50]\}$ under varying noise levels; (b) Evolution of the Euclidean position error $e_{\mathrm{xy}}=\sqrt{e_{x}^{2}+e_{y}^{2}}$ corresponding to (a); (c) Tracking performance for a circular reference trajectory $\{[8\cos(0.125t),8\sin(0.125t)]^{\top}\in\mathbb{R}^{2}\mid t\in[0,50]\}$ using the policy trained on straight lines; (d) Tracking performance for a sinusoidal reference trajectory $\{[t,\sin(0.2t)]^{\top}\in\mathbb{R}^{2}\mid t\in[0,50]\}$; (e) Evolution of the position error $e_{\mathrm{xy}}$ for the circular and sinusoidal cases shown in (c) and (d). The results demonstrate robust convergence and effective generalization to unseen path geometries.
Figure 8: Robustness (a)–(b) and generalization (c)–(e) evaluation for the QuadrotorTracking task: (a) Quadrotor trajectories versus the horizontal helical reference $\mathbf{x}_{t}^{\text{ref}}=\{[0.4t,0.4\sin t,0.6\cos t]^{\top}\mid t\in[0,40]\}$ under varying noise levels; (b) Evolution of the position error $e_{\mathrm{xyz}}=\sqrt{e_{x}^{2}+e_{y}^{2}+e_{z}^{2}}$ corresponding to (a); (c) Tracking performance for a vertical helical reference $\{[0.5\cos t,0.5\sin t,0.3t]^{\top}\in\mathbb{R}^{3}\mid t\in[0,40]\}$ using the policy trained on the horizontal helix; (d) Tracking performance for a Lissajous reference trajectory $\{[0.8\sin(0.4t),0.8\cos(0.4t),0.4\sin(1.2t)+1.0]^{\top}\in\mathbb{R}^{3}\mid t\in[0,40]\}$; (e) Evolution of the position error $e_{\mathrm{xyz}}$ for the cases in (c) and (d). The results highlight the policy’s robust tracking capability under environmental noise and its effective generalization to geometrically distinct reference trajectories.
Based on these results, we conclude that policies trained via the MSACL algorithm exhibit robust performance against moderate uncertainties and favorable generalization to unseen reference trajectories. However, our testing also reveals that the policies’ efficacy diminishes under extreme disturbances or environmental parameters that deviate substantially from the training distribution. In such cases of significant domain shift, the MSACL algorithm can be readily redeployed to retrain the policy, ensuring stable and high-precision control for the specific operational scenario.
V-E Sensitivity Analysis of the Multi-Step Horizon $n$
In this subsection, we investigate the influence of the multi-step horizon parameter $n$ on the performance of the MSACL algorithm across all benchmark environments. Following the evaluation protocol detailed in Section V-C, we utilize both performance metrics (AMCR and AMCC) and stability metrics (RR, ARS, and AHS) to assess the policies trained under varying configurations of $n$. Specifically, we conduct 100 randomized experiments for each benchmark by generating and fixing a set of 100 random initial states. Policies trained with different $n$ values interact with the environment starting from these fixed states to compute the aforementioned evaluation metrics.
The results across all environments are summarized in Table IV. The experimental findings reveal a consistent trend: as $n$ increases, the stabilization and tracking performance of the MSACL-trained policies generally improves. This improvement occurs because a larger $n$ provides a longer-horizon trajectory during training, enabling the policy $\pi$ to more accurately identify the stability direction, namely the descent direction of the Lyapunov certificate $V_{\psi}$. This reduces the training bias and reinforces the stability guarantees for the resulting neural policy.
Empirically, setting $n=20$ yields satisfactory performance across all benchmark environments, suggesting it as a versatile default parameter. However, an excessively large $n$ may not always lead to optimal results. As analyzed in Section III-B, longer temporal horizons introduce greater cumulative variance into the training data, which can distort the optimization landscape and result in suboptimal policies. Therefore, selecting an appropriate $n$ is essential to balancing the bias-variance trade-off. In summary, for the control tasks presented in this article, $n=20$ delivers consistently robust results; for specific applications requiring peak performance, fine-tuning around this value is recommended to achieve the optimal balance between training bias and variance.
TABLE IV: Experimental Results of MSACL Algorithm with different multi-step horizon Parameters $n$
Benchmark $n$ Values AMCR AMCC RR / ARS / AHS at different radii
0.2 0.1 0.05 0.01
VanderPol $n=1$ 30.2 $\pm$ 64.5 31.1 $\pm$ 29.5 1.0 / 56.3 / 944.6 1.0 / 61.9 / 939.0 1.0 / 68.9 / 932.0 1.0 / 88.2 / 912.7
$n=5$ 31.5 $\pm$ 64.5 30.9 $\pm$ 29.4 1.0 / 58.4 / 942.4 1.0 / 63.4 / 937.5 1.0 / 67.6 / 933.3 1.0 / 78.7 / 922.2
$n=10$ 32.4 $\pm$ 64.6 30.9 $\pm$ 29.4 1.0 / 57.9 / 942.9 1.0 / 61.4 / 939.5 1.0 / 64.1 / 936.8 1.0 / 68.9 / 932.0
$n=15$ 33.1 $\pm$ 64.4 31.2 $\pm$ 29.5 1.0 / 54.1 / 946.7 1.0 / 55.1 / 945.8 1.0 / 55.7 / 945.3 1.0 / 57.0 / 943.9
$n=20$ 33.2 $\pm$ 64.5 31.0 $\pm$ 29.5 1.0 / 57.8 / 943.1 1.0 / 58.7 / 942.2 1.0 / 59.2 / 941.7 1.0 / 59.5 / 941.4
Pendulum $n=1$ -578.5 $\pm$ 1232.9 444.2 $\pm$ 815.1 0.77 / 37.1 / 963.5 0.77 / 42.4 / 958.6 0.77 / 44.4 / 956.5 0.77 / 48.8 / 952.1
$n=5$ 34.7 $\pm$ 360.0 30.7 $\pm$ 171.8 0.98 / 36.8 / 964.0 0.98 / 44.7 / 956.2 0.98 / 46.0 / 954.9 0.98 / 49.4 / 951.5
$n=10$ 35.7 $\pm$ 360.1 31.0 $\pm$ 171.7 0.98 / 34.9 / 966.0 0.98 / 35.6 / 965.4 0.98 / 36.1 / 964.9 0.98 / 38.5 / 962.3
$n=15$ 35.2 $\pm$ 360.0 31.0 $\pm$ 171.7 0.98 / 35.2 / 965.6 0.98 / 36.9 / 964.0 0.98 / 38.4 / 962.6 0.98 / 44.0 / 956.9
$n=20$ 35.8 $\pm$ 355.7 30.3 $\pm$ 167.9 0.98 / 38.0 / 962.8 0.98 / 39.3 / 961.6 0.98 / 40.7 / 960.2 0.98 / 44.1 / 956.8
Ducted Fan $n=1$ -116.3 $\pm$ 584.8 163.2 $\pm$ 463.3 0.89 / 32.3 / 967.8 0.89 / 49.8 / 950.6 0.89 / 57.0 / 943.9 0.89 / 83.5 / 917.3
$n=5$ 87.4 $\pm$ 3.7 2.0 $\pm$ 1.3 1.0 / 35.5 / 965.2 1.0 / 49.8 / 950.9 1.0 / 57.9 / 942.9 1.0 / 78.4 / 922.1
$n=10$ 88.6 $\pm$ 3.5 2.1 $\pm$ 1.3 1.0 / 35.3 / 965.3 1.0 / 52.7 / 948.0 1.0 / 59.6 / 941.3 1.0 / 68.9 / 932.0
$n=15$ 88.1 $\pm$ 4.0 2.0 $\pm$ 1.3 1.0 / 36.1 / 964.6 1.0 / 52.0 / 949.0 1.0 / 61.4 / 939.5 1.0 / 74.3 / 926.6
$n=20$ 73.7 $\pm$ 37.7 4.4 $\pm$ 8.5 0.95 / 56.3 / 944.2 0.95 / 79.2 / 921.6 0.95 / 94.5 / 906.4 0.95 / 127.0 / 873.9
Two-link $n=1$ -208.6 $\pm$ 372.5 216.0 $\pm$ 267.0 0.60 / 13.8 / 987.2 0.60 / 22.3 / 976.7 0.60 / 29.6 / 971.3 0.60 / 40.4 / 960.5
$n=5$ -142.4 $\pm$ 340.0 167.6 $\pm$ 241.6 0.67 / 13.7 / 987.0 0.67 / 20.3 / 980.4 0.67 / 24.7 / 976.1 0.67 / 29.9 / 971.0
$n=10$ -163.8 $\pm$ 346.6 183.1 $\pm$ 247.1 0.64 / 10.7 / 990.2 0.64 / 16.4 / 984.5 0.64 / 23.2 / 977.6 0.64 / 37.0 / 963.9
$n=15$ -131.0 $\pm$ 354.1 158.3 $\pm$ 251.4 0.71 / 12.1 / 988.7 0.71 / 18.5 / 982.3 0.71 / 24.8 / 976.1 0.71 / 40.6 / 960.3
$n=20$ 89.8 $\pm$ 7.5 0.4 $\pm$ 0.2 1.0 / 18.1 / 982.0 1.0 / 25.0 / 975.8 1.0 / 28.6 / 972.2 1.0 / 35.2 / 965.7
SingleTrackCar $n=1$ -51.0 $\pm$ 194.9 46.4 $\pm$ 62.6 0.64 / 32.0 / 964.0 0.64 / 64.6 / 931.1 0.64 / 79.7 / 921.2 0.64 / 92.1 / 908.8
$n=5$ -40.2 $\pm$ 217.9 35.0 $\pm$ 58.8 0.74 / 31.6 / 968.7 0.74 / 56.1 / 943.8 0.74 / 78.9 / 920.9 0.74 / 113.0 / 887.8
$n=10$ 84.7 $\pm$ 32.7 2.9 $\pm$ 14.1 0.99 / 37.6 / 961.6 0.99 / 55.3 / 940.7 0.99 / 76.5 / 924.1 0.99 / 96.4 / 904.5
$n=15$ 67.5 $\pm$ 84.7 8.8 $\pm$ 32.5 0.95 / 35.2 / 964.8 0.95 / 57.8 / 938.7 0.95 / 82.6 / 917.0 0.95 / 109.1 / 891.7
$n=20$ 82.6 $\pm$ 55.1 3.0 $\pm$ 15.6 0.99 / 37.6 / 960.5 0.99 / 63.1 / 934.3 0.99 / 85.0 / 915.8 0.99 / 95.0 / 905.6
QuadrotorTracking $n=1$ -12708.9 $\pm$ 645.4 12582.7 $\pm$ 645.2 0.0 / – / – 0.0 / – / – 0.0 / – / – 0.0 / – / –
$n=5$ -4708.7 $\pm$ 3158.6 4646.3 $\pm$ 3135.9 0.0 / – / – 0.0 / – / – 0.0 / – / – 0.0 / – / –
$n=10$ 534.3 $\pm$ 5.0 11.8 $\pm$ 0.9 1.0 / 54.4 / 946.5 1.0 / 66.2 / 934.7 1.0 / 116.5 / 650.9 0.0 / – / –
$n=15$ 858.5 $\pm$ 1.0 10.8 $\pm$ 0.5 1.0 / 31.3 / 969.6 1.0 / 36.8 / 964.1 1.0 / 41.2 / 959.7 1.0 / 115.4 / 747.9
$n=20$ 838.5 $\pm$ 1.5 14.7 $\pm$ 2.0 1.0 / 33.1 / 967.8 1.0 / 34.1 / 966.8 1.0 / 37.0 / 963.9 1.0 / 49.8 / 620.0
VI Discussion and Conclusion
VI-A Discussion
This subsection provides an analytical discussion on the functional roles of the various components within the MSACL framework. As highlighted in Remark 4, MSACL extends the Soft Actor-Critic (SAC) algorithm by integrating multi-step Lyapunov certificate learning, thereby establishing formal stability guarantees while preserving the exploration efficiency of maximum entropy reinforcement learning. In model-free stabilization and tracking tasks, conventional SAC, which lacks the structural guidance provided by Lyapunov theory, may fail to converge to stable policies under improper reward structures. Even with well-designed rewards, standard RL often requires exhaustive exploration of the action space to identify stabilizing directions, leading to prohibitive training times. To address these limitations, our approach introduces a Lyapunov network that serves as a stability-oriented critic, providing explicit directional guidance during policy updates by identifying action sequences that facilitate the descent of the Lyapunov certificate.
The theoretical foundation of this mechanism is rooted in Lemma 1, which establishes the equivalence between the Lyapunov stability conditions and exponential state convergence. To ensure the Lyapunov certificate $V_{\psi}$ is learned accurately, we developed a training methodology that aligns theoretical requirements with empirical trajectories. Specifically, the introduction of Exponential Stability Labels (ESL) in Definition 2 allows the algorithm to categorize trajectory segments as positive or negative samples based on their adherence to exponential decay criteria. This classification enables the construction of the loss function (11), which effectively bridges abstract stability conditions with observed system dynamics.
The policy optimization process leverages this learned certificate through the stability advantage mechanism defined in (17). This advantage function quantifies the rate of Lyapunov descent along trajectories, prioritizing policy updates that simultaneously ensure stability and accelerate convergence. Our experimental results confirm that this dual-objective optimization yields policies that achieve both stability and rapid convergence across diverse and challenging benchmarks. A critical insight is that the efficacy of the policy is fundamentally tied to the fidelity of the learned Lyapunov certificate. The multi-step sampling strategy, combined with the ESL-based loss formulation, ensures that $V_{\psi}$ accurately reflects the theoretical stability criteria, providing reliable gradients for policy improvement.
While developed for stabilization and tracking, the MSACL framework is highly extensible to other stability-critical applications. Potential extensions include trajectory planning for autonomous vehicles [42], spacecraft proximity operations [43], and robotic manipulation tasks requiring rigorous safety guarantees [44, 45]. The integration of Lyapunov-based guarantees within maximum entropy RL represents a significant step toward developing verifiable, learning-based control systems for real-world deployment.
VI-B Conclusion
This article presented the MSACL (Multi-Step Actor-Critic Learning with Lyapunov certificates) algorithm, a novel framework that integrates exponential stability guarantees into the maximum entropy RL paradigm. By embedding Lyapunov stability theory within the SAC architecture, MSACL achieves stable control for model-free dynamical systems without sacrificing the exploration benefits of entropy regularization.
To rigorously evaluate the proposed methodology, we developed six benchmark environments covering four stabilization and two tracking tasks. Comprehensive comparisons against established baselines demonstrate the consistent superiority of MSACL across both performance and stability metrics. Notably, these results were achieved using a simple reward structure that penalizes state deviations and control effort while providing sparse positive rewards near the equilibrium. Beyond standard performance, we conducted extensive robustness analyses against environmental uncertainties and generalization tests on unseen reference trajectories. The MSACL-trained neural policies exhibited remarkable resilience to disturbances and effective adaptation to geometrically distinct paths.
A key algorithmic contribution is the systematic analysis of the multi-step horizon parameter $n$. We established that $n=20$ serves as a robust default setting, effectively balancing the bias-variance trade-off in stability learning. This work advances the integration of formal stability guarantees within learning-based control, providing a foundation for verifiable policies in safety-critical applications. Future research will explore extensions to partially observable systems [46, 47], adaptive Lyapunov architectures [48, 49], and transfer learning mechanisms for complex, cross-domain dynamical systems [50, 51].
[Benchmark environments information]
This appendix provides a detailed description of the six nonlinear dynamical systems selected as benchmarks for this study: VanderPol, Pendulum, DuctedFan, Two-link, SingleCarTracking, and QuadrotorTracking. Widely recognized as canonical benchmarks in nonlinear control (NLC) [52, 53, 54], reinforcement learning (RL) [55, 56], and optimal control literature [57], these environments constitute a comprehensive testbed for evaluating the proposed methods. They encompass a broad spectrum of dynamic characteristics, ranging from self-sustaining oscillations and underactuated mechanics to complex tire–road friction interactions and rigid-body dynamics on the $\text{SE}(3)$ manifold.
To provide a unified evaluation framework, we categorize these benchmarks based on their control objectives: Stabilization and Tracking. The formal task definitions and their corresponding reward structures are detailed below.
-C Unified Control Objectives
Despite the distinct physical dynamics inherent to each system, the core control tasks are formulated using a consistent mathematical structure.
Stabilization Tasks [58, 55, 59]
The first four systems (VanderPol, Pendulum, DuctedFan, and Two-link) are configured as stabilization tasks. For these systems, the control objective is to find a policy that generates a sequence of control inputs $\{\mathbf{u}_{t}\}_{t=0}^{T-1}$ to drive the system state $\mathbf{x}_{t}$ from a randomly sampled initial state $\mathbf{x}_{0}$ toward a fixed equilibrium point $\mathbf{x}_{g}$ (typically the origin, $\mathbf{x}_{g}=\mathbf{0}$) and maintain stability at this equilibrium for the duration of the episode.
Tracking Tasks [60, 54, 61]
The remaining two systems (SingleCarTracking, and QuadrotorTracking) are formulated as trajectory tracking tasks. Here, the objective is defined in terms of the tracking error $\mathbf{e}_{t}=\mathbf{x}_{t}-\mathbf{x}_{t}^{\text{ref}}$. The goal is to determine a control sequence such that the error trajectory $\{\mathbf{e}_{t}\}_{t=0}^{T}$ converges to the zero-error state $\mathbf{e}_{g}=\mathbf{0}$, thereby ensuring that the system state $\mathbf{x}_{t}$ accurately follows a time-varying reference trajectory $\mathbf{x}_{t}^{\text{ref}}$.
-D Unified Reward Function Design
To align the learning process with the aforementioned control objectives, we employ a consistent reward function structure across all six environments:
$r_{t}=\underbrace{-\left(\mathbf{s}_{t}^{\top}\mathbf{Q}_{r}\mathbf{s}_{t}+\mathbf{u}_{t}^{\top}\mathbf{R}_{r}\mathbf{u}_{t}\right)}_{\text{Penalty}}+\underbrace{r_{\text{approach}}(\mathbf{s}_{t})}_{\text{Encouragement}},$
where $\mathbf{s}_{t}$ represents the generalized state vector, defined as $\mathbf{s}_{t}=\mathbf{x}_{t}$ for stabilization tasks and $\mathbf{s}_{t}=\mathbf{e}_{t}$ for tracking tasks. The matrices $\mathbf{Q}_{r}$ and $\mathbf{R}_{r}$ are positive semi-definite weighting matrices that penalize state deviations and control effort, respectively.
The specific parameter configurations for each environment, including the diagonal elements of $\mathbf{Q}_{r}$ and $\mathbf{R}_{r}$, are summarized in Table V. The term $\mathbb{I}(\cdot)$ denotes the indicator function, which provides a sparse encouragement reward when the system enters a predefined tolerance region around the target state. For all environments, the state cost matrix $\mathbf{Q}_{c}$ is simply defined as a identity matrix $\mathbf{I}_{n}$ of appropriate dimensions, where $n$ denotes the dimension of state space.
TABLE V: Reward Function Parameters for All Benchmark Systems
System State Penalty $\mathbf{Q}_{r}$ (diag) Control Penalty $\mathbf{R}_{r}$ (diag) Encouragement ($r_{\text{approach}}$)
VanderPol $[2,1]$ $[0.1]$ $\mathbb{I}(\|\mathbf{x}_{t}\|_{\infty}\leq 10^{-2})$
Pendulum $[2,1]$ $[0.1]$ $\mathbb{I}(\|\mathbf{x}_{t}\|_{\infty}\leq 10^{-2})$
DuctedFan $[2,2,2,1,1,1]$ $[0.1,0.1]$ $\mathbb{I}(\|\mathbf{x}_{t}\|_{\infty}\leq 10^{-2})$
Two-link $[2,2,1,1]$ $[0.1,0.1]$ $\mathbb{I}(\|\mathbf{x}_{t}\|_{\infty}\leq 10^{-2})$
SingleCarTracking $[2,2,1,1,1,1,1]$ $[0.1,0.1]$ $\mathbb{I}(\|\mathbf{e}_{t}\|_{\infty}\leq 10^{-2})$
QuadrotorTracking $\mathbf{I}_{12}$ $10^{-4},0.01,0.01,0.01]$ $10(1-\frac{\|\mathbf{e}_{t}\|_{\infty}}{0.1})\cdot\mathbb{I}(\|\mathbf{e}_{t}\|_{\infty}\leq 0.1)$
-E Details of VanderPol
-E1 System Overview
The VanderPol oscillator, originally proposed by Balthasar van der Pol in 1920 [62], is a classic model describing self-sustaining oscillatory circuits characterized by nonlinear damping. This system serves as a fundamental benchmark for electronic circuit stabilization (e.g., vacuum tube oscillators), mechanical vibration control, and biological modeling, such as cardiac pacemaker dynamics [63]. In this study, we extend the autonomous system with a control input to evaluate stabilization performance within a nonlinear regime.
-E2 Dynamics and Physical Parameters
The continuous-time dynamics of controlled VanderPol are governed by the following second-order nonlinear differential equations:
$\dot{x}_{1}$ $=x_{2},$
$\dot{x}_{2}$ $=\mu\left(1-x_{1}^{2}\right)x_{2}-x_{1}+u,$
where $x_{1}$ and $x_{2}$ represent the oscillator displacement and velocity, respectively. The physical meanings of the variables are summarized in Table VI, and the specific parameter values and simulation settings used in our experiments are detailed in Table VII.
TABLE VI: Variable Definitions for Controlled VanderPol
Variable Physical Meaning
$x_{1}$ Displacement of the oscillator (m)
$x_{2}$ Velocity of the oscillator (m/s)
$\dot{x}_{2}$ Acceleration of the oscillator (m/s2)
$\mu$ Nonlinear damping strength parameter (dimensionless)
$u$ Control input force applied to the oscillator (N)
TABLE VII: Parameter Values for Controlled VanderPol
Parameter Value
$\mu$ (nonlinear strength) 1.0 (dimensionless)
Simulation time step ($dt$) 0.01 s
Control update interval 5 simulation steps
Maximum episode steps 1000
-E3 State and Action Spaces
For the reinforcement learning formulation, the system state vector is defined as $\mathbf{x}=[x_{1},x_{2}]^{\top}$. To reflect the physical limits of the oscillator and ensure numerical stability during training, the state space is constrained to $\mathcal{X}=[-10,10]\times[-10,10]$, where both displacement and velocity are bounded within the range of $\pm 10$. The control objective is to drive the system from a random initial configuration to the equilibrium state (goal state) defined as $\mathbf{x}_{g}=[0,0]^{\top}$, corresponding to the stable origin of the phase plane.
The control action $u$ is a one-dimensional continuous input restricted to the action space $\mathcal{U}=[-5,5]$. These bounds are selected to provide sufficient authority for stabilization while preventing excessively large inputs that could violate physical actuation constraints or lead to divergent gradients. At the start of each episode, the initial state $\mathbf{x}_{0}$ is sampled from a uniform distribution $\mathbf{x}_{0}\sim\text{Uniform}([-5,5]\times[-5,5])$. This initialization range is intentionally chosen to place the system within a significant nonlinear regime, ensuring that the controller must actively counteract the self-sustaining oscillations to achieve stabilization.
-F Details of Pendulum
-F1 System Overview
The Pendulum is a classic rigid-body system with a pivot point constrained to a vertical plane, where the primary control objective is to stabilize the system in an inverted (upright) position against gravity. Beyond its role as a fundamental pedagogical tool for illustrating nonlinear stability and feedback control [64], its underlying principles are essential for robotic locomotion in bipedal systems [65], aerospace attitude control, and industrial crane load stabilization [58].
-F2 Dynamics and Physical Parameters
The continuous-time dynamics of the Pendulum are derived from Newtonian mechanics and are defined by the following equations of motion:
$\dot{\theta}$ $=\dot{\theta},$
$\ddot{\theta}$ $=\frac{mgL\sin(\theta)-b\dot{\theta}+u}{mL^{2}},$
where $\theta$ and $\dot{\theta}$ denote the angular displacement and angular velocity, respectively. The physical variables and their definitions are summarized in Table VIII, while the specific parameter values and simulation settings used in this study are provided in Table IX.
TABLE VIII: Variable Definitions for Pendulum
Variable Physical Meaning
$\theta$ Angular displacement of the pendulum (rad)
$\dot{\theta}$ Angular velocity of the pendulum (rad/s)
$\ddot{\theta}$ Angular acceleration of the pendulum (rad/s2)
$m$ Mass of the pendulum bob (kg)
$g$ Gravitational acceleration (m/s2)
$L$ Length of the pendulum rod (m)
[…]
TABLE IX: Physical Parameter Values for Pendulum
Parameter Value
$g$ 9.81 m/s2
$L$ 0.5 m
$m$ 0.15 kg
[…]
Simulation time step ($dt$) 0.01 s
Control update interval 5 simulation steps
Maximum episode steps 1000
-F3 State and Action Spaces
The system state vector is defined as $\mathbf{x}=[\theta,\dot{\theta}]^{\top}$, with the state space constrained to $\mathcal{X}=[-\pi,\pi]\times[-10,10]$. This range covers the full angular displacement and accounts for physical angular velocity limits. The control objective is to drive the system to the unstable equilibrium point (upright position) defined as $\mathbf{x}_{g}=[0,0]^{\top}$.
The control action $u$ is a one-dimensional continuous torque restricted to the action space $\mathcal{U}=[-5,5]$. These torque bounds are selected to ensure the system has sufficient authority for the swing-up maneuver while remaining within realistic actuation limits. At the start of each episode, the initial state $\mathbf{x}_{0}$ is uniformly sampled from the entire state space, $\mathbf{x}_{0}\sim\text{Uniform}([-\pi,\pi]\times[-10,10])$, requiring the controller to handle both stabilization from near-upright positions and complex swing-up maneuvers from high-energy initial states.
-F4 Comparison with OpenAI Gym Pendulum
Our Pendulum implementation is significantly more rigorous than the standard OpenAI Gym version [40]. Specifically, we expand the initial state space to the full range $\pi,\pi]\times[-10,10]$ and increase the actuation limit to [5,5][-5,5], whereas the Gym environment utilizes a narrower velocity initialization ([1,1][-1,1]) and limited torque range ([2,2][-2,2]). Furthermore, we impose strict boundary-based termination and extend the episode horizon to $T=1000$ steps. These more stringent constraints are intentionally designed to expose the limitations of standard RL baselines; for instance, while PPO converges readily in the Gym setting, it often fails to achieve robust stabilization in our high-energy regime.
-G Details of DuctedFan
-G1 System Overview
The DuctedFan system models the coupled translational and rotational dynamics of a ducted-fan aircraft under the influence of gravitational and control forces [66]. As a canonical benchmark for Vertical Takeoff and Landing (VTOL) research, it provides a rigorous platform for validating hover stabilization and attitude control algorithms [61, 67]. The system is underactuated and exhibits strong nonlinear coupling between its lateral displacement and angular orientation.
-G2 Dynamics and Physical Parameters
The continuous-time dynamics of the DuctedFan are derived from Newtonian mechanics and rigid-body dynamics, defined by the following state-space equations:
$\dot{x}$ $=\dot{x},$
$\dot{y}$ $=\dot{y},$
$\dot{\theta}$ $=\dot{\theta},$
$\ddot{x}$ $=\frac{-mg\sin(\theta)-d\dot{x}+u_{1}\cos(\theta)-u_{2}\sin(\theta)}{m},$
$\ddot{y}$ $=\frac{mg(\cos(\theta)-1)-d\dot{y}+u_{1}\sin(\theta)+u_{2}\cos(\theta)}{m},$
$\ddot{\theta}$ $=\frac{ru_{1}}{J},$
where $u_{1}=f_{1}$ and $u_{2}=f_{2}-mg$, with $f_{1}$ and $f_{2}$ representing the control forces applied perpendicular and parallel to the fan axis, respectively. The variable definitions and physical parameters used for the simulation are summarized in Table X and Table XI.
TABLE X: Variable Definitions for DuctedFan
Variable Physical Meaning
$x,y$ Horizontal and vertical displacement (m)
$\theta$ Angular orientation (rad)
$\dot{x},\dot{y}$ Horizontal and vertical velocity (m/s)
$\dot{\theta}$ Angular velocity (rad/s)
$m$ System mass (kg)
$g$ Gravitational acceleration (m/s2)
$r$ Moment arm of the control force (m)
[…]
$u_{1},u_{2}$ Control force inputs (N)
TABLE XI: Physical Parameter Values for DuctedFan
Parameter Value
$g$ 9.81 m/s2
$m$ 8.5 kg
$r$ 0.26 m
[…]
Simulation time step ($dt$) 0.01 s
Control update interval 5 simulation steps
Maximum episode steps 1000
-G3 State and Action Spaces
The system state vector is defined as $\mathbf{x}=[x,y,\theta,\dot{x},\dot{y},\dot{\theta}]^{\top}$. The admissible state space is $\mathcal{X}=[-5,5]^{2}\times[-\pi/2,\pi/2]\times[-5,5]^{3}$, which covers the translational workspace, the pitch angle range, and the respective velocity limits. The control objective is to stabilize the vehicle at the equilibrium state $\mathbf{x}_{g}=[0,0,0,0,0,0]^{\top}$, corresponding to a steady hover at the origin.
The control vector $\mathbf{u}=[u_{1},u_{2}]^{\top}$ consists of two continuous force inputs, with the action space defined as $\mathcal{U}=[-5,5]\times[-5,5]$. These bounds reflect physical actuation limits while ensuring sufficient control authority to counteract gravity and damping. At the start of each episode, the initial state $\mathbf{x}_{0}$ is sampled from a uniform distribution $\mathbf{x}_{0}\sim\text{Uniform}([-0.5,0.5]^{6})$. This initialization provides a localized yet challenging set of starting conditions that require the controller to coordinate multi-axis forces to achieve precise stabilization.
-H Details of Two-link
-H1 System Overview
The Two-link system captures the rigid-body rotational dynamics of a serial two-link planar structure subject to gravitational effects and joint torque inputs. This system serves as a representative benchmark for underactuated or fully actuated mechanical systems, where the nonlinear coupling between links presents a significant control challenge [59]. Key applications of this model include robotic manipulator trajectory tracking, humanoid robot joint coordination, and the study of multi-degree-of-freedom Lagrangian systems [68].
-H2 Dynamics and Physical Parameters
The continuous-time dynamics of the Two-link Planar Robot are derived from Lagrangian mechanics and are expressed by the standard Euler-Lagrange equation:
$\mathbf{M}(\mathbf{q})\ddot{\mathbf{q}}+\mathbf{C}(\mathbf{q},\dot{\mathbf{q}})\dot{\mathbf{q}}+\mathbf{G}(\mathbf{q})=\mathbf{u},$
where $\mathbf{q}=[\theta_{1},\theta_{2}]^{\top}$ denotes the vector of joint angles, $\mathbf{M}(\mathbf{q})$ is the $2\times 2$ symmetric positive-definite mass matrix, $\mathbf{C}(\mathbf{q},\dot{\mathbf{q}})$ is the Coriolis and centrifugal matrix, $\mathbf{G}(\mathbf{q})$ is the gravity vector, and $\mathbf{u}=[u_{1},u_{2}]^{\top}$ is the control vector of joint torques. The explicit forms of these matrices are given by:
$\mathbf{M}(\mathbf{q})=\begin{bmatrix}M_{11}&M_{12}\\ M_{21}&M_{22}\in\mathbb{R}^{2\times 2}\end{bmatrix}$,
$\mathbf{C}(\mathbf{q},\dot{\mathbf{q}})=\begin{bmatrix}-h\dot{\theta}_{2}&-h(\dot{\theta}_{1}+\dot{\theta}_{2})\\ h\dot{\theta}_{1}&0\end{bmatrix}$,
$\mathbf{G}(\mathbf{q})=\begin{bmatrix}-(m_{1}l_{c1}+m_{2}l_{1})g\sin\theta_{1}-m_{2}l_{c2}g\sin(\theta_{1}+\theta_{2})\\ -m_{2}l_{c2}g\sin(\theta_{1}+\theta_{2})\end{bmatrix}$,
where $h=m_{2}l_{1}l_{c2}\sin\theta_{2}$. The individual elements of the mass matrix are defined as:
$M_{11}$ $=I_{1}+I_{2}+m_{1}l_{c1}^{2}+m_{2}(l_{1}^{2}+l_{c2}^{2}+2l_{1}l_{c2}\cos\theta_{2}),$
$M_{12}$ $=M_{21}=I_{2}+m_{2}(l_{c2}^{2}+l_{1}l_{c2}\cos\theta_{2}),$
$M_{22}$ $=I_{2}+m_{2}l_{c2}^{2}.$
The variable definitions and the specific physical parameters used in our simulation are summarized in Table XII and Table XIII, respectively.
TABLE XII: Variable Definitions for Two-link Planar Robot
Variable Physical Meaning
$\theta_{1},\theta_{2}$ Angular orientation of link 1 and link 2 (rad)
$\dot{\theta}_{1},\dot{\theta}_{2}$ Angular velocity of link 1 and link 2 (rad/s)
$\ddot{\theta}_{1},\ddot{\theta}_{2}$ Angular acceleration of link 1 and link 2 (rad/s2)
$l_{1},l_{2}$ Length of link 1 and link 2 (m)
$m_{1},m_{2}$ Mass of link 1 and link 2 (kg)
$l_{c1},l_{c2}$ Distance from joint to center of mass (COM) (m)
[…]
TABLE XIII: Physical Parameter Values for Two-link Planar Robot
Parameter Value
$g$ 9.81 m/s2
$l_{1},l_{2}$ 1.0 m, 1.0 m
$m_{1},m_{2}$ 1.0 kg, 1.0 kg
$l_{c1},l_{c2}$ 0.5 m, 0.5 m
[…]
Simulation time step ($dt$) 0.01 s
Control update interval 5 simulation steps
Maximum episode steps 1000
-H3 State and Action Spaces
The system state vector is defined as $\mathbf{x}=[\theta_{1},\theta_{2},\dot{\theta}_{1},\dot{\theta}_{2}]^{\top}$. To incorporate physical constraints, the state space is bounded by $\mathcal{X}=[-\pi/2,\pi/2]^{2}\times[-20,20]^{2}$, where the joint angles are limited to a $\pm\pi/2$ range and the angular velocities are capped at $\pm 20$ rad/s. The control objective is to achieve stabilization at the equilibrium state $\mathbf{x}_{g}=[0,0,0,0]^{\top}$, representing the system at rest in the downward vertical position.
The control action vector $\mathbf{u}=[u_{1},u_{2}]^{\top}$ consists of continuous joint torques within the action space $\mathcal{U}=[-20,20]\times[-20,20]$. These torque limits are established to prevent excessively high control efforts while providing sufficient authority to manage the gravitational and inertial coupling effects. For each training episode, the initial state $\mathbf{x}_{0}$ is sampled from a uniform distribution $\mathbf{x}_{0}\sim\text{Uniform}([-0.5,0.5]^{4})$, ensuring that the controller begins each run in a neighborhood of the target equilibrium while requiring active correction to maintain stability.
-I Details of SingleCarTracking
-I1 System Overview
The SingleCarTracking environment is designed for high-fidelity ground vehicle trajectory . It utilizes a single-track (bicycle) model that incorporates both lateral and longitudinal dynamics of a rear-wheel-drive vehicle, accounting for nonlinear tire-road friction [69] and steering constraints. This benchmark is crucial for validating control strategies in autonomous driving [60, 70], such as path following and high-precision maneuvers in mobile robotics [10].
-I2 Reference Trajectory and Tracking Error Formulation
In this study, the reference trajectory focuses on the spatial path in the 2D plane. Let $\bm{p}_{t}^{\text{ref}}=[x_{t}^{\text{ref}},y_{t}^{\text{ref}}]^{\top}\in\mathbb{R}^{2}$ denote the target position on the horizontal plane. While the full reference state $\mathbf{x}_{t}^{\text{ref}}$ includes heading and velocity, the core objective is defined by the straight-line trajectory $\bm{p}_{t}^{\text{ref}}=[t,0]^{\top}$ for $t\in[0,50]$. The reference parameters are set to a longitudinal velocity $v_{\text{ref}}=1.0$ m/s, zero acceleration $a_{\text{ref}}=0.0$ m/s2, and a zero heading angle $\psi_{\text{ref}}=0.0$ with zero angular velocity $\omega_{\text{ref}}=0.0$ rad/s. A friction scaling factor $\mu_{\text{scale}}=0.1$ is applied to adjust tire nonlinearity.
The physical state of the vehicle is described by the vector $\mathbf{x}=[x,y,\delta,v,\psi,\dot{\psi},\beta]^{\top}$, representing position, steering angle, longitudinal speed, heading angle, yaw rate, and side-slip angle. For control purposes, we define the tracking error vector as $\mathbf{e}=[s_{xe},s_{ye},\delta_{e},v_{e},\psi_{e},\dot{\psi}_{e},\beta_{e}]^{\top}$, where $s_{xe}=x-x_{\text{ref}}$ and $s_{ye}=y-y_{\text{ref}}$ are the longitudinal and lateral position errors (also denoted as $e_{x}$ and $e_{y}$ respectively in the main text). The admissible error space $\mathcal{E}$ is constrained to $1,1]^{2}\times[-1.066,1.066]\times[-1,1]\times[-\pi/2,\pi/2]^{2}\times[-\pi/3,\pi/3]$, reflecting realistic tracking tolerances and physical actuation limits. The control objective is to maintain the system at the equilibrium error state $\mathbf{e}_{g}=\mathbf{0}$. At the start of each episode, the error is initialized via $\mathbf{e}_{0}\sim\text{Uniform}([-0.5,0.5]^{7})$.
-I3 Dynamics and Physical Parameters
The system dynamics are modeled as an affine nonlinear system $\dot{\mathbf{e}}=\mathbf{f}(\mathbf{e})+\mathbf{g}(\mathbf{e})\mathbf{u}$, where the control input $\mathbf{u}=[\dot{\delta},a_{\text{long}}]^{\top}$ consists of the steering angular velocity and longitudinal acceleration, bounded by $\mathcal{U}=[-5,5]^{2}$. To ensure numerical stability and physical accuracy across different speed regimes, the model transitions between a dynamic tire model and a kinematic bicycle model.
Dynamic Model ($|v|\geq 0.1\mathrm{m/s}$)
In the high-speed regime, the drift vector field is defined as $\mathbf{f}(\mathbf{e})=[v\cos(\psi_{e}+\beta_{e})-v_{\text{ref}}+\omega_{\text{ref}}s_{ye},v\sin(\psi_{e}+\beta_{e})-\omega_{\text{ref}}s_{xe},0,-a_{\text{ref}},\dot{\psi}_{e},f_{6}(\mathbf{e}),f_{7}(\mathbf{e})]^{\top}$. The nonlinear yaw and side-slip terms are:
$f_{6}(\mathbf{e})=$ $\frac{\mu m}{I_{z}(l_{f}+l_{r})}\Bigg[-\frac{l_{f}^{2}C_{Sf}gl_{r}+l_{r}^{2}C_{Sr}gl_{f}}{v}\dot{\psi}$
$+(l_{r}C_{Sr}gl_{f}-l_{f}C_{Sf}gl_{r})\beta_{e}+(l_{f}C_{Sf}gl_{r})\delta_{e}\Bigg],$
$f_{7}(\mathbf{e})=$ $\left[\frac{\mu(C_{Sr}gl_{f}l_{r}-C_{Sf}gl_{r}l_{f})}{v^{2}(l_{f}+l_{r})}-1\right]\dot{\psi}$
$-\frac{\mu(C_{Sr}gl_{f}+C_{Sf}gl_{r})}{v(l_{f}+l_{r})}\beta_{e}+\frac{\mu C_{Sf}gl_{r}}{v(l_{f}+l_{r})}\delta_{e}.$
where
$C_{Sf}=C_{Sr}=-\frac{\text{tire\_p\_ky1}}{\text{tire\_p\_dy1}}$
The control gain matrix is $\mathbf{g}(\mathbf{e})=[\mathbf{0}_{2\times 2},[1,0],[0,1],\mathbf{0},[0,\mathbf{g}_{6,2}(\mathbf{e})],[0,\mathbf{g}_{7,2}(\mathbf{e})]]^{\top}$. The terms $g_{6,2}$ and $g_{7,2}$ account for CG height $h_{s}$ and vertical load shifts:
$g_{6,2}(\mathbf{e})=$ $\frac{\mu m}{I_{z}(l_{f}+l_{r})}\Bigg[-\frac{-l_{f}^{2}C_{Sf}h_{s}+l_{r}^{2}C_{Sr}h_{s}}{v}\dot{\psi}$
$+(l_{r}C_{Sr}h_{s}+l_{f}C_{Sf}h_{s})\beta_{e}-(l_{f}C_{Sf}h_{s})\delta_{e}\Bigg],$
$g_{7,2}(\mathbf{e})=$ $\frac{\mu(C_{Sr}h_{s}l_{r}+C_{Sf}h_{s}l_{f})}{v^{2}(l_{f}+l_{r})}\dot{\psi}$
$-\frac{\mu(C_{Sr}h_{s}-C_{Sf}h_{s})}{v(l_{f}+l_{r})}\beta_{e}-\frac{\mu C_{Sf}h_{s}}{v(l_{f}+l_{r})}\delta_{e}.$
Kinematic Model ($|v|<0.1\mathrm{m/s}$)
To avoid singularities at low speeds, the model simplifies to a kinematic representation where side-slip is negligible. Let $L=l_{f}+l_{r}$ denote the wheelbase. The drift field $\mathbf{f}(\mathbf{e})$ and input matrix $\mathbf{g}(\mathbf{e})$ are expressed as:
$\mathbf{f}(\mathbf{e})=\Big[$ $v\cos(\psi_{e}+\beta_{e})-v_{\text{ref}}+\omega_{\text{ref}}s_{ye},$
$v\sin(\psi_{e}+\beta_{e})-\omega_{\text{ref}}s_{xe},$
$0,\qquad-a_{\text{ref}},\qquad\frac{v\cos(\beta_{e})}{L}\tan(\delta_{e})-\omega_{\text{ref}},\qquad 0,\qquad 0\Big]^{\top},$
$\mathbf{g}(\mathbf{e})=\Big[$ $\mathbf{0}_{2\times 2},\qquad[1,0],\qquad[0,1],\qquad\mathbf{0},$
$\frac{\cos(\beta_{e})\tan(\delta_{e})}{L},\qquad\beta_{e}^{\prime}\Big]^{\top},$
where the side-slip rate auxiliary term is
$\beta_{e}^{\prime}=\frac{l_{r}/(L\cos^{2}(\delta_{e}))}{1+\left(\tan(\delta_{e})\frac{l_{r}}{L}\right)^{2}}.$
The system parameters are detailed in Tables XIV and XV.
TABLE XIV: Variable Definitions for SingleCarTracking
Category Variable Physical Meaning
True State $\mathbf{x}\in\mathbb{R}^{7}$ $\mathbf{x}=[x,y,\delta,v,\psi,\dot{\psi},\beta]^{\top}$: Position ($x,y$), steering angle, speed, heading, yaw rate, side-slip.
Tracking Error $\mathbf{e}\in\mathbb{R}^{7}$ $\mathbf{e}=[s_{xe},s_{ye},\delta_{e},v_{e},\psi_{e},\dot{\psi}_{e},\beta_{e}]^{\top}$: Tracking errors relative to reference trajectory.
Control $\mathbf{u}\in\mathbb{R}^{2}$ $\mathbf{u}=[\dot{\delta},a_{\text{long}}]^{\top}$: Steering rate (rad/s) and longitudinal acceleration (m/s2).
Parameters $m,I_{z},l_{f},l_{r},h_{s}$ Mass, yaw inertia, distance to front/rear axles, and CG height.
$\mu,K_{\text{tire}}$ Tire-road friction coefficient and tire stiffness parameters.
TABLE XV: Physical Parameter Values for SingleCarTracking
Parameter Value Parameter Value
[…]
$l_{f},l_{r}$ 1.155 m, 1.423 m $h_{s}$ 0.614 m
$\delta_{\text{max}}$ $\pm 1.066$ rad $a_{\text{long, max}}$ $\pm 5$ m/s2
tire_p_dy1 1.0489 tire_p_ky1 -21.92
dtdt 0.01 s Horizon $T$ 1000
-J Details of QuadrotorTracking
-J1 System Overview
The QuadrotorTracking environment evaluates the 6-degree-of-freedom (6-DoF) rigid-body dynamics of a quadrotor UAV [71]. This benchmark incorporates complex rotor thrust-torque coupling and attitude kinematics defined on the special Euclidean group $\text{SE}(3)$ [54]. It serves as a standard testbed for geometric control and reinforcement learning algorithms in high-agility scenarios[72], such as autonomous aerial inspection and precise path following [73, 74].
-J2 Reference Trajectory and Tracking Error Formulation
The control objective is to track a time-varying reference trajectory that specifies position, velocity, and heading. Let $\bm{p}_{t}^{\text{ref}}=[x_{t}^{\text{ref}},y_{t}^{\text{ref}},z_{t}^{\text{ref}}]^{\top}$ denote the desired 3D position at time step $t$. The benchmark utilizes a dynamic helical trajectory defined by:
$\bm{p}_{t}^{\text{ref}}$ $=[0.4t,0.4\sin(t),0.6\cos(t)]^{\top},$
$\bm{v}_{t}^{\text{ref}}$ $=[0.4,0.4\cos(t),-0.6\sin(t)]^{\top},$
where $\bm{v}_{t}^{\text{ref}}$ is the reference velocity. The desired heading is determined by a rotating unit vector $\bm{b}_{1,t}^{\text{ref}}=[\cos(t),\sin(t),0]^{\top}$.
The system state is $\mathbf{x}=[\mathbf{x}_{p}^{\top},\mathbf{v}^{\top},\mathbf{R},\bm{\Omega}^{\top}]^{\top}$, where $\mathbf{x}_{p}\in\mathbb{R}^{3}$ is the position, $\mathbf{v}\in\mathbb{R}^{3}$ is the translational velocity, $\mathbf{R}\in\text{SO}(3)$ is the rotation matrix, and $\bm{\Omega}\in\mathbb{R}^{3}$ is the body-frame angular velocity. Although $\mathbf{R}$ is a $3\times 3$ matrix, it fundamentally represents three degrees of freedom (3-DoF). Thus for the sake of notational consistency, with a slight abuse of notation, we treat $\mathbf{R}$ as a three-dimensional vector, thereby concatenating it with $\mathbf{x}_{p}^{\top}$, $\mathbf{v}^{\top}$, and $\bm{\Omega}^{\top}$ to constitute the system state $\mathbf{x}$.
The tracking error $\mathbf{e}\in\mathbb{R}^{12}$ is defined by four components: position error $\mathbf{e}_{p}=\mathbf{x}_{p}-\bm{p}_{t}^{\text{ref}}$, velocity error $\mathbf{e}_{v}=\mathbf{v}-\bm{v}_{t}^{\text{ref}}$, attitude error $\mathbf{e}_{R}$, and angular velocity error $\mathbf{e}_{\Omega}$. Following [54], the attitude error is computed via the skew-symmetric matrix $\mathbf{S}=\mathbf{R}_{t}^{\text{ref}\top}\mathbf{R}-\mathbf{R}^{\top}\mathbf{R}_{t}^{\text{ref}}$, where $\mathbf{e}_{R}=0.5[\mathbf{S}_{2,1},\mathbf{S}_{0,2},\mathbf{S}_{1,0}]^{\top}$, and $\mathbf{S}_{i,j}$ denotes the $(i,j)$-th element of the $\mathbf{S}$ matrix. The angular velocity error is $\mathbf{e}_{\Omega}=\bm{\Omega}-\mathbf{R}^{\top}\mathbf{R}_{t}^{\text{ref}}\bm{\Omega}_{t}^{\text{ref}}$. The composite error vector $\mathbf{e}=[\mathbf{e}_{p}^{\top},\mathbf{e}_{v}^{\top},\mathbf{e}_{R}^{\top},\mathbf{e}_{\Omega}^{\top}]^{\top}$ is used as the input for the controller. At the start of each episode, the initial state $\mathbf{x}_{0}$ is uniformly randomly sampled to calculate $\mathbf{e}_{0}$, specifically
$\left\{\begin{array}[]{l}\mathbf{x}_{p}^{\top}\sim\text{Uniform}\left([-0.01,0.01]^{3}\right),\\ \mathbf{v}^{\top}\sim\text{Uniform}\left([-0.01,0.01]^{3}\right),\\ \text{vec}(\mathbf{R}-\mathbf{I}_{3})\sim\text{Uniform}\left([-0.01,0.01]^{9}\right),\mathbf{R}\text{ is orthogonal},\\ \bm{\Omega}^{\top}\sim\text{Uniform}\left([-0.01,0.01]^{3}\right).\end{array}\right$.
-J3 Dynamics and Physical Parameters
The continuous-time dynamics follow the Newton-Euler equations for rigid bodies:
$\begin{cases}\dot{\mathbf{x}}_{p}=\mathbf{v},&\dot{\mathbf{R}}=\mathbf{R}[\bm{\Omega}]_{\times},\\ \dot{\mathbf{v}}=-\mathbf{g}+\frac{F}{m}\mathbf{b}_{3},&\dot{\bm{\Omega}}=\mathbf{J}^{-1}(\mathbf{M}-\bm{\Omega}\times\mathbf{J}\bm{\Omega}),\end{cases}$
[…]
$\begin{bmatrix}F\\ M_{x}\\ M_{y}\\ M_{z}\end{bmatrix}=\begin{bmatrix}1&1&1&1\\ 0&-d&0&d\\ d&0&-d&0\\ -c_{tf}&c_{tf}&-c_{tf}&c_{tf}\end{bmatrix}\begin{bmatrix}f_{1}\\ f_{2}\\ f_{3}\\ f_{4}\end{bmatrix}$
where $d$ is the moment arm and $c_{tf}$ is the thrust-torque coefficient. Detailed parameter values are provided in Tables XVI and XVII.
TABLE XVI: Variable Definitions for QuadrotorTracking
Category Variable Physical Meaning
True State $\mathbf{x}=[\mathbf{x}_{p}^{\top},\mathbf{v}^{\top},\mathbf{R},\bm{\Omega}^{\top}]^{\top}$ 3D position, velocity, rotation matrix elements, and angular velocity.
Tracking Error $\mathbf{e}=[\mathbf{e}_{p}^{\top},\mathbf{e}_{v}^{\top},\mathbf{e}_{R}^{\top},\mathbf{e}_{\Omega}^{\top}]^{\top}$ Geometric tracking errors in position, velocity, attitude, and angular rates.
[…]
Parameters $m,\mathbf{J},d,c_{tf}$ Mass, inertia matrix, rotor distance, and torque coefficient.
TABLE XVII: Physical Parameter Values for QuadrotorTracking
Parameter Value Parameter Value
$m$ 4.34 kg $\mathbf{g}$ $0,0,9.8]^{\top}$
$\mathbf{J}$ $\text{diag}(0.08,0.09,0.14)$ $d$ 0.315 m
[…]
dtdt 0.01 s $T$ 1000

### v2

@@
 Abstract
-Achieving provable stability guarantees in model-free reinforcement learning (RL) for control systems remains a fundamental challenge, particularly in balancing exploration efficiency with rigorous safety requirements.
-This article introduces MSACL, a novel framework that seamlessly integrates exponential stability theory with maximum entropy reinforcement learning through multi-step Lyapunov certificate learning.
-Unlike conventional approaches that rely on complex reward engineering or single-step constraints, MSACL leverages off-policy multi-step trajectory data to learn Lyapunov certificates that strictly satisfy theoretical stability conditions.
-By introducing the concept of Exponential Stability Labels (ESL) and a $\lambda$-weighted aggregation mechanism, the framework effectively balances the bias-variance trade-off inherent in multi-step learning.
-Policy optimization is subsequently guided by a stability-aware advantage function, which ensures that the learned policy $\pi$ promotes the rapid descent of the Lyapunov certificate.
-We rigorously evaluate the proposed method across six benchmark environments, encompassing four stabilization tasks and two high-dimensional nonlinear tracking scenarios.
-The experimental results demonstrate consistent superiority over baseline RL algorithms and state-of-the-art Lyapunov-based RL algorithms.
-Remarkably, MSACL achieves both exponential stability and rapid convergence under simple, intuitive reward structures, while exhibiting significant robustness to environmental uncertainties and exceptional generalization to unseen reference trajectories.
-Our sensitivity analysis reveals that the multi-step horizon $n$ is a critical factor in stability learning, with $n=20$ established as a robust default across diverse dynamical systems.
-By establishing a rigorous connection between Lyapunov stability theory and off-policy actor-critic frameworks, MSACL provides a foundational approach for developing verifiably safe learning-based controllers in safety-critical applications.
-To facilitate reproducibility and further research, the source code and benchmark environments of MSACL will be made publicly available soon.
+For safety-critical applications, model-free reinforcement learning (RL) faces numerous challenges, particularly the difficulty of establishing verifiable stability guarantees while maintaining high exploration efficiency.
+To address these challenges, we present Multi-Step Actor-Critic Learning with Lyapunov Certificates (MSACL), a novel approach that seamlessly integrates exponential stability with maximum entropy reinforcement learning (MERL).
+In contrast to existing methods that rely on complex reward engineering and single-step constraints, MSACL utilizes intuitive rewards and multi-step data for actor-critic learning.
+Specifically, we first introduce Exponential Stability Labels (ESLs) to categorize samples and propose a $\lambda$-weighted aggregation mechanism to learn Lyapunov certificates.
+Leveraging these certificates, we then develop a stability-aware advantage function to guide policy optimization, thereby ensuring rapid Lyapunov descent and robust state convergence.
+We evaluate MSACL across six benchmarks, comprising four stabilization and two high-dimensional tracking tasks.
+Experimental results demonstrate its consistent superiority over both standard RL baselines and state-of-the-art Lyapunov-based RL algorithms.
+Beyond rapid convergence, MSACL exhibits significant robustness against environmental uncertainties and remarkable generalization to unseen reference signals.
+The source code and benchmarking environments are available at https://github.com/YuanZhe-Xing/MSACL.
 I Introduction
-Reinforcement learning (RL) has emerged as a powerful paradigm for controlling complex, nonlinear dynamical systems without requiring analytical models [1].
-Despite its empirical success in high-dimensional tasks [2, 3], the widespread deployment of RL in safety-critical applications, including autonomous driving and human-robot interaction, remains constrained by the lack of formal reliability guarantees [4, 5].
-In these domains, stability is not merely a desirable metric but an imperative constraint; failure to maintain the system within a safe operating region can lead to catastrophic consequences.
-While approaches like constrained Markov Decision Processes (cMDPs) [6, 7] and Hamilton-Jacobi reachability analysis [8, 9] offer partial solutions, certificate-based methods have established themselves as the gold standard for verifying learning agents.
-By leveraging rigorous control-theoretic tools, specifically Lyapunov functions for asymptotic stability [10, 11] and control barrier functions (CBFs) for set invariance [12, 13], these methods provide concise, data-driven proofs that guarantee the system will converge to an equilibrium or remain strictly within safety boundaries, offering a level of assurance that standard reward-driven learning cannot match.
-The synthesis of stability and safety certificates has undergone a paradigm shift from computationally expensive analytical methods, such as sum-of-squares (SOS) optimization with its inherent restriction to polynomial dynamics [14], to scalable neural approaches.
-Around 2018, foundational work on neural certificates emerged, marking a turning point in bridging black-box learning with verifiable control.
-Early efforts by Berkenkamp et al.
-[15] employed Gaussian Processes (GP) to learn Lyapunov functions for safe exploration, while Richards et al.
-[16] and Chow et al.
-[17] developed theoretical frameworks for synthesizing Lyapunov functions to estimate regions of attraction.
-These studies established that Lyapunov certificates as well as barrier-based counterparts provide a rigorous mathematical foundation for analyzing the stability and safety of nonlinear systems.
-Building on this verification-centric perspective, recent research has advanced toward the joint synthesis of policies and certificates: Chang et al.
-[18], for instance, showed that a neural certificate can act as a “critic” or supervision signal during training, directly guiding policy learning to satisfy formal stability conditions.
-This evolution underscores the dual role of learned certificates, which not only enable post-hoc safety verification but also actively shape the learning process to ensure stability and safety by design, offering guarantees that standard trial-and-error reinforcement learning cannot provide.
-To enable effective learning in complex continuous spaces, we build our approach upon the framework of Maximum Entropy Reinforcement Learning (MERL).
-Algorithms such as Soft Actor-Critic (SAC) [19] and its distributional variants like DSAC [20, 21] maximize not just the expected return but also the entropy of the policy.
-This entropy regularization encourages robust exploration and prevents the policy from converging prematurely to suboptimal deterministic behaviors.
-Given its superior sample efficiency and robustness to disturbances, the maximum entropy framework serves as an ideal substrate for developing advanced, stability-aware control algorithms.
-However, standard model-free RL algorithms, including DDPG [22], TRPO [23], PPO [24], TD3 [25], D4PG [26], SAC, and DSAC, face significant limitations when applied to stabilization or tracking tasks.
-First, they lack intrinsic stability guarantees; since policies trained via these methods optimize solely for scalar reward maximization, stability is essentially an incidental byproduct that is achieved only when the agent implicitly discovers stable behaviors during the exploration process, rather than being structurally enforced.
-Second, the efficacy of these algorithms relies heavily on exhaustive reward engineering, requiring practitioners to laboriously calibrate coefficients for state deviation, control effort, and auxiliary terms to elicit stable behavior.
-Our proposed framework, MSACL (Multi-Step Actor-Critic Learning with Lyapunov Certificates), directly addresses these shortcomings.
-By adopting a formulation rooted in quadratic performance functionals such as LQR-like costs, MSACL eliminates the need for heuristic reward shaping.
-Most importantly, it guarantees stability by design, ensuring that the synthesized controller is not only high-performing but also mathematically certifiable.
-While Lyapunov-based RL algorithms have been proposed to mitigate the two issues above, existing methods exhibit distinct shortcomings.
-The Policy Optimization with Lyapunov Certificates (POLYC) [27] algorithm offers theoretical guarantees but relies on on-policy learning, which results in prohibitive sample complexity for physical systems.
-Conversely, the Lyapunov-based Actor-Critic (LAC) [28] algorithm introduces an off-policy approach but fundamentally misaligns the objective: it minimizes a stability risk rather than a comprehensive control cost.
-This often leads to aggressive controllers that stabilize the system at the cost of excessive control effort, rendering them impractical for real-world actuators.
-Moreover, POLYC and LAC typically certify only asymptotic stability, a relatively weaker stability notion that ensures convergence but lacks the guaranteed exponential rate of decay.
-In practice, this guarantees eventual convergence but fails to impose a minimum convergence rate, often resulting in sluggish transient responses unacceptable for high-performance tasks.
-To overcome these limitations, MSACL is built upon an off-policy framework [29, 30], significantly improving data efficiency by leveraging historical agent-environment interaction data.
-Unlike LAC, MSACL optimizes a comprehensive objective covering both state regulation and control effort while enforcing exponential stability, which guarantees faster convergence rates and rigorous stability margins.
-Another critical innovation of our approach is the utilization of multi-step data.
-Existing RL and Lyapunov-based methods typically rely on single-step transitions to enforce learning constraints.
-This pointwise reliance is susceptible to significant discretization errors when approximating the continuous-time Lie derivative of the Lyapunov function, particularly in the presence of local noise.
-To mitigate this and capture long-horizon dynamics, we introduce a multi-step learning mechanism.
-By training on $n$-step trajectories, we effectively bridge the gap between temporal difference learning and rigorous stability verification.
-We integrate this multi-step data with exponential stability theory to form a robust constraint mechanism.
-This allows the agent to learn a Lyapunov function that certifies exponential convergence over trajectory segments, rather than just instantaneous descent, thereby reconciling rigorous stability requirements with the need for flexible exploration.
-In summary, this article presents a comprehensive RL framework for model-free exponentially stabilizing control. The main contributions are summarized as follows:
+Reinforcement learning (RL) has emerged as a transformative paradigm for the control of complex, nonlinear dynamical systems [1, 2], demonstrating remarkable empirical success across high-dimensional tasks without requiring explicit analytical models [3, 4].
+Despite these advancements, the deployment of RL in safety-critical applications, such as autonomous driving and human-robot interaction, remains fundamentally constrained by a lack of formal reliability guarantees [5, 6].
+While alternative frameworks like constrained Markov Decision Processes (cMDPs) [7, 8] and Hamilton-Jacobi reachability analysis [9, 10] offer potential solutions, certificate-based methods constitute a fundamental framework for verifying and stabilizing learning agents.
+By leveraging rigorous control-theoretic tools, such as Lyapunov certificates for establishing asymptotic stability [11, 12, 14, 13] and barrier certificates for ensuring set invariance [15, 16], these methods provide concise, data-driven proofs of system behavior.
+The synthesis of stability certificates has undergone a paradigm shift from computationally intensive analytical methods toward scalable neural architectures.
+Traditional approaches, such as sum-of-squares (SOS) optimization [17, 18, 19] and algebraic synthesis[20], are often restricted to polynomial dynamics.
+To bridge the gap between black-box learning and verifiable control, the field has evolved from probabilistic methods, notably the use of Gaussian Processes for safe exploration [21], to formal frameworks for synthesizing neural certificates that estimate regions of attraction (ROA) [22, 23].
+Building on this verification-centric perspective, recent research has advanced toward the joint synthesis of neural certificates and policies.
+For instance, Chang et al.
+[24, 25] and Quan et al.
+[26, 27, 28] demonstrated that neural certificates can function as critics during training, directly guiding policy optimization to satisfy formal stability constraints.
+In contrast to the stability-by-design paradigms presented above, standard model-free RL algorithms, such as DDPG [29], PPO [31], TD3 [32], SAC [33], and SAC’s distributional variant DSAC [34, 35], lack intrinsic stability guarantees, thus facing fundamental limitations in precision control tasks.
+These methods treat stability as an incidental byproduct of reward maximization rather than a structural property.
+Consequently, employing general RL algorithms for safety-critical tasks necessitates extensive reward engineering to achieve stable behaviors [1, 2].
+To bridge this gap, Lyapunov-based RL algorithms have been developed to integrate stability constraints into the learning process.
+For instance, Policy Optimization with Lyapunov Certificates (POLYC) [25] incorporates Lyapunov certificates within the PPO framework, while Lyapunov-based Actor-Critic (LAC) [36] introduces an off-policy approach for learning stability-constrained critics.
+By augmenting standard RL architectures with learned Lyapunov certificates, these methods provide structural stability guarantees for nonlinear control systems.
+However, several critical limitations persist.
+First, these approaches suffer from low sample efficiency under on-policy training paradigms.
+Second, they often fail to account for control costs during policy optimization.
+More critically, most existing methods focus solely on asymptotic stability, which lacks the rapid stabilization characteristics necessary for high-performance response in dynamic control scenarios.
+To overcome these limitations, we propose Multi-Step Actor-Critic Learning with Lyapunov Certificates (MSACL), a Lyapunov-based off-policy RL algorithm that enhances sample efficiency [37, 38] while optimizing a comprehensive objective that balances state regulation and control cost.
+By integrating exponential stability constraints into the maximum entropy reinforcement learning (MERL) architecture, MSACL ensures rapid stabilization and enhanced robustness.
+A key feature of our approach is the multi-step learning mechanism, which captures long-horizon dynamics to establish connections between temporal difference learning and formal stability verification.
+By mitigating discretization errors and stochastic noise associated with single-step transitions, this mechanism enables the synthesis of policies that are both high-performing and mathematically certifiable.
+In summary, this article presents a Lyapunov-based RL approach for model-free exponentially stabilizing control. The main contributions are summarized as follows:
   1. (i)
-    Exponential Stability Guarantee via Multi-Step Learning: We introduce the MSACL framework, which integrates exponential stability conditions into Lyapunov certificate learning using a multi-step data utilization mechanism.
-By employing Exponential Stability Labels (ESL), our approach enforces that the system state converges to the equilibrium exponentially, providing stronger performance guarantees than conventional asymptotic convergence methods.
+    Exponential Stability via Multi-Step Learning.
+We propose the MSACL algorithm, which integrates exponential stability with Lyapunov certificate learning via a multi-step mechanism.
+This algorithm enforces exponential convergence rates, providing stability-guaranteed performance for complex nonlinear control, surpassing conventional asymptotic methods.
   2. (ii)
-    Sample-Efficient Off-Policy Framework: Developed within the maximum entropy RL paradigm, our framework enables efficient off-policy learning.
-By incorporating trajectory-wise importance sampling to correct for distribution shifts, MSACL allows for effective reuse of historical interaction data, significantly reducing sample complexity compared to existing on-policy Lyapunov-based RL algorithms.
+    Sample-Efficient Off-Policy Framework.
+We develop a sample-efficient off-policy framework that leverages the maximum entropy paradigm and importance sampling.
+This framework effectively manages distribution shifts and interaction data reuse, thus enhancing training stability compared to existing on-policy algorithms.
   3. (iii)
-    Comprehensive Benchmarking System: We developed a diverse suite of six nonlinear dynamical environments, spanning four stabilization tasks (e.g., VanderPol) and two high-dimensional tracking tasks (e.g., Quadrotor Trajectory Tracking), to rigorously evaluate and provide a standard for Lyapunov-based RL algorithms.
-  4. (iv)
-    Robustness and Generalization Verification: Our methodology demonstrates significant resilience to stochastic process noise and parametric uncertainties.
-Furthermore, in tracking scenarios, the learned neural policies exhibit robust generalization capabilities, maintaining high-precision performance even when navigating geometrically distinct, unseen reference trajectories.
-  5. (v)
-    Bias-Variance Trade-off Analysis: We provide a systematic analysis of the multi-step horizon parameter $n$ and its impact on stability learning.
-Through the introduction of $\lambda$-weighted aggregation, we effectively manage the trade-off between estimation bias (associated with small $n$) and sampling variance (associated with large $n$), establishing as a robust heuristic for diverse control applications.
+    Systematic Evaluation and Robustness Analysis.
+We establish an evaluation platform across six benchmarks comprising stabilization and high-dimensional tracking tasks.
+Experiment results demonstrate superior performance, robustness and generalization of MSACL, while sensitivity analyses elucidate the multi-step bias-variance trade-off to guide diverse control tasks.
 The remainder of this article is organized as follows.
-Section II formulates the stabilization and tracking problems and establishes the theoretical foundations, including the criteria for exponential stability and the maximum entropy reinforcement learning framework.
-Section III details the methodology for learning the Lyapunov certificate, specifically describing the multi-step data collection strategy and the loss functions designed to enforce formal stability conditions.
-Section IV presents the learning process for the neural control policy, deriving the optimization objectives for the soft Q-networks, the policy network, and the temperature coefficient.
-Section V reports the experimental results, providing a comprehensive comparison with baseline algorithms across six benchmark environments, followed by an evaluation of robustness, generalization capabilities, and parameter sensitivity.
-Finally, Section VI provides a discussion on the underlying design philosophy and concludes the article.
+Section II formulates the control problems and establishes necessary theoretical foundations.
+Section III describes the Lyapunov certificate learning methodology.
+Section IV presents the optimization process within the actor-critic framework.
+Section V provides comprehensive experimental evaluations across six benchmarks.
+Finally, Section VI discusses the design philosophy and concludes this article.
 II Preliminaries
-In this section, we establish the theoretical foundations for the proposed MSACL framework.
-We first define the stabilization and tracking problems for unknown discrete-time systems.
-Subsequently, the concept of exponential stability is introduced along with a Lyapunov-based certificate lemma, providing the criteria for convergence guarantees.
-Finally, we present the maximum entropy reinforcement learning framework, bridging Lyapunov-based control with model-free actor-critic optimization.
+This section establishes the theoretical foundations for MSACL.
+We first define the stabilization and tracking problems for model-free discrete-time systems, then introduce Lyapunov-based exponential stability criteria, and finally review the MERL framework.
 II-A Problem Statement
-In this article, we study the stabilization and tracking problems for general discrete-time deterministic control systems:
+In this article, we study the stabilization problem for general discrete-time deterministic control systems:
 $\mathbf{x}_{t+1}=f(\mathbf{x}_{t},\mathbf{u}_{t}),$ (1)
-where $\mathbf{x}_{t}\in\mathcal{X}\subseteq\mathbb{R}^{n}$ denotes the state at time $t$, $\mathbf{u}_{t}\in\mathcal{U}\subseteq\mathbb{R}^{m}$ represents the control input, $\mathcal{X}$ and $\mathcal{U}$ denote the sets of admissible states and control inputs, respectively, and $f:\mathcal{X}\times\mathcal{U}\to\mathcal{X}$ is the flow map.
-We assume that $f$ is unknown in our setting, thus focusing on model-free control.
-It is worth noting that the discrete-time framework presented in this article can be similarly extended to the continuous-time case; interested readers may refer to [10] and [27] for further details.
-In the context of control systems, our objective is to find a state feedback policy $\pi:\mathcal{X}\to\mathcal{U}$ such that the control input $\mathbf{u}_{t}=\pi(\mathbf{x}_{t})$ renders the closed-loop system $\mathbf{x}_{t+1}=f(\mathbf{x}_{t},\pi(\mathbf{x}_{t}))$ to satisfy desired properties (e.g., stability, safety).
-For clarity, we consider the scenario where the state $\mathbf{x}_{t}$ is fully observable, and the policy $\pi$ is a time-invariant mapping.
-Specifically, we focus on tasks including
-  • Stabilization: The objective is to find a policy $\pi$ such that the state ultimately converges to a target state $\mathbf{x}_{g}\in\mathcal{X}$.
-Without loss of generality, we assume $\mathbf{x}_{g}$ to be the origin, i.e., $\mathbf{x}_{g}=\mathbf{0}$.
-  • Tracking: The goal is to find a policy $\pi$ such that the state $\mathbf{x}_{t}$ eventually converges to a reference signal $\mathbf{x}_{t}^{\text{ref}}$ (which may represent a desired trajectory, velocity, or attitude).
-We define the tracking error as $\mathbf{e}_{t}=\mathbf{x}_{t}-\mathbf{x}_{t}^{\text{ref}}$, thereby transforming the tracking task into a stabilization task where $\mathbf{e}_{t}$ must converge to zero.
-In brief, the core problem addressed in this article is formulated as follows:
+where $\mathbf{x}_{t}\in\mathcal{X}\subseteq\mathbb{R}^{d}$ denotes the state at time $t$, $\mathbf{u}_{t}\in\mathcal{U}\subseteq\mathbb{R}^{m}$ denotes the control input, $\mathcal{X}$ and $\mathcal{U}$ represent the sets of admissible states and control inputs, respectively, and $f:\mathcal{X}\times\mathcal{U}\to\mathcal{X}$ is the flow map.
+Given that $f$ is unknown, we adopt a model-free control paradigm.
+Let $\mathcal{X}_{0}\subseteq\mathcal{X}$ denote the set of initial states.
+For any $\mathbf{x}_{0}\in\mathcal{X}_{0}$, let $\mathbf{x}(t,\mathbf{x}_{0})$ denote the state at time $t$ along the trajectory starting from $\mathbf{x}_{0}$.
+For notational simplicity, $\mathbf{x}_{t}$ is used to denote $\mathbf{x}(t,\mathbf{x}_{0})$ whenever the context is unambiguous.
+It is worth noting that the discrete-time case presented in this article can be similarly extended to the continuous-time case; interested readers may refer to [11] and [25] for details.
+In the context of control systems, our objective is to find a state feedback policy $\pi:\mathcal{X}\to\mathcal{U}$ such that the control input $\mathbf{u}_{t}=\pi(\mathbf{x}_{t})$ renders the closed-loop system $\mathbf{x}_{t+1}=f(\mathbf{x}_{t},\pi(\mathbf{x}_{t}))$ to satisfy exponential stability.
+For clarity, we consider the scenario where the state $\mathbf{x}_{t}$ is fully observable, and the policy $\pi$ is stationary, i.e., it depends only on the state and not explicitly on time.
+Specifically, the problem studied in this article is formulated as:
 Problem 1
-For the control system (1), find a policy $\pi$ such that: (i) in stabilization tasks, $\mathbf{x}_{t}\to\mathbf{0}$; (ii) in tracking tasks, $\mathbf{e}_{t}\to\mathbf{0}$.
-Moreover, the policy should be optimized to achieve the fastest possible convergence rate (e.g., exponential stability).
+(Exponential Stabilization) For the control system (1), design a stationary policy $\pi$ such that for any initial state $\mathbf{x}_{0}\in\mathcal{X}_{0}$, the resulting closed-loop trajectory $\mathbf{x}(t,\mathbf{x}_{0})$ exponentially converges to the equilibrium state $\mathbf{x}_{g}$.
+Without loss of generality, let $\mathbf{x}_{g}$ denote the origin, i.e., $\mathbf{x}_{g}=\mathbf{0}$.
+Furthermore, the tracking problem is inherently encompassed within the stabilization framework, as tracking a time-varying reference signal $\mathbf{x}_{t}^{\text{ref}}$ reduces to stabilizing the error $\mathbf{e}_{t}=\mathbf{x}_{t}-\mathbf{x}_{t}^{\text{ref}}$ at zero.
 II-B Exponential Stability and Lyapunov Certificates
 Definition 1
-For the control system (1), the equilibrium point $\mathbf{x}_{g}=\mathbf{0}$ is said to be exponentially stable if there exist constants $C>0$ and $0<\eta<1$ such that
-$\|\mathbf{x}_{t}-\mathbf{x}_{g}\|\leq C\eta^{t}\|\mathbf{x}_{0}-\mathbf{x}_{g}\|,\quad\forall\mathbf{x}_{0}\in\mathcal{X}_{0},$ (2)
-where $\mathbf{x}_{0}\in\mathcal{X}_{0}\subseteq\mathcal{X}$ denotes the initial state and $t\in\mathbb{Z}_{\geq 0}$.
-To guarantee that a closed-loop system is exponentially stable with respect to $\mathbf{x}_{g}$, we introduce the following lemma:
+For the control system (1), the equilibrium state $\mathbf{x}_{g}$ is said to be exponentially stable if there exist constants $C>0$ and $0<\eta<1$ such that
+$\|\mathbf{x}_{t}-\mathbf{x}_{g}\|\leq C\eta^{t}\|\mathbf{x}_{0}-\mathbf{x}_{g}\|,\quad\forall\mathbf{x}_{0}\in\mathcal{X}_{0},\forall t\geq 0.$ (2)
+To guarantee that the closed-loop system is exponentially stable with respect to $\mathbf{x}_{g}$, we introduce the following lemma:
 Lemma 1
@@
 for all $t\in\mathbb{Z}_{\geq 0}$, where $\mathbf{x}_{t+1}=f(\mathbf{x}_{t},\mathbf{u}_{t})$, then the origin $\mathbf{x}_{g}=\mathbf{0}$ is exponentially stable.
-The function $V$ is referred to as an exponentially stabilizing control Lyapunov function (ESCLF) [10, 31].
-Proof: From the condition $V(\mathbf{x}_{t+1})\leq(1-\alpha_{3})V(\mathbf{x}_{t})$, we can iteratively derive the multi-step relationship:
-$V(\mathbf{x}_{t+1})$ $\leq(1-\alpha_{3})V(\mathbf{x}_{t}),$ (4)
-$V(\mathbf{x}_{t+2})$ $\leq(1-\alpha_{3})V(\mathbf{x}_{t+1})\leq(1-\alpha_{3})^{2}V(\mathbf{x}_{t}),$
-$\vdots$
-$V(\mathbf{x}_{t+n})$ $\leq(1-\alpha_{3})V(\mathbf{x}_{t+n-1})\leq\dots\leq(1-\alpha_{3})^{n}V(\mathbf{x}_{t}).$
-By incorporating the lower and upper quadratic bounds $\alpha_{1}\|\mathbf{x}_{t}\|^{2}\leq V(\mathbf{x}_{t})\leq\alpha_{2}\|\mathbf{x}_{t}\|^{2}$, it follows that:
+The function $V$ is referred to as an exponentially stabilizing control Lyapunov function.
+Lemma 1 is a classic result in control theory, while its continuous-time counterpart is discussed in [39], we provide a concise derivation of the discrete-time version to elucidate its underlying mechanism.
+Specifically, starting from the descent condition $V(\mathbf{x}_{t+1})\leq(1-\alpha_{3})V(\mathbf{x}_{t})$ in (3), the multi-step relationship can be established iteratively as:
+$V(\mathbf{x}_{t+n})\leq(1-\alpha_{3})V(\mathbf{x}_{t+n-1})\leq\dots\leq(1-\alpha_{3})^{n}V(\mathbf{x}_{t}).$ (4)
+By incorporating the lower and upper quadratic bounds, i.e., $\alpha_{1}\|\mathbf{x}_{t+n}\|^{2}\leq V(\mathbf{x}_{t+n})$ and $V(\mathbf{x}_{t})\leq\alpha_{2}\|\mathbf{x}_{t}\|^{2}$, the relationship between state norms across multiple steps follows as:
 $\|\mathbf{x}_{t+n}\|\leq\sqrt{\frac{\alpha_{2}}{\alpha_{1}}}(1-\alpha_{3})^{n/2}\|\mathbf{x}_{t}\|.$ (5)
-Specifically, by setting the initial time to $t=0$ and denoting the current time step as $n=t\in\mathbb{Z}_{\geq 0}$, we obtain:
+Furthermore, by setting the initial time to $t=0$ and denoting the current time step as $n=t\in\mathbb{Z}_{\geq 0}$, we obtain the global trajectory bound:
 $\|\mathbf{x}_{t}\|\leq\sqrt{\frac{\alpha_{2}}{\alpha_{1}}}(1-\alpha_{3})^{t/2}\|\mathbf{x}_{0}\|.$
-Defining $C=\sqrt{\alpha_{2}/\alpha_{1}}$ and $\eta=\sqrt{1-\alpha_{3}}$, the result satisfies Definition 1, thereby proving that the control system (1) is exponentially stable with respect to $\mathbf{x}_{g}=\mathbf{0}$.
+By defining the constants $C=\sqrt{\alpha_{2}/\alpha_{1}}$ and $\eta=\sqrt{1-\alpha_{3}}$, the result satisfies Definition 1, thereby proving that the control system (1) is exponentially stable with respect to $\mathbf{x}_{g}=\mathbf{0}$.
 II-C Standard RL and Maximum Entropy RL
-In the reinforcement learning framework, the control system (1) is interpreted as a discrete-time deterministic environment, where $\mathcal{X}$ and $\mathcal{U}$ represent continuous state and action spaces, respectively.
-The application of control inputs to the system (1) corresponds to an agent interacting with this environment.
-Specifically, at each time step $t$, the agent observes the state $\mathbf{x}_{t}$ and executes an action $\mathbf{u}_{t}$ according to a policy $\pi$.
-Given a reward function $r:\mathcal{X}\times\mathcal{U}\to\mathbb{R}$, which is typically designed to penalize deviations from the target state $\mathbf{x}_{g}$ or reference $\mathbf{x}_{t}^{\text{ref}}$ in accordance with Problem 1, the environment returns a scalar reward $r_{t}:=r(\mathbf{x}_{t},\mathbf{u}_{t})$ and transitions to the next state $\mathbf{x}_{t+1}$.
-Consequently, the interaction sequence is denoted by the transition tuple $(\mathbf{x}_{t},\mathbf{u}_{t},r_{t},\mathbf{x}_{t+1})$.
-While classical control often employs deterministic laws, this article utilizes stochastic policies to facilitate exploration and robustness within the maximum entropy RL (MERL) framework.
-A stochastic policy is defined as a mapping $\pi:\mathcal{X}\to\mathcal{P}(\mathcal{U})$, where $\mathbf{u}_{t}\sim\pi(\cdot|\mathbf{x}_{t})$ is sampled from a probability distribution.
+To begin the analysis, we interpret the control system (1) as a discrete-time environment with continuous state space $\mathcal{X}$ and action space $\mathcal{U}$ under the standard RL paradigm.
+At each time step $t$, the RL agent observes the state $\mathbf{x}_{t}\in\mathcal{X}$ and executes an action $\mathbf{u}_{t}\in\mathcal{U}$, causing the environment to transition to the next state $\mathbf{x}_{t+1}$ and yield a scalar reward $r_{t}=r(\mathbf{x}_{t},\mathbf{u}_{t})$.
+This reward is formulated to penalize state deviations from the target $\mathbf{x}_{g}$ (or reference signal $\mathbf{x}_{t}^{\text{ref}}$) while accounting for the associated control effort.
+Furthermore, we employ stochastic policies in this article to facilitate exploration within the maximum entropy RL (MERL) framework.
+A stochastic policy is defined as a mapping $\pi:\mathcal{X}\to\mathcal{P}(\mathcal{U})$, where $\mathcal{P}(\mathcal{U})$ denotes the set of probability distributions over $\mathcal{U}$, and $\mathbf{u}_{t}\sim\pi(\cdot|\mathbf{x}_{t})$ is sampled from the probability distribution.
 We denote the state and state-action marginals induced by $\pi$ as $\rho_{\pi}(\mathbf{x}_{t})$ and $\rho_{\pi}(\mathbf{x}_{t},\mathbf{u}_{t})$, respectively.
-Unlike standard RL, which maximizes the expected cumulative discounted reward, MERL optimizes an entropy-augmented objective:
+In contrast to standard RL, which maximizes the expected cumulative discounted reward, MERL optimizes an entropy-augmented objective [33, 34, 35]:
 $J_{\pi}=\underset{(\mathbf{x}_{i\geq t},\mathbf{u}_{i\geq t})\sim\rho_{\pi}}{\mathbb{E}}\left[\sum_{i=t}^{\infty}\gamma^{i-t}\left[r_{i}+\alpha\mathcal{H}(\pi(\cdot|\mathbf{x}_{i}))\right]\right],$ (6)
-where $\gamma\in[0,1)$ is the discount factor, $\alpha>0$ is the temperature coefficient governing the trade-off between reward and exploration, and $\mathcal{H}(\pi(\cdot|\mathbf{x}))=\mathbb{E}_{\mathbf{u}\sim\pi(\cdot|\mathbf{x})}[-\log\pi(\mathbf{u}|\mathbf{x})]$ is the policy entropy.
-The corresponding soft Q-function, $Q^{\pi}(\mathbf{x}_{t},\mathbf{u}_{t})$, represents the expected soft return, which comprises both rewards and entropy terms, starting from $(\mathbf{x}_{t},\mathbf{u}_{t})$ and following $\pi$ thereafter.
+where $\gamma\in[0,1)$ is the discount factor, $\alpha>0$ is the temperature coefficient, and $\mathcal{H}(\pi(\cdot|\mathbf{x}))=\mathbb{E}_{\mathbf{u}\sim\pi(\cdot|\mathbf{x})}[-\log\pi(\mathbf{u}|\mathbf{x})]$ is the policy entropy.
+The corresponding soft Q-function, $Q^{\pi}(\mathbf{x}_{t},\mathbf{u}_{t})$, represents the expected soft return, which comprises both rewards and entropy terms.
 The soft Bellman operator $\mathcal{T}^{\pi}$ is defined as:
@@
 where $\mathbf{x}_{t+1}=f(\mathbf{x}_{t},\mathbf{u}_{t})$.
-In the soft policy iteration (SPI) algorithm [19], the policy is optimized by alternating between soft policy evaluation, where $Q^{\pi}$ is updated via iterative application of $\mathcal{T}^{\pi}$, and soft policy improvement, which is equivalent to finding a policy $\pi$ that maximizes the soft Q-function augmented with an entropy term for $\mathbf{x}_{t}$:
+In the soft policy iteration (SPI) algorithm [33], the policy is optimized by alternating between soft policy evaluation, where $Q^{\pi}$ is updated via $\mathcal{T}^{\pi}$, and soft policy improvement, which is equivalent to finding a policy $\pi$ that maximizes the soft Q-function for $\mathbf{x}_{t}$:
 $\pi_{\rm new}=\arg\max_{\pi}\underset{\mathbf{x}_{t}\sim\rho_{\pi},\mathbf{u}_{t}\sim\pi}{\mathbb{E}}\left[Q^{\pi_{\rm old}}(\mathbf{x}_{t},\mathbf{u}_{t})-\alpha\log\pi(\mathbf{u}_{t}|\mathbf{x}_{t})\right].$ (8)
-Although SPI converges to the optimal maximum entropy policy $\pi^{*}$ in tabular settings, practical control problems involve continuous spaces requiring function approximation.
+Although SPI converges to the optimal policy $\pi^{*}$ in tabular settings, the control problems studied in this article involve continuous state and action spaces necessitating function approximation.
 Thus, we parameterize the soft Q-function $Q_{\theta}(\mathbf{x}_{t},\mathbf{u}_{t})$ and the policy $\pi_{\phi}(\mathbf{u}_{t}|\mathbf{x}_{t})$ using neural networks with parameters $\theta$ and $\phi$.
-We assume $\pi_{\phi}$ follows a diagonal Gaussian distribution with mean $\mu_{\phi}(\mathbf{x}_{t})$ and standard deviation $\sigma_{\phi}(\mathbf{x}_{t})$.
-Actions are sampled using the reparameterization trick: $\mathbf{u}_{t}=g_{\phi}(\mathbf{x}_{t};\epsilon_{t})$, where $\epsilon_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{m})$, ensuring that the objective remains differentiable with respect to $\phi$.
+We assume $\pi_{\phi}$ follows a diagonal Gaussian distribution, and actions are sampled using the reparameterization trick [33]: $\mathbf{u}_{t}=g_{\phi}(\mathbf{x}_{t};\epsilon_{t})$, where $\epsilon_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{m})$, ensuring that the optimization objectives remain differentiable with respect to $\phi$.
 III Learning Lyapunov Certificate
-In this section, we develop the learning mechanism for the Lyapunov certificate within a model-free framework.
-We first introduce a sliding-window data collection strategy to satisfy multi-step trajectory requirements.
-Subsequently, we design a comprehensive loss function for the Lyapunov network, incorporating boundedness constraints and a multi-step stability loss.
-To enhance training efficiency and robustness, we employ trajectory-wise importance sampling for off-policy correction, a labeling mechanism for positive and negative samples, and a $\lambda$-weighted aggregation to balance the bias-variance trade-off.
+This section details the model-free Lyapunov certificate learning mechanism.
+We first introduce a sliding-window strategy for multi-step data collection, and then design a loss function for the Lyapunov network (also referred to as the Lyapunov certificate), incorporating both boundedness and stability constraints.
 III-A Data Collection
-Our algorithm is developed within the maximum entropy RL framework, where data is acquired through agent-environment interactions with the system (1).
-Unlike the standard SAC algorithm [19] that samples single-step transitions, we leverage the multi-step Lyapunov condition presented in (4), which necessitates data sequences spanning $n$ consecutive steps.
-We define a single-step interaction tuple as $\mathbf{d}_{t}=(\mathbf{x}_{t},\mathbf{u}_{t},r_{t},\pi_{\phi}(\mathbf{u}_{t}|\mathbf{x}_{t}),\mathbf{x}_{t+1})$, where $\pi_{\phi}(\mathbf{u}_{t}|\mathbf{x}_{t})$ represents the probability density of sampling action $\mathbf{u}_{t}$ at state $\mathbf{x}_{t}$.
-The collected multi-step data consists of sequences $\{\mathbf{d}_{t},\mathbf{d}_{t+1},\dots,\mathbf{d}_{t+n-1}\}$, which are stored in the experience replay buffer $\mathcal{D}$.
-Data storage commences only after $n$ consecutive steps have been accumulated.
-Specifically, during the initial n1n-1 steps of any rollout, no data is committed to $\mathcal{D}$.
-Once the rollout length reaches or exceeds $n$, the collection process begins.
-As illustrated in Fig.
-1, we employ a double-ended queue (deque) to temporarily buffer interaction data.
-During each agent-environment interaction, new data enters the deque.
-Once the deque reaches a length of $n$, the complete $n$-step sequence is transferred to $\mathcal{D}$, and the oldest data point is removed to maintain the deque size, enabling a continuous sliding window for data collection.
-Figure 1: Schematic of the $n$-step sliding window data collection process.
-A double-ended queue (deque) of size $n$ is utilized to buffer consecutive interaction transitions $\mathbf{d}_{t}$, ensuring that complete $n$-step trajectories are available for the verification of the multi-step Lyapunov certificate.
+In contrast to conventional Lyapunov-based RL algorithms [25, 36], which primarily target asymptotic stability through single-step transitions, our proposed MSACL enforces the stricter exponential stability requirement.
+Guided by the multi-step relationship (4), establishing exponential stability necessitates leveraging multi-step sequence samples from state trajectories.
+This sequential data structure provides a sufficient temporal horizon to validate the contraction properties of Lyapunov certificates.
+To operationalize this, we define the fundamental transition tuple as $\mathbf{d}_{t}=(\mathbf{x}_{t},\mathbf{u}_{t},r_{t},\pi_{\phi}(\mathbf{u}_{t}|\mathbf{x}_{t}),\mathbf{x}_{t+1})$.
+Unlike standard replay buffers that store these tuples individually, we employ a double-ended queue (deque) as a sliding window to capture temporal correlations.
+Upon reaching a length of $n$, the complete multi-step sequence $\{\mathbf{d}_{t},\mathbf{d}_{t+1},\dots,\mathbf{d}_{t+n-1}\}$ is packaged and transferred to the replay buffer $\mathcal{D}$, while the oldest entry is concurrently removed to maintain the window size.
+Figure 1: Schematic of $n$-step sliding window data collection. A double-ended queue (deque) of size $n$ buffers consecutive transitions $\mathbf{d}_{t}$ to provide complete multi-step trajectories.
 III-B Loss Function Design of Lyapunov Network
-Similar to the soft Q-network and policy network, we introduce a Lyapunov network parameterized by $\psi$, denoted as $V_{\psi}$, to estimate the Lyapunov certificate.
-In designing the loss function for the Lyapunov certificate, we strictly adhere to the conditions established in Lemma 1, which encompass the boundedness and multi-step stability requirements.
-(1) Boundedness Loss: Based on the first condition in Lemma 1, we construct the boundedness loss function to ensure the Lyapunov certificate is quadratically constrained:
+Similar to the soft Q-network and policy network, we introduce a Lyapunov network $V_{\psi}$ parameterized by $\psi$ to estimate the Lyapunov certificate.
+In designing the loss function, We strictly adhere to the boundedness and multi-step stability conditions in Lemma 1.
+(1) Boundedness Loss. Based on the boundedness condition in (3), we construct the boundedness loss function to ensure the Lyapunov certificate to be quadratically constrained:
 $\mathcal{L}_{\text{Bnd}}(\psi)=\frac{1}{Nn}$ $\sum_{i=1}^{N}\sum_{k=0}^{n-1}\Big[\max\left(\alpha_{1}\|\mathbf{x}_{t+k}^{i}\|^{2}-V_{\psi}(\mathbf{x}_{t+k}^{i}),0\right)$ (9)
 $+\max\left(V_{\psi}(\mathbf{x}_{t+k}^{i})-\alpha_{2}\|\mathbf{x}_{t+k}^{i}\|^{2},0\right)\Big],$
-where $N$ denotes the number of multi-step trajectories in a batch, $i$ denotes the $i$-th multi-step sample in the batch, and $n$ represents the number of steps within each trajectory.
-This loss term explicitly constrains $V_{\psi}$ to satisfy the boundedness condition defined in (3).
-(2) Multi-step Stability Loss: Consider a sampled multi-step trajectory $\{\mathbf{d}_{t},\dots,\mathbf{d}_{t+n-1}\}$ generated by a previous policy $\pi_{\phi_{\text{old}}}$.
-Since our goal is to train the Lyapunov certificate $V_{\psi}$ to be compatible with the current policy $\pi_{\phi_{\text{new}}}$ for subsequent policy updates, we must account for the off-policy nature of the data.
-We introduce trajectory-wise importance sampling to correct the distribution shift.
-For an $n$-step trajectory, the importance sampling ratio for the segment from $t$ to $t+k$ is defined as ($k=1,2,\dots,n-1$):
+where $N$ denotes the batch size of multi-step sequence samples, $i$ denotes the $i$-th sample, and $n$ is the sequence length. This loss enforces the boundedness condition on $V_{\psi}$.
+(2) Multi-step Stability Loss.
+A major challenge in learning Lyapunov certificates from off-policy data is the discrepancy between the behavior policy $\pi_{\phi_{\text{old}}}$ and the target policy $\pi_{\phi}$.
+To mitigate this discrepancy, we employ a sequence-wise importance sampling strategy.
+Specifically, for a multi-step sequence $\{\mathbf{d}_{t},\dots,\mathbf{d}_{t+n-1}\}$ generated by $\pi_{\phi_{\text{old}}}$, we compute the cumulative importance sampling ratio for the transition from $t$ to $t+k$ ($k=1,\dots,n-1$) as:
 $\text{IS}_{\text{clip},k}=\prod_{j=t}^{t+k-1}\min\left(\frac{\pi_{\phi}(\mathbf{u}_{j}|\mathbf{x}_{j})}{\pi_{\phi_{\text{old}}}(\mathbf{u}_{j}|\mathbf{x}_{j})},1\right).$ (10)
 Note that we apply a clipping operation to the importance sampling ratios to prevent variance explosion and enhance training stability.
-Furthermore, we address the second condition in (3), namely the stability requirement.
-While (4) represents the ideal theoretical decay, empirical trajectories may not strictly satisfy the exponential stability bound in (5) during early stages of training.
-To bridge the gap between the stability condition and the stochastic samples, we introduce the concepts of positive and negative samples.
+On the other hand, learning Lyapunov certificates directly from off-policy data may not be accurate, as empirical sequences in the early stages of training often fail to satisfy the exponential decay condition (5), leading to unreliable certificate estimation.
+To establish a rigorous correspondence between empirical data and theoretical stability requirements, we propose the Exponential Stability Label (ESL) mechanism.
+Specifically, we introduce positive and negative samples to bridge the gap between theory and practice:
 Definition 2
-Within an $n$-step sample, for a state $\mathbf{x}_{t+k}$, if it satisfies the exponential decay condition:
+Within an $n$-step sequence sample, for the state $\mathbf{x}_{t+k}$, if it satisfies the exponential decay condition:
 $\|\mathbf{x}_{t+k}\|\leq\sqrt{\frac{\alpha_{2}}{\alpha_{1}}}(1-\alpha_{3})^{k/2}\|\mathbf{x}_{t}\|,$
 then $\mathbf{d}_{t+k}$ is classified as a positive sample; otherwise, it is classified as a negative sample.
-Accordingly, we define the Exponential Stability Label (ESL) as:
+Accordingly, we define the Exponential Stability Label as:
 $\text{ESL}_{k}=\begin{cases}1,&\text{if }\|\mathbf{x}_{t+k}\|\leq\sqrt{\frac{\alpha_{2}}{\alpha_{1}}}(1-\alpha_{3})^{k/2}\|\mathbf{x}_{t}\|\\ -1,&\text{otherwise}.\end{cases}$
 This labeling mechanism allows the Lyapunov network to distinguish between transitions that contribute to system stability and those that violate the certificate requirements.
-For trajectories composed of multi-step interaction data, we introduce a weighting factor $\lambda\in(0,1)$, inspired by the (truncated) $\lambda$-return in reinforcement learning [33], to balance the trade-off between bias and variance.
-Specifically, in the stability condition $V_{\psi}(\mathbf{x}_{t+k})\leq(1-\alpha_{3})^{k}V_{\psi}(\mathbf{x}_{t})$ from (4), the parameter $k$ determines the temporal horizon of the system’s evolution.
-When $k$ is small, the short trajectory horizon provides a limited view of the system’s long-term dynamics under policy $\pi_{\phi}$.
-Given that the Lyapunov network is randomly initialized and potentially inaccurate in the early stages of training, relying on short-horizon data can introduce significant estimation bias.
-Conversely, a larger $k$ captures more extensive information regarding the trajectory evolution (extended horizon), thereby providing a stronger and more accurate learning signal that reduces bias.
-However, since each transition is sampled from the environment, every step introduces stochasticity; consequently, longer trajectories accumulate higher cumulative variance.
-Therefore, a small $k$ leads to high bias but low variance, while a large $k$ results in low bias but high variance.
-To optimize this trade-off, we employ the weighting factor $\lambda$ to aggregate multi-step information.
-By integrating the positive/negative sample labels with the weighting factor $\lambda$, we construct the Lyapunov stability loss to enforce a rigorous correspondence between empirical samples and the stability condition (4):
+Furthermore, to balance the bias-variance trade-off in multi-step sequences, we introduce a weighting factor $\lambda\in(0,1)$, inspired by the (truncated) $\lambda$-return in classic RL literature [40].
+Specifically, for the stability condition $V_{\psi}(\mathbf{x}_{t+k})\leq(1-\alpha_{3})^{k}V_{\psi}(\mathbf{x}_{t})$ from (4), small $k$ determines a short horizon with limited information.
+Given that the Lyapunov network is randomly initialized and potentially inaccurate in the early stages of training, relying on short-horizon data may introduce high estimation bias.
+Conversely, large $k$ captures extended information to reduce bias but accumulates environmental stochasticity into higher variance.
+To this end, we employ the weighting factor $\lambda$ to aggregate multi-step information and balance bias-variance trade-off.
+Finally, by integrating the positive/negative sample labels with the weighting factor $\lambda$, we construct the Lyapunov stability loss to enforce a rigorous correspondence between empirical samples and the stability condition (4):
 $\mathcal{L}_{\text{Stab}}(\psi)=\frac{1}{N}\sum_{i=1}^{N}\left[\sum_{k=1}^{n-1}\frac{\lambda^{k-1}\mathcal{L}_{\text{diff},k}^{i}(\psi)}{\sum_{j=1}^{n-1}\lambda^{j-1}}\right],$ (11)
-and the step-wise difference loss $\mathcal{L}_{\text{diff},k}^{i}(\psi)$ is defined as:
+where the step-wise difference loss $\mathcal{L}_{\text{diff},k}^{i}(\psi)$ is defined as:
 $\mathcal{L}_{\text{diff},k}^{i}(\psi)=$ $\text{IS}_{\text{clip},k}^{i}\cdot\max\Big(0,$
 $\text{ESL}_{k}^{i}\cdot\left(V_{\psi}(\mathbf{x}_{t+k}^{i})-(1-\alpha_{3})^{k}V_{\psi}(\mathbf{x}_{t}^{i})\right)\Big).$
-(3) Combined Lyapunov Loss: By integrating the boundedness and multi-step stability losses, the total loss function for optimizing the Lyapunov network is formulated as:
+(3) Combined Lyapunov Loss. By integrating the boundedness loss and the multi-step stability loss, the comprehensive loss function for optimizing the Lyapunov network $V_{\psi}$ is formulated as:
 $\mathcal{L}_{\text{Lya}}(\psi)=\omega_{\text{Bnd}}\mathcal{L}_{\text{Bnd}}(\psi)+\omega_{\text{Stab}}\mathcal{L}_{\text{Stab}}(\psi),$ (12)
@@
 IV Learning Policy with Exponential Stability Constraints
-This section presents the optimization framework for the proposed MSACL algorithm.
-We first establish the soft Q-network updates using multi-step Bellman residuals to enhance data efficiency.
-We then introduce a novel policy loss function that incorporates a stability advantage derived from the Lyapunov certificate, utilizing a PPO-style clipping mechanism to ensure robust convergence toward exponentially stable behavior.
-Finally, we describe the automated temperature adjustment and synthesize the complete actor-critic learning procedure.
+This section details the MSACL optimization framework.
+We first establish soft Q-network updates using multi-step Bellman residuals to enhance data efficiency.
+Then we introduce a novel policy loss incorporating Lyapunov-based stability advantages with PPO-style clipping for robust exponential stabilization.
+Finally, we describe automated temperature adjustment and the complete actor-critic procedure.
 IV-A Loss Function Design of Soft Q Networks
-Based on the soft Bellman operator (7), we define the sampled Bellman residual for the $i$-th transition within a multi-step sequence for the soft Q-network $Q_{\theta}$ as:
+Following the soft Bellman operator (7), we define the sampled Bellman residual for the $i$-th multi-step sequence sample of $Q_{\theta}$ as:
 $\mathcal{B}_{t}^{i}(\theta)=Q_{\theta}(\mathbf{x}_{t}^{i},\mathbf{u}_{t}^{i})-\Big[$ $r_{t}^{i}+\gamma Q_{\bar{\theta}}(\mathbf{x}_{t+1}^{i},\mathbf{u}_{t+1}^{i})$ (13)
 $-\alpha\log\pi_{\phi}(\mathbf{u}_{t+1}^{i}|\mathbf{x}_{t+1}^{i})\Big],$
-where $Q_{\bar{\theta}}$ denotes the target network parameterized by $\bar{\theta}$, which is updated via polyak averaging to stabilize the learning process.
-Utilizing the collected multi-step sequences, the loss function for the soft Q-network is defined as the mean squared Bellman error (MSBE) across the entire trajectory:
+where $Q_{\bar{\theta}}$ denotes the target network parameterized by $\bar{\theta}$.
+Utilizing collected multi-step sequence samples, the loss function for the soft Q-network is defined as the mean squared Bellman error (MSBE) across the entire sequence:
 $\mathcal{L}_{\text{SoftQ}}(\theta)=\frac{1}{Nn}\sum_{i=1}^{N}\sum_{k=0}^{n-1}(\mathcal{B}_{t+k}^{i}(\theta))^{2}.$ (14)
-While standard implementations of maximum entropy RL algorithms, such as SAC [19], typically set the upper limit of the inner summation in (14) to zero (i.e., $n=1$) to utilize only single-step transitions, our formulation leverages the complete $n$-step sequence for parameter updates.
-Theoretically, minimizing the MSBE requires state-action pairs $(\mathbf{x}_{t},\mathbf{u}_{t})$ to be sampled from a distribution that ensures an unbiased stochastic gradient estimate.
-Specifically, the objective is given by:
-$\mathcal{J}_{\text{SoftQ}}(\theta)=$ $\underset{\mathbf{x}_{t}\sim\rho_{\pi},\mathbf{u}_{t}\sim\pi}{\mathbb{E}}\Big[Q_{\theta}(\mathbf{x}_{t},\mathbf{u}_{t})-\Big(r_{t}+\gamma\underset{\mathbf{x}_{t+1}\sim\rho_{\pi},\mathbf{u}_{t+1}\sim\pi}{\mathbb{E}}$
+In contrast to standard MERL algorithms such as SAC [33], which typically utilize single-step transition (i.e., $n=1$), our approach leverages the complete $n$-step sequence to minimize the following objective function:
+$\mathcal{J}_{\text{SoftQ}}(\theta)=$ $\underset{\mathbf{x}_{t}\sim\rho_{\pi},\mathbf{u}_{t}\sim\pi}{\mathbb{E}}\Big[Q_{\theta}(\mathbf{x}_{t},\mathbf{u}_{t})-\Big(r_{t}+\gamma\underset{\mathbf{x}_{t+1}\sim\rho_{\pi},\mathbf{u}_{t+1}\sim\pi}{\mathbb{E}}$ (15)
 $\big(Q_{\bar{\theta}}(\mathbf{x}_{t+1},\mathbf{u}_{t+1})-\alpha\log\pi_{\phi}(\mathbf{u}_{t+1}|\mathbf{x}_{t+1})\big)\Big)\Big]^{2}.$
-Strictly speaking, the transitions within a multi-step sequence are temporally correlated, and utilizing the full sequence simultaneously deviates from the independent and identically distributed (i.i.d.) sampling requirement commonly assumed in stochastic optimization.
-However, this approach represents a deliberate trade-off between mathematical rigor and sample efficiency.
-By incorporating all transitions within each collected sequence into the loss function, we significantly enhance the data utilization rate and the information density of each training batch.
-Experimental evidence suggests that as long as the sequence length $n$ remains moderate, the bias introduced by these short-term correlations does not undermine the stability of the Q-function learning.
-In practice, the benefits of improved sample efficiency and accelerated convergence often outweigh the theoretical concerns regarding strict uniform sampling, ultimately contributing to more robust control performance.
 Remark 1
-In the practical implementation of Algorithm 1, we utilize clipped double-Q learning with two networks, $Q_{\theta_{1}}$ and $Q_{\theta_{2}}$, to mitigate the overestimation bias of soft Q-values, as motivated by the TD3 framework [25].
-Furthermore, we incorporate the “delayed policy update” technique to ensure that the policy is updated against a more stable value function, thereby facilitating smoother convergence of the actor-critic process.
+Although minimizing (15) theoretically necessitates independent and identically distributed (i.i.d.) sampling for unbiased gradient estimation [41], our proposed multi-step approach strategically balances mathematical rigor with enhanced sample efficiency.
+Indeed, while $n$-step sequences introduce temporal correlations that deviate from strict i.i.d.
+assumptions, experimental evidence demonstrates that moderate sequence lengths maintain learning stability.
+Ultimately, the substantial gains in batch data information outweigh the minor bias introduced by short-term correlations, leading to accelerated convergence and more robust control performance.
 IV-B Loss Function Design of Policy Network
-This subsection details the optimization process for the policy network, which integrates the maximum entropy objective with a Lyapunov-based stability constraint.
-In the standard SAC framework [19], policy optimization aims to maximize the expected soft Q-value.
-For the $i$-th multi-step sequence, the SAC-component of the loss function is expressed as:
+This subsection details the policy network optimization, which seamlessly integrates the maximum entropy objective with a novel Lyapunov-based stability guidance mechanism.
+(1) Soft-Q Value Maximization.
+Following the policy optimization (8) in SPI, the first objective is to maximize the soft-Q value.
+For the $i$-th multi-step sequence sample $\{\mathbf{d}_{t}^{i},\dots,\mathbf{d}_{t+n-1}^{i}\}$, the SAC-style component of the loss function is expressed as:
 $\mathcal{L}_{\pi,\text{SAC}}^{i}(\phi)=\frac{1}{n}\sum_{k=0}^{n-1}\Big[$ $Q_{\theta}(\mathbf{x}_{t+k}^{i},g_{\phi}(\mathbf{x}_{t+k}^{i};\epsilon_{t+k}^{i}))$
 $-\alpha\log\pi_{\phi}(g_{\phi}(\mathbf{x}_{t+k}^{i};\epsilon_{t+k}^{i})|\mathbf{x}_{t+k}^{i})\Big].$
-To ensure the learned policy satisfies the exponential stability requirements, we utilize the learned Lyapunov certificate $V_{\psi}$ to guide the policy updates.
-Inspired by the Generalized Advantage Estimation (GAE) in PPO [24], we define a stability advantage function based on $V_{\psi}$, which serves as a metric for convergence performance:
+Note that $g_{\phi}(\mathbf{x}_{t+k}^{i};\epsilon_{t+k}^{i})$ is obtained via the reparameterization trick [33].
+(2) Stability-Aware Advantage Guidance.
+To ensure exponential stability, we attempt to utilize the learned Lyapunov certificate $V_{\psi}$ to guide policy updates.
+Inspired by the Generalized Advantage Estimation (GAE) in PPO [31], we transform the Lyapunov descent property into an active guidance signal.
+Specifically, we define the stability advantage function based on $V_{\psi}$ at $t+k$ as a metric for convergence performance:
 $A_{t,k}=(1-\alpha_{3})^{k}V_{\psi}(\mathbf{x}_{t})-V_{\psi}(\mathbf{x}_{t+k}).$ (16)
-A positive stability advantage ($A_{t,k}\geq 0$) indicates that the system trajectory from $\mathbf{x}_{t}$ to $\mathbf{x}_{t+k}$ strictly adheres to the exponential decay condition.
-Maximizing this advantage encourages the policy to select actions that accelerate the decay of $V_{\psi}$, thereby directly increasing the state convergence rate toward the equilibrium $\mathbf{x}_{g}$.
-Conversely, $A_{t,k}<0$ signifies a violation of the stability certificate, representing a disadvantage that the policy must learn to avoid.
-This stability-driven optimization paradigm aligns with recent developments in stable model-free control, such as D-learning [34, 35].
-To robustly handle the bias-variance trade-off across varying horizons, we apply the weighting factor $\lambda$ to obtain the aggregated stability advantage:
+A positive stability advantage ($A_{t,k}\geq 0$) indicates that the system trajectory adheres to the exponential decay condition (4).
+Maximizing this advantage encourages actions that accelerate $V_{\psi}$ decay, thereby increasing the state convergence rate toward $\mathbf{x}_{g}$.
+Conversely, $A_{t,k}<0$ signifies a certificate condition violation that the policy must learn to avoid.
+This stability-aware optimization aligns with recent developments in stable model-free control, such as D-learning [26, 27, 28].
+Similar to (11), in order to balance the bias-variance trade-off across varying horizons, we apply the weighting factor $\lambda$ to obtain the aggregated stability advantage:
 $A_{t,\lambda}=\sum_{k=1}^{n-1}\frac{\lambda^{k-1}A_{t,k}}{\sum_{j=1}^{n-1}\lambda^{j-1}}.$ (17)
-For a multi-step sample $\{\mathbf{d}_{t},\dots,\mathbf{d}_{t+n-1}\}$, we define the policy importance sampling ratio based on the initial transition as $\rho^{i}(\phi)=\frac{\pi_{\phi}(\mathbf{u}_{t}^{i}|\mathbf{x}_{t}^{i})}{\pi_{\phi_{\text{old}}}(\mathbf{u}_{t}^{i}|\mathbf{x}_{t}^{i})}$.
-Incorporating the clipping mechanism to prevent excessively large policy updates, the final policy loss function is formulated as:
+(3) Integrated Policy Optimization Objective.
+For the $i$-th multi-step sequence sample, we define the first-step importance sampling (IS) ratio as $\rho^{i}(\phi)=\frac{\pi_{\phi}(\mathbf{u}_{t}^{i}|\mathbf{x}_{t}^{i})}{\pi_{\phi_{\text{old}}}(\mathbf{u}_{t}^{i}|\mathbf{x}_{t}^{i})}$.
+By incorporating the aggregated stability advantage (17) with PPO-style clipping to prevent excessively large updates, the policy loss function is formulated as:
 $\mathcal{L}_{\pi}(\phi)=-\frac{1}{N}\sum_{i=1}^{N}\Big[$ $\mathcal{L}_{\pi,\text{SAC}}^{i}(\phi)+\min\Big(\rho^{i}(\phi)A_{t,\lambda}^{i},$ (18)
@@
 Remark 2
-It is important to distinguish between the two importance sampling (IS) strategies employed in this article.
-In (11), sequence-level IS (from $t$ to t+k1t+k-1) is used to train the Lyapunov certificate $V_{\psi}$.
-While sequence-level IS typically introduces higher variance, it provides the low-bias correction necessary for the Lyapunov network to accurately reflect the stability of the current policy $\pi_{\phi}$ using off-policy data.
-Since the policy parameters are fixed during the Lyapunov update phase, this variance remains manageable via the truncation in (10).
-However, when optimizing the policy $\pi_{\phi}$ itself, sequence-level IS would introduce prohibitive variance into the gradient estimates, often leading to training divergence.
-Consequently, following the design principle of PPO, we utilize only the first-step IS ratio $\rho(\phi)$ in (17).
-This approach effectively reduces the variance of the policy gradient, ensuring stable optimization while still benefiting from the multi-step stability information embedded within the weighted advantage $A_{t,\lambda}$.
-IV-C Automated Temperature Adjustment
-In maximum entropy RL, the temperature coefficient $\alpha$ in (6) is critical for balancing exploration and exploitation.
+Notably, we employ distinct importance sampling strategies for the critic and the actor.
+In (11), sequence-level IS provides low-bias correction for learning $V_{\psi}$ from off-policy data.
+This variance remains manageable via truncation in (10) since policy parameters are fixed during Lyapunov certificate updates.
+However, during policy optimization, where policy parameters are actively updated, sequence-level IS is avoided to prevent prohibitive variance and potential divergence.
+Following the PPO principle, we utilize only the first-step IS ratio $\rho^{i}(\phi)$ in (18).
+This ensures stable optimization while retaining multi-step stability information via the aggregated stability advantage $A_{t,\lambda}$.
+IV-C Unified Actor-Critic Optimization Framework
+Under the MERL paradigm, the temperature coefficient $\alpha$ in (6) is critical for balancing exploration and exploitation.
 Following the automated entropy adjustment mechanism introduced in [33], we update $\alpha$ by minimizing the following loss function:
 $\mathcal{L}_{\text{Ent}}(\alpha)=\frac{1}{Nn}\sum_{i=1}^{N}\sum_{k=0}^{n-1}\left[-\alpha\log\pi_{\phi}(\mathbf{u}_{t+k}^{i}|\mathbf{x}_{t+k}^{i})-\alpha\bar{\mathcal{H}}\right],$ (19)
-where $\bar{\mathcal{H}}$ denotes the target entropy, typically set as $\bar{\mathcal{H}}=-m$, with $m$ representing the dimension of the action space.
-By integrating the loss functions for the soft Q-network $Q_{\theta}$, the Lyapunov network $V_{\psi}$, the policy network $\pi_{\phi}$, and the temperature coefficient $\alpha$, we iteratively update the parameters within an actor-critic framework using the Adam optimizer [37].
-The complete algorithmic procedure is summarized in Algorithm 1, and the overall flowchart is illustrated in Fig.
+where $\bar{\mathcal{H}}$ denotes the target entropy.
+By integrating the loss functions for the soft Q-network $Q_{\theta}$, Lyapunov network $V_{\psi}$, policy network $\pi_{\phi}$, and temperature coefficient $\alpha$, we iteratively update these parameters within an actor-critic framework.
+The complete procedure is summarized in Algorithm 1 with flowchart illustrated in Fig.
 2.
-Figure 2: Flowchart of the proposed MSACL algorithm.
+Figure 2: Flowchart of the MSACL algorithm.
 Algorithm 1 Multi-Step Actor-Critic Learning with Lyapunov Certificates (MSACL)
@@
    Update $V_{\psi}$ via $\psi\leftarrow\psi-\beta_{\psi}\nabla_{\psi}\mathcal{L}_{\text{Lya}}(\psi)$
-   Update $Q_{\theta_{i}}$ via $\theta_{i}\leftarrow\theta_{i}-\beta_{\theta}\nabla_{\theta_{i}}\mathcal{L}_{\text{SoftQ}}(\theta_{i})$ 11footnotemark: 1
-   if $\left(\text{iteration count}\mod d\right)=0$ then 22footnotemark: 2
+   Update $Q_{\theta_{i}}$ via $\theta_{i}\leftarrow\theta_{i}-\beta_{\theta}\nabla_{\theta_{i}}\mathcal{L}_{\text{SoftQ}}(\theta_{i})$
+   if $\left(\text{iteration count}\mod d\right)=0$ then
      Update $\pi_{\phi}$ via $\phi\leftarrow\phi-\beta_{\phi}\nabla_{\phi}\mathcal{L}_{\pi}(\phi)$
@@
 Remark 3
-In the flowchart of Algorithm 1, we incorporate two implementation refinements.
-Specifically, regarding footnote 1: for the soft Q-loss function (14), we replace the target value $Q_{\bar{\theta}}$ with $\min\{Q_{\bar{\theta}_{1}},Q_{\bar{\theta}_{2}}\}$ to mitigate overestimation bias, as noted in Remark 1.
-Regarding footnote 2: we implement delayed updates for both the policy network $\pi_{\phi}$ and the temperature coefficient $\alpha$.
-However, within each delayed update cycle, we perform $d$ consecutive updates for these parameters as a “delayed update compensation” mechanism to ensure adequate policy improvement.
-These techniques are adapted from the cleanrl framework [38], and our experiments demonstrate that integrating these refinements significantly enhances overall performance and training stability.
+Algorithm 1 incorporates two implementation refinements.
+First, for the soft Q-loss (14), we replace $Q_{\theta}$ with $\min\{Q_{\theta_{1}},Q_{\theta_{2}}\}$ to mitigate overestimation bias.
+Second, we implement delayed updates for both policy $\pi_{\phi}$ and temperature $\alpha$, performing $d$ consecutive updates within each delayed cycle to ensure sufficient policy improvement.
+These techniques, adapted from the cleanrl framework [42], significantly improve training stability and performance.
 V Experiments
-In this section, we conduct a comprehensive experimental evaluation of the MSACL algorithm across six benchmark environments, encompassing both stabilization and high-dimensional nonlinear tracking tasks.
-We compare our approach against baseline model-free RL algorithms and Lyapunov-based RL algorithms using rigorous performance and stability metrics.
-The analysis further includes the visualization of learned Lyapunov certificates, extensive robustness and generalization testing under parametric uncertainties and stochastic noise, and a sensitivity study investigating the impact of the multi-step horizon $n$ on the learning process.
-V-A Benchmark Environments
-Figure 3: Benchmark environments. (a) Uncontrolled VanderPol (Phase Portrait), (b) Pendulum, (c) Planar DuctedFan, (d) Two-link planar robot, (e) SingleCarTracking, (f) QuadrotorTracking.
-In this subsection, we present the benchmark environments utilized to evaluate the experimental performance of the proposed and baseline algorithms, encompassing both stabilization and tracking tasks.
-We categorize these environments by their primary control objectives.
-For stabilization tasks, the environments include: Controlled VanderPol (hereafter referred to as VanderPol), Pendulum, Planar DuctedFan (DuctedFan), and Two-link planar robot (Two-link).
-For tracking tasks, the benchmark comprises SingleCarTracking and QuadrotorTracking.
-As illustrated in Figure 3, comprehensive specifications and dynamical parameters for all environments are provided in Appendix VI-B.
-In stabilization tasks, the control objective is to drive the system state to the equilibrium point, i.e., $\mathbf{x}_{g}=\mathbf{0}\in\mathbb{R}^{n}$.
-In tracking tasks, the goal is to ensure the system state follows a time-varying reference trajectory $\mathbf{x}_{t}^{\text{ref}}$, which may represent desired coordinates, velocities, or attitudes.
-Across all tasks, we employ a unified reward structure:
+This section evaluates the MSACL algorithm across six benchmarking environments, covering stabilization and high-dimensional tracking tasks.
+We compare our approach against model-free and Lyapunov-based RL baselines using performance and stability metrics.
+Further analysis includes Lyapunov certificate visualization, robustness assessment under parametric uncertainty and stochastic noise, and sensitivity studies on the multi-step horizon $n$.
+V-A Benchmarking Environments
+Figure 3: Benchmarking environments. (a) VanderPol (Uncontrolled Phase Portrait), (b) Pendulum, (c) Planar DuctedFan, (d) Two-link Planar Robot, (e) SingleCarTracking, (f) QuadrotorTracking.
+This subsection presents the benchmarking environments for evaluating MSACL and baseline algorithms.
+We categorize environments by control objective: stabilization tasks include Controlled VanderPol (VanderPol)[43], Pendulum[44], Planar DuctedFan (DuctedFan)[45], and Two-link planar robot (Two-link)[46]; tracking tasks comprise SingleCarTracking[47] and QuadrotorTracking[48], as illustrated in Figure 3.
+For stabilization tasks, the goal is to drive the state to the origin $\mathbf{x}_{g}=\mathbf{0}\in\mathbb{R}^{d}$; for tracking tasks, the system follows a time-varying reference signal $\mathbf{x}_{t}^{\text{ref}}$.
+All tasks use a unified reward structure:
 $r_{t}=\underbrace{-\left(\mathbf{x}_{t}^{\top}\mathbf{Q_{r}}\mathbf{x}_{t}+\mathbf{u}_{t}^{\top}\mathbf{R_{r}}\mathbf{u}_{t}\right)}_{\text{Penalty}}+\underbrace{r_{\mathrm{approach}}}_{\text{Encouragement}}$ (20)
-where the first term is a penalty term inspired by the performance functionals commonly utilized in optimal control [39], with $\mathbf{Q_{r}}$ and $\mathbf{R_{r}}$ being positive definite weight matrices.
-This term penalizes deviations from the equilibrium and excessive control effort at each time step, effectively minimizing the cumulative cost $\sum_{t=0}^{\infty}\left[\mathbf{x}_{t}^{\top}\mathbf{Q_{r}}\mathbf{x}_{t}+\mathbf{u}_{t}^{\top}\mathbf{R_{r}}\mathbf{u}_{t}\right]$.
-For tracking tasks, the state $\mathbf{x}_{t}$ in the penalty term is replaced by the tracking error $\mathbf{e}_{t}=\mathbf{x}_{t}-\mathbf{x}_{t}^{\text{ref}}$.
-The second term provides sparse encouragement, defined as:
+where the first term penalizes state deviations and control effort using positive definite matrices $\mathbf{Q_{r}}$ and $\mathbf{R_{r}}$, minimizing the cumulative cost $\sum_{t=0}^{\infty}\left[\mathbf{x}_{t}^{\top}\mathbf{Q_{r}}\mathbf{x}_{t}+\mathbf{u}_{t}^{\top}\mathbf{R_{r}}\mathbf{u}_{t}\right]$ used in optimal control [49].
+For tracking tasks, $\mathbf{x}_{t}$ is replaced by the tracking error $\mathbf{e}_{t}=\mathbf{x}_{t}-\mathbf{x}_{t}^{\text{ref}}$.
+The second term $r_{\text{approach}}$ provides sparse encouragement, defined as:
 $r_{\mathrm{approach}}=\begin{cases}r_{\delta},&\text{if }\|\mathbf{x}_{t}\|_{\infty}\leq\delta\\ 0,&\text{otherwise}\end{cases}$
-Specifically, when the state approaches the equilibrium (i.e., its infinity norm falls below a small threshold $\delta$, such as 0.01), the environment provides a constant positive reward $r_{\delta}$ (e.g., +1) to further incentivize convergence or minimize the steady-state tracking error.
-This reward design is intentionally parsimonious and intuitive.
-Subsequent experiments demonstrate that under this straightforward scheme, our algorithm successfully achieves the control objectives across all benchmark environments.
-These six benchmark environments are constructed following the OpenAI Gym framework [40].
-Notably, our implementation of the Pendulum environment is significantly more rigorous than the standard OpenAI Gym version.
-Specifically, we expand the initial state space to $\pi,\pi]\times[-10,10]$ and the action range to [5,5][-5,5], while introducing boundary-based termination conditions and extending the maximum episode length to 1000 steps.
-These modifications increase the control difficulty, intentionally designed to highlight the performance disparities between different algorithms.
-Detailed specifications for this and other environments are provided in Appendix VI-B.
-During each agent-environment interaction, the environment returns a standardized 5-tuple $(\mathbf{x}_{t},\mathbf{u}_{t},r_{t},\pi_{\phi}(\mathbf{u}_{t}|\mathbf{x}_{t}),\mathbf{x}_{t+1})$, where the state transition from $\mathbf{x}_{t}$ to $\mathbf{x}_{t+1}$ is updated using the standard explicit Euler integration method.
+Specifically, when $\|\mathbf{x}\|_{\infty}<\delta$ (e.g., 0.01), a positive reward $r_{\delta}$ is provided to incentivize convergence and mitigate steady-state error.
+All benchmarks follow the OpenAI Gym framework [50].
+Notably, the Pendulum implementation is modified for increased difficulty, featuring expanded state and action spaces, boundary-based termination, and longer episodes, to better differentiate algorithmic performance.
+Specific details can be found in our code repository.
+During interaction, the environment returns a 5-tuple $(\mathbf{x}_{t},\mathbf{u}_{t},r_{t},\pi_{\phi}(\mathbf{u}_{t}|\mathbf{x}_{t}),\mathbf{x}_{t+1})$, with state transitions updated via explicit Euler integration.
 V-B Baseline Algorithms
-To evaluate the effectiveness of the proposed MSACL algorithm, we compare it against classic model-free RL algorithms, including SAC [19] and PPO [24], as well as state-of-the-art Lyapunov-based RL algorithms, specifically LAC [28] and POLYC [27].
-While SAC and PPO have demonstrated superior performance across various benchmarks, they are evaluated here on stabilization and tracking tasks to provide a performance baseline.
-LAC and POLYC incorporate Lyapunov-based constraints into standard RL frameworks to provide stability guarantees; these approaches serve as primary inspirations for our method.
-In the LAC algorithm, the authors replace the standard reward signal with a quadratic cost function, which is defined across environments as:
+We evaluate MSACL against classic model-free RL baselines (SAC [33], PPO [31]) and state-of-the-art Lyapunov-based algorithms (LAC [36], POLYC [25]).
+SAC and PPO serve as performance baselines for stabilization and tracking tasks, while LAC and POLYC provide stability guarantees through Lyapunov constraints.
+For LAC, the standard reward is replaced with a quadratic cost function defined as:
 $c_{t}=\mathbf{x}_{t}^{\top}\mathbf{Q_{c}}\mathbf{x}_{t},$
-where $\mathbf{Q_{c}}$ is a positive definite matrix.
-Notably, this formulation considers only state-related penalties, whereas our reward definition $r_{t}$ explicitly accounts for both state deviations and control effort.
-By comparing against these baselines, we aim to demonstrate the advantages of MSACL in terms of stability metrics and convergence speed.
-To ensure a fair and consistent comparison, all algorithms are implemented within the General Optimal control Problem Solver (GOPS) framework [41], an open-source PyTorch-based RL library.
-Within this unified architecture, the actor and critic networks utilize identical structures consisting of multi-layer perceptrons (MLPs) with two hidden layers of 256 units each and ReLU activation functions.
-All actor networks employ stochastic policies parameterized by diagonal Gaussian distributions.
-Regarding the critic architectures: LAC utilizes only a Lyapunov network, while POLYC incorporates both a value function and a Lyapunov certificate.
-For fairness, SAC and LAC also implement clipped double Q-learning and delayed policy updates as discussed in Remark 1.
-Parameter updates for all algorithms are performed using the Adam optimizer.
-The core hyperparameter configurations are summarized in Table I, with full specifications available in our code repository, which will be publicly released.
+where $\mathbf{Q_{c}}$ is positive definite.
+Note that in contrast to this state-only penalty formulation, our reward $r_{t}$ accounts for both state deviations and control effort.
+By comparing against these baselines, we aim to demonstrate the advantages of MSACL in terms of performance and stability metrics.
+For fair comparison, all algorithms are implemented in the GOPS framework [51].
+Actor and critic networks utilize MLPs with two 256-unit hidden layers and ReLU activations.
+All actors employ stochastic diagonal Gaussian policies.
+Critic architectures vary: LAC uses only a Lyapunov network, while POLYC incorporates both value and Lyapunov functions.
+SAC and LAC also implement clipped double Q-learning and delayed policy updates according to Remark 3.
+Optimization is performed via the Adam optimizer [30], with core hyperparameters listed in Table I.
 TABLE I: Hyperparameters in the algorithms
@@
    Target smoothing coefficient ($\tau$) 0.05
-   Reward/Cost scale 100
    Optimizer Adam (default $\beta_{1}$ and $\beta_{2}$)
@@
 V-C Performance Evaluation and Stability Analysis
-In this subsection, we present the training outcomes, evaluate stability metrics under the learned policies, and visualize the obtained Lyapunov certificates.
-To ensure statistical significance, each algorithm was trained using five random seeds per benchmark.
-The training progress is illustrated in Fig.
-4, utilizing the following performance metrics:
-  • AMCR (Average Mean Cumulative Reward): The sample mean of the Mean Cumulative Reward (MCR) computed across multiple independent training trials.
-The MCR is defined as the total undiscounted reward accumulated throughout an episode, normalized by the episode length to provide a consistent per-step measure of policy performance.
-  • AMCC (Average Mean Cumulative Cost): The sample mean of the Mean Cumulative Cost (MCC) computed across multiple independent training trials.
-The MCC is defined as the total undiscounted cost accumulated throughout an episode, normalized by the episode length to quantify the average per-step magnitude of state deviations and control effort.
+This subsection presents training outcomes, stability metrics, and Lyapunov certificate visualizations.
+Each algorithm is trained across five random seeds, with training results illustrated in Fig.
+4 using the following performance metrics:
+  • AMCR (Average Mean Cumulative Reward): The average of the Mean Cumulative Reward (MCR) computed across five independent training trials.
+MCR represents the total undiscounted reward accumulated per episode, normalized by the episode length to measure per-step performance.
+  • AMCC (Average Mean Cumulative Cost): The average of the Mean Cumulative Cost (MCC) computed across five independent training trials.
+MCC represents the total undiscounted cost accumulated per episode, normalized by the episode length to quantify the average per-step magnitude of state deviations.
 (a) VanderPol
@@
 MSACL (Ours)            SAC [33]            LAC [36]            PPO [31]            POLYC [25]
-Figure 4: Training performance across six benchmark environments.
+Figure 4: Training performance across six benchmarks.
 The results show the average MCR and MCC over five independent runs.
@@
 The horizontal axis indicates the number of training steps.
-The training curves demonstrate that MSACL outperforms or matches all baseline algorithms across all environments.
-Notably, the AMCR for MSACL remains positive and stable during the late training phase.
-Given our reward design, a positive MCR indicates that the system state has successfully entered the target $\delta$-neighborhood.
-Higher reward values further imply faster convergence, as the agent spends a greater proportion of the episode receiving the positive reward $r_{\delta}$.
-For subsequent comparative analysis, we select the optimal policy model for each algorithm, defined as the specific model that achieved the highest MCR among the five independent training runs.
-These optimal policies are evaluated from 100 fixed random initial states across all benchmarks.
-To rigorously assess stability and convergence, we utilize the following metrics across four target radii ($0.2,0.1,0.05,\text{and }0.01$):
-  • Reach Rate (RR): The percentage of trajectories that enter the specified radius.
-  • Average Reach Step (ARS): The mean number of steps required to first enter the radius.
-  • Average Hold Step (AHS): The mean number of steps the state remains within the radius after the first entry.
+Training curves show MSACL outperforms or at least matches all baselines across six benchmarks.
+A positive AMCR indicates that the system state successfully enters the target $\delta$-neighborhood, while higher reward values imply faster convergence as the agent spends more time receiving $r_{\delta}$.
+For comparison, we select the optimal policy for each algorithm, defined as the model that achieved the highest MCR among five training runs.
+These optimal policies are evaluated from 100 fixed initial states, with stability and convergence are assessed across four target radii ($0.2,0.1,0.05,0.01$) using the following stability metrics:
+  • Reach Rate (RR): Percentage of trajectories that enter the specified radius.
+  • Average Reach Step (ARS): Mean steps that required for the state to first enter the radius.
+  • Average Hold Step (AHS): Mean steps that the state remains within the radius after the initial entry.
 The evaluation results are summarized in Table II.
-TABLE II: Experimental Results of Baseline Algorithms Across Benchmark Environments
+TABLE II: Experimental Results of Baseline Algorithms Across Benchmarking Environments
 Benchmark Algorithm AMCR AMCC RR / ARS / AHS at different radii
@@
 POLYC -17528.1 $\pm$ 2162.6 17318.7 $\pm$ 2155.6 0.0 / – / – 0.0 / – / – 0.0 / – / – 0.0 / – / –
-  • Note: The values for AMCR and AMCC represent the mean performance of the optimal policy (defined as the policy achieving the highest evaluation MCR) averaged over 100 independent trials, with $\pm$ denoting one standard deviation.
-Stability metrics (RR, ARS, and AHS) are evaluated across four distinct radii, where “–” indicates that the system failed to enter the specified radius within the episode duration.
-Bold entries highlight the superior results: for performance metrics, the maximum AMCR and minimum AMCC are bolded; for stability metrics, the highest RR is prioritized for bolding, with the shortest ARS and longest AHS selected in the event of identical RR values.
-These results highlight the superiority of MSACL.
-Across all six environments, MSACL consistently achieves the highest AMCR.
-In terms of AMCC, MSACL attains the absolute minimum in the Pendulum, SingleTrackCar, and QuadrotorTracking tasks, while remaining highly competitive in others.
-Furthermore, MSACL achieves the highest RR across all radii.
-Most significantly, in the high-dimensional nonlinear QuadrotorTracking task, MSACL is the only method to achieve a 100% reach rate at the $0.01$ precision level, whereas baselines fail to meet this stringent requirement.
-These results underscore the robustness and precision of MSACL in complex control applications.
-Finally, we visualize the Lyapunov networks $V_{\psi}$ corresponding to the optimal policies.
-The learned Lyapunov certificates and their associated contour plots are shown in Fig.
+  • Note: AMCR and AMCC values represent the mean performance of the optimal policy (highest evaluation MCR) averaged over 100 independent trials, with $\pm$ denoting one standard deviation.
+Stability metrics (RR, ARS, and AHS) are evaluated across four radii; – indicates the radius was not reached.
+Bold entries highlight superior results: maximum AMCR, minimum AMCC, and highest RR.
+For stability metrics, the highest RR is prioritized for bolding, with the shortest ARS and longest AHS bolded for identical RR values.
+Evaluation results demonstrate MSACL’s superiority.
+Specifically, MSACL consistently achieves the highest AMCR across all six benchmarks.
+Regarding AMCC, MSACL attains the absolute minimum in the Pendulum, SingleCarTracking, and QuadrotorTracking tasks, while remaining competitive in others.
+Furthermore, MSACL maintains the highest RR across all radii.
+Notably, in the high-dimensional QuadrotorTracking task, MSACL is the only approach to achieve a 100% reach rate at the $0.01$ precision level, whereas baselines fail.
+These results demonstrate the speed and precision of MSACL in complex control systems.
+Finally, we visualize the Lyapunov networks $V_{\psi}$ for the optimal policies.
+The learned certificates and associated contour plots are shown in Fig.
 5.
-For high-dimensional systems, the plots represent the relationship between $V_{\psi}$ and two specific state components while setting all other components to zero.
+For high-dimensional systems, the plots illustrate the relationship between $V_{\psi}$ and two specific state components with all other components set to zero.
 Remark 4
-Since MSACL is built upon the SAC framework, SAC serves as the primary baseline for our ablation study.
+Since MSACL builds on the SAC framework, SAC serves as the primary baseline for our ablation study.
 As shown in Fig.
-4 and Table II, the integration of the Lyapunov certificate significantly enhances both training performance and stability metrics compared to the standard SAC algorithm.
-This validates that embedding Lyapunov-based constraints into maximum entropy RL effectively yields neural policies with superior stability guarantees for deterministic control systems.
-Figure 5: Visualization of learned Lyapunov certificates and their corresponding contour plots across various benchmark environments: (a) VanderPol; (b) Pendulum; (c)–(e) DuctedFan (projections in the $(x,y)$-plane with $\theta=0$, $(y,\theta)$-plane with $x=0$, and $(x,\theta)$-plane with $y=0$); (f) Two-link; (g) SingleCarTracking; (h)–(j) QuadrotorTracking (projections in the $(e_{x},e_{y})$-plane with $e_{z}=0$, $(e_{y},e_{z})$-plane with $e_{x}=0$, and $(e_{x},e_{z})$-plane with $e_{y}=0$).
+4 and Table II, the integration of the Lyapunov certificate significantly enhances both training performance and stability metrics compared to the standard SAC.
+This validates that embedding Lyapunov-based constraints into MERL effectively yields neural policies with superior stability guarantees for model-free control systems.
+Figure 5: Visualization of learned Lyapunov certificates and their corresponding contour plots across six benchmarks: (a) VanderPol; (b) Pendulum; (c)–(e) DuctedFan (projections in the $(x,y)$-plane with $\theta=0$, $(y,\theta)$-plane with $x=0$, and $(x,\theta)$-plane with $y=0$); (f) Two-link; (g) SingleCarTracking; (h)–(j) QuadrotorTracking (projections in the $(e_{x},e_{y})$-plane with $e_{z}=0$, $(e_{y},e_{z})$-plane with $e_{x}=0$, and $(e_{x},e_{z})$-plane with $e_{y}=0$).
 V-D Evaluation of Robustness and Generalization
-It is well known that neural policies, characterized by high-dimensional parameter spaces, are susceptible to overfitting within specific training environments.
-This can lead to degraded robustness against stochastic noise and parametric variations, as well as limited generalization to unseen scenarios.
-In this subsection, we evaluate the robustness and generalization capabilities of the policies trained via MSACL to demonstrate their applicability in real-world environments with inherent uncertainties.
-For stabilization tasks (VanderPol, Pendulum, DuctedFan, and Two-link), we introduce process noise $\bm{\epsilon}_{t}\sim\mathcal{N}(\mathbf{0},\bm{\sigma}_{\text{env}})$ to the system dynamics, resulting in the perturbed form:
+Neural policies with high-dimensional parameter spaces are prone to overfitting, often resulting in reduced robustness to noise and parametric variations, and limited generalization to unseen scenarios.
+This subsection evaluates the robustness and generalization of MSACL-trained policies, with the aim of demonstrating their real-world applicability under inherent uncertainties.
+For stabilization tasks (VanderPol, Pendulum, DuctedFan, Two-link), we introduce process noise $\bm{\epsilon}_{t}\sim\mathcal{N}(\mathbf{0},\sigma_{\text{env}}\cdot\mathbf{1}_{d})$ to the system dynamics, where $\mathbf{1}_{d}$ denotes the all-ones vector of dimension $d$, yielding the perturbed form:
 $\mathbf{x}_{t+1}=f(\mathbf{x}_{t},\mathbf{u}_{t})+\bm{\epsilon}_{t}.$
 Additionally, we modify key physical parameters of the systems to assess whether the MSACL-trained policies can maintain stability under environmental variations.
-For tracking tasks (SingleCarTracking and QuadrotorTracking), we evaluate robustness by injecting dynamical noise and assess generalization by requiring the policies to track reference trajectories $\mathbf{x}_{t}^{\text{ref}}$ that differ significantly from those encountered during training.
-The specific parameter modifications and noise magnitudes utilized for these evaluations are detailed in Table III.
+For tracking tasks (SingleCarTracking and QuadrotorTracking), robustness is evaluated via dynamical noise, while generalization is assessed using reference trajectories $\mathbf{x}_{t}^{\text{ref}}$ that are significantly different from those in training.
+Table III details the specific parameter modifications and noise magnitudes.
 TABLE III: System Parameters and Noise Magnitudes
@@
 QuadrotorTracking $0.01,0.02$
-The robustness testing results for the stabilization tasks are illustrated in Fig.
-6.
-Despite the introduction of parametric perturbations and process noise, all system trajectories originating from the specified initial states successfully converge to the neighborhood of the equilibrium $\mathbf{x}_{g}=\mathbf{0}$.
-The results for the tracking tasks are presented in Fig.
-7 and Fig.
-8.
-It is evident that the trained neural policies effectively mitigate noise interference within a reasonable range.
-In the SingleCarTracking environment, while the policy was trained on straight-line references, we evaluate its generalization using circular and sinusoidal trajectories.
-Experimental validation demonstrates excellent tracking performance; as shown in the position error plot in Fig.
-7e, the tracking error $e_{\mathrm{xy}}=\sqrt{e_{x}^{2}+e_{y}^{2}}$ remains within 0.1 during the steady-state phase.
-For the QuadrotorTracking task, the policy was trained using a horizontal helical trajectory.
-During generalization testing, we evaluate the policy against two unseen trajectories: a vertical helix and a Lissajous curve, as shown in Fig.
-8c and Fig.
-8d.
-Notably, the MSACL-trained policy demonstrates superior generalization, with the tracking error $e_{\mathrm{xyz}}=\sqrt{e_{x}^{2}+e_{y}^{2}+e_{z}^{2}}$ stabilizing around 0.02 in the later stages.
-Figure 6: System trajectories under the neural policies in stabilization tasks subjected to environmental parametric uncertainties and process noise: (a) VanderPol with state variables $\mathbf{x}=[x,\dot{x}]^{\top}$; (b) Pendulum with angular position $\theta$ and angular velocity $\dot{\theta}$; (c) DuctedFan with coordinates $x,y$ and orientation $\theta$; (d) Two-link with joint angles $q_{1}$ and $q_{2}$.
-All trajectories exhibit robust convergence to the neighborhood of the equilibrium $\mathbf{x}_{g}=\mathbf{0}$ despite external disturbances.
+Robustness results for stabilization tasks in Fig.
+6 show that all trajectories converge to the equilibrium $\mathbf{x}_{g}=\mathbf{0}$ despite parametric perturbations and process noise.
+Tracking task results (Fig.
+7, 8) demonstrate that MSACL-trained policies effectively mitigate process noise within reasonable ranges.
+For SingleCarTracking, though trained on straight-line references, the policy generalizes well to circular and sinusoidal trajectories (Fig.
+7c, 7d), with position error $e_{\mathrm{xy}}=\sqrt{e_{x}^{2}+e_{y}^{2}}$ remaining below 0.1 in steady state (Fig.
+7e).
+For QuadrotorTracking, though trained on horizontal helical trajectories, the policy successfully tracks unseen vertical helix and Lissajous curves (Fig.
+8c, 8d), with tracking error $e_{\mathrm{xyz}}=\sqrt{e_{x}^{2}+e_{y}^{2}+e_{z}^{2}}$ stabilizing around 0.02 (Fig.
+8e), demonstrating superior generalization capability.
+Figure 6: System trajectories under the MSACL-trained policies in stabilization tasks subjected to environmental parametric uncertainties and process noise: (a) VanderPol with state variables $\mathbf{x}=[x,\dot{x}]^{\top}$; (b) Pendulum with angular position $\theta$ and angular velocity $\dot{\theta}$; (c) DuctedFan with coordinates $x,y$ and orientation $\theta$; (d) Two-link with joint angles $q_{1}$ and $q_{2}$.
+All trajectories exhibit robust convergence to the neighborhood of the equilibrium $\mathbf{x}_{g}=\mathbf{0}$.
 Figure 7: Robustness (a)–(b) and generalization (c)–(e) evaluation for the SingleCarTracking task: (a) Car trajectories versus the straight-line reference $\mathbf{x}_{t}^{\text{ref}}=\{[t,0]\in\mathbb{R}^{2}\mid t\in[0,50]\}$ under varying noise levels; (b) Evolution of the Euclidean position error $e_{\mathrm{xy}}=\sqrt{e_{x}^{2}+e_{y}^{2}}$ corresponding to (a); (c) Tracking performance for a circular reference trajectory $\{[8\cos(0.125t),8\sin(0.125t)]^{\top}\in\mathbb{R}^{2}\mid t\in[0,50]\}$ using the policy trained on straight lines; (d) Tracking performance for a sinusoidal reference trajectory $\{[t,\sin(0.2t)]^{\top}\in\mathbb{R}^{2}\mid t\in[0,50]\}$; (e) Evolution of the position error $e_{\mathrm{xy}}$ for the circular and sinusoidal cases shown in (c) and (d).
@@
 The results highlight the policy’s robust tracking capability under environmental noise and its effective generalization to geometrically distinct reference trajectories.
-Based on these results, we conclude that policies trained via the MSACL algorithm exhibit robust performance against moderate uncertainties and favorable generalization to unseen reference trajectories.
-However, our testing also reveals that the policies’ efficacy diminishes under extreme disturbances or environmental parameters that deviate substantially from the training distribution.
-In such cases of significant domain shift, the MSACL algorithm can be readily redeployed to retrain the policy, ensuring stable and high-precision control for the specific operational scenario.
-V-E Sensitivity Analysis of the Multi-Step Horizon $n$
-In this subsection, we investigate the influence of the multi-step horizon parameter $n$ on the performance of the MSACL algorithm across all benchmark environments.
-Following the evaluation protocol detailed in Section V-C, we utilize both performance metrics (AMCR and AMCC) and stability metrics (RR, ARS, and AHS) to assess the policies trained under varying configurations of $n$.
-Specifically, we conduct 100 randomized experiments for each benchmark by generating and fixing a set of 100 random initial states.
-Policies trained with different $n$ values interact with the environment starting from these fixed states to compute the aforementioned evaluation metrics.
-The results across all environments are summarized in Table IV.
-The experimental findings reveal a consistent trend: as $n$ increases, the stabilization and tracking performance of the MSACL-trained policies generally improves.
-This improvement occurs because a larger $n$ provides a longer-horizon trajectory during training, enabling the policy $\pi$ to more accurately identify the stability direction, namely the descent direction of the Lyapunov certificate $V_{\psi}$.
-This reduces the training bias and reinforces the stability guarantees for the resulting neural policy.
-Empirically, setting $n=20$ yields satisfactory performance across all benchmark environments, suggesting it as a versatile default parameter.
-However, an excessively large $n$ may not always lead to optimal results.
-As analyzed in Section III-B, longer temporal horizons introduce greater cumulative variance into the training data, which can distort the optimization landscape and result in suboptimal policies.
-Therefore, selecting an appropriate $n$ is essential to balancing the bias-variance trade-off.
-In summary, for the control tasks presented in this article, $n=20$ delivers consistently robust results; for specific applications requiring peak performance, fine-tuning around this value is recommended to achieve the optimal balance between training bias and variance.
+These results demonstrate that MSACL-trained policies exhibit robustness to model uncertainties and process noise, while generalizing effectively to unseen reference signals.
+However, performance degrades under extreme disturbances or significant deviations from the training distribution.
+In such cases, MSACL can be redeployed to retrain policies for stable, high-precision control in specific operational scenarios.
+V-E Sensitivity Analysis of the Multi-Step Horizon
+This subsection examines the impact of multi-step horizon $n$ on MSACL performance across all benchmarks.
+Following the protocol in Section V-C, we evaluate policies trained with varying $n$ using performance metrics (AMCR, AMCC) and stability metrics (RR, ARS, AHS).
+For each benchmark, we conduct 100 experiments from fixed random initial states to compute these metrics.
+As summarized in Table IV, performance generally improves with larger $n$, as a longer horizon enables the policy to better identify the stability direction, i.e., the descent of $V_{\psi}$.
+This reduces training bias and reinforces stability guarantees.
+Empirically, $n=20$ delivers satisfactory performance across all benchmarks, serving as a versatile default.
+However, excessively large $n$ may degrade performance due to increased cumulative variance in training data, potentially distorting the optimization landscape as analyzed in Section III-B.
+Therefore, selecting an appropriate $n$ is essential to balance the bias-variance trade-off.
+For the control tasks herein, $n=20$ provides consistently robust results; for better performance, fine-tuning around this value is recommended.
 TABLE IV: Experimental Results of MSACL Algorithm with different multi-step horizon Parameters $n$
@@
 $n=20$ 89.8 $\pm$ 7.5 0.4 $\pm$ 0.2 1.0 / 18.1 / 982.0 1.0 / 25.0 / 975.8 1.0 / 28.6 / 972.2 1.0 / 35.2 / 965.7
-SingleTrackCar $n=1$ -51.0 $\pm$ 194.9 46.4 $\pm$ 62.6 0.64 / 32.0 / 964.0 0.64 / 64.6 / 931.1 0.64 / 79.7 / 921.2 0.64 / 92.1 / 908.8
+SingleCarTracking $n=1$ -51.0 $\pm$ 194.9 46.4 $\pm$ 62.6 0.64 / 32.0 / 964.0 0.64 / 64.6 / 931.1 0.64 / 79.7 / 921.2 0.64 / 92.1 / 908.8
 $n=5$ -40.2 $\pm$ 217.9 35.0 $\pm$ 58.8 0.74 / 31.6 / 968.7 0.74 / 56.1 / 943.8 0.74 / 78.9 / 920.9 0.74 / 113.0 / 887.8
@@
 VI-A Discussion
-This subsection provides an analytical discussion on the functional roles of the various components within the MSACL framework.
-As highlighted in Remark 4, MSACL extends the Soft Actor-Critic (SAC) algorithm by integrating multi-step Lyapunov certificate learning, thereby establishing formal stability guarantees while preserving the exploration efficiency of maximum entropy reinforcement learning.
-In model-free stabilization and tracking tasks, conventional SAC, which lacks the structural guidance provided by Lyapunov theory, may fail to converge to stable policies under improper reward structures.
-Even with well-designed rewards, standard RL often requires exhaustive exploration of the action space to identify stabilizing directions, leading to prohibitive training times.
-To address these limitations, our approach introduces a Lyapunov network that serves as a stability-oriented critic, providing explicit directional guidance during policy updates by identifying action sequences that facilitate the descent of the Lyapunov certificate.
-The theoretical foundation of this mechanism is rooted in Lemma 1, which establishes the equivalence between the Lyapunov stability conditions and exponential state convergence.
-To ensure the Lyapunov certificate $V_{\psi}$ is learned accurately, we developed a training methodology that aligns theoretical requirements with empirical trajectories.
-Specifically, the introduction of Exponential Stability Labels (ESL) in Definition 2 allows the algorithm to categorize trajectory segments as positive or negative samples based on their adherence to exponential decay criteria.
-This classification enables the construction of the loss function (11), which effectively bridges abstract stability conditions with observed system dynamics.
-The policy optimization process leverages this learned certificate through the stability advantage mechanism defined in (17).
-This advantage function quantifies the rate of Lyapunov descent along trajectories, prioritizing policy updates that simultaneously ensure stability and accelerate convergence.
-Our experimental results confirm that this dual-objective optimization yields policies that achieve both stability and rapid convergence across diverse and challenging benchmarks.
-A critical insight is that the efficacy of the policy is fundamentally tied to the fidelity of the learned Lyapunov certificate.
-The multi-step sampling strategy, combined with the ESL-based loss formulation, ensures that $V_{\psi}$ accurately reflects the theoretical stability criteria, providing reliable gradients for policy improvement.
-While developed for stabilization and tracking, the MSACL framework is highly extensible to other stability-critical applications.
-Potential extensions include trajectory planning for autonomous vehicles [42], spacecraft proximity operations [43], and robotic manipulation tasks requiring rigorous safety guarantees [44, 45].
-The integration of Lyapunov-based guarantees within maximum entropy RL represents a significant step toward developing verifiable, learning-based control systems for real-world deployment.
+This subsection elucidates the functional roles of MSACL components.
+As highlighted in Remark 4, MSACL augments the SAC framework with a stability-oriented critic, departing from standard RL that relies on exhaustive exploration to discover stabilizing behaviors.
+Instead, MSACL leverages a learned Lyapunov certificate to provide explicit, structured guidance for policy updates.
+Theoretically, MSACL is grounded in Lemma 1, which establishes the equivalence between Lyapunov stability and exponential convergence.
+To operationalize this exponential stability requirement, the ESL mechanism (Definition 2) categorizes trajectory segments based on their contraction properties, transforming the stability condition into a supervised learning task.
+This ensures the synthesis of a high-fidelity Lyapunov certificate $V_{\psi}$ that accurately captures the underlying system dynamics.
+Such fidelity is crucial for providing the actor with theoretically sound, stability-aware gradients.
+Consequently, the stability advantage in (18) prioritizes actions that facilitate Lyapunov descent, effectively guiding the policy toward the region of stability.
+While developed for stabilization and tracking, MSACL is extensible to safety-critical applications including autonomous vehicle trajectory planning [52], spacecraft proximity operations [53], and safety-guaranteed robotic manipulation [54, 55].
+The integration of Lyapunov guarantees within MERL represents a step toward verifiable learning-based control.
 VI-B Conclusion
-This article presented the MSACL (Multi-Step Actor-Critic Learning with Lyapunov certificates) algorithm, a novel framework that integrates exponential stability guarantees into the maximum entropy RL paradigm.
-By embedding Lyapunov stability theory within the SAC architecture, MSACL achieves stable control for model-free dynamical systems without sacrificing the exploration benefits of entropy regularization.
-To rigorously evaluate the proposed methodology, we developed six benchmark environments covering four stabilization and two tracking tasks.
-Comprehensive comparisons against established baselines demonstrate the consistent superiority of MSACL across both performance and stability metrics.
-Notably, these results were achieved using a simple reward structure that penalizes state deviations and control effort while providing sparse positive rewards near the equilibrium.
-Beyond standard performance, we conducted extensive robustness analyses against environmental uncertainties and generalization tests on unseen reference trajectories.
-The MSACL-trained neural policies exhibited remarkable resilience to disturbances and effective adaptation to geometrically distinct paths.
-A key algorithmic contribution is the systematic analysis of the multi-step horizon parameter $n$.
-We established that $n=20$ serves as a robust default setting, effectively balancing the bias-variance trade-off in stability learning.
-This work advances the integration of formal stability guarantees within learning-based control, providing a foundation for verifiable policies in safety-critical applications.
-Future research will explore extensions to partially observable systems [46, 47], adaptive Lyapunov architectures [48, 49], and transfer learning mechanisms for complex, cross-domain dynamical systems [50, 51].
-[Benchmark environments information]
-This appendix provides a detailed description of the six nonlinear dynamical systems selected as benchmarks for this study: VanderPol, Pendulum, DuctedFan, Two-link, SingleCarTracking, and QuadrotorTracking.
-Widely recognized as canonical benchmarks in nonlinear control (NLC) [52, 53, 54], reinforcement learning (RL) [55, 56], and optimal control literature [57], these environments constitute a comprehensive testbed for evaluating the proposed methods.
-They encompass a broad spectrum of dynamic characteristics, ranging from self-sustaining oscillations and underactuated mechanics to complex tire–road friction interactions and rigid-body dynamics on the $\text{SE}(3)$ manifold.
-To provide a unified evaluation framework, we categorize these benchmarks based on their control objectives: Stabilization and Tracking.
-The formal task definitions and their corresponding reward structures are detailed below.
--C Unified Control Objectives
-Despite the distinct physical dynamics inherent to each system, the core control tasks are formulated using a consistent mathematical structure.
-Stabilization Tasks [58, 55, 59]
-The first four systems (VanderPol, Pendulum, DuctedFan, and Two-link) are configured as stabilization tasks.
-For these systems, the control objective is to find a policy that generates a sequence of control inputs $\{\mathbf{u}_{t}\}_{t=0}^{T-1}$ to drive the system state $\mathbf{x}_{t}$ from a randomly sampled initial state $\mathbf{x}_{0}$ toward a fixed equilibrium point $\mathbf{x}_{g}$ (typically the origin, $\mathbf{x}_{g}=\mathbf{0}$) and maintain stability at this equilibrium for the duration of the episode.
-Tracking Tasks [60, 54, 61]
-The remaining two systems (SingleCarTracking, and QuadrotorTracking) are formulated as trajectory tracking tasks.
-Here, the objective is defined in terms of the tracking error $\mathbf{e}_{t}=\mathbf{x}_{t}-\mathbf{x}_{t}^{\text{ref}}$.
-The goal is to determine a control sequence such that the error trajectory $\{\mathbf{e}_{t}\}_{t=0}^{T}$ converges to the zero-error state $\mathbf{e}_{g}=\mathbf{0}$, thereby ensuring that the system state $\mathbf{x}_{t}$ accurately follows a time-varying reference trajectory $\mathbf{x}_{t}^{\text{ref}}$.
--D Unified Reward Function Design
-To align the learning process with the aforementioned control objectives, we employ a consistent reward function structure across all six environments:
-$r_{t}=\underbrace{-\left(\mathbf{s}_{t}^{\top}\mathbf{Q}_{r}\mathbf{s}_{t}+\mathbf{u}_{t}^{\top}\mathbf{R}_{r}\mathbf{u}_{t}\right)}_{\text{Penalty}}+\underbrace{r_{\text{approach}}(\mathbf{s}_{t})}_{\text{Encouragement}},$
-where $\mathbf{s}_{t}$ represents the generalized state vector, defined as $\mathbf{s}_{t}=\mathbf{x}_{t}$ for stabilization tasks and $\mathbf{s}_{t}=\mathbf{e}_{t}$ for tracking tasks.
-The matrices $\mathbf{Q}_{r}$ and $\mathbf{R}_{r}$ are positive semi-definite weighting matrices that penalize state deviations and control effort, respectively.
-The specific parameter configurations for each environment, including the diagonal elements of $\mathbf{Q}_{r}$ and $\mathbf{R}_{r}$, are summarized in Table V.
-The term $\mathbb{I}(\cdot)$ denotes the indicator function, which provides a sparse encouragement reward when the system enters a predefined tolerance region around the target state.
-For all environments, the state cost matrix $\mathbf{Q}_{c}$ is simply defined as a identity matrix $\mathbf{I}_{n}$ of appropriate dimensions, where $n$ denotes the dimension of state space.
-TABLE V: Reward Function Parameters for All Benchmark Systems
-System State Penalty $\mathbf{Q}_{r}$ (diag) Control Penalty $\mathbf{R}_{r}$ (diag) Encouragement ($r_{\text{approach}}$)
-VanderPol $[2,1]$ $[0.1]$ $\mathbb{I}(\|\mathbf{x}_{t}\|_{\infty}\leq 10^{-2})$
-Pendulum $[2,1]$ $[0.1]$ $\mathbb{I}(\|\mathbf{x}_{t}\|_{\infty}\leq 10^{-2})$
-DuctedFan $[2,2,2,1,1,1]$ $[0.1,0.1]$ $\mathbb{I}(\|\mathbf{x}_{t}\|_{\infty}\leq 10^{-2})$
-Two-link $[2,2,1,1]$ $[0.1,0.1]$ $\mathbb{I}(\|\mathbf{x}_{t}\|_{\infty}\leq 10^{-2})$
-SingleCarTracking $[2,2,1,1,1,1,1]$ $[0.1,0.1]$ $\mathbb{I}(\|\mathbf{e}_{t}\|_{\infty}\leq 10^{-2})$
-QuadrotorTracking $\mathbf{I}_{12}$ $10^{-4},0.01,0.01,0.01]$ $10(1-\frac{\|\mathbf{e}_{t}\|_{\infty}}{0.1})\cdot\mathbb{I}(\|\mathbf{e}_{t}\|_{\infty}\leq 0.1)$
--E Details of VanderPol
--E1 System Overview
-The VanderPol oscillator, originally proposed by Balthasar van der Pol in 1920 [62], is a classic model describing self-sustaining oscillatory circuits characterized by nonlinear damping.
-This system serves as a fundamental benchmark for electronic circuit stabilization (e.g., vacuum tube oscillators), mechanical vibration control, and biological modeling, such as cardiac pacemaker dynamics [63].
-In this study, we extend the autonomous system with a control input to evaluate stabilization performance within a nonlinear regime.
--E2 Dynamics and Physical Parameters
-The continuous-time dynamics of controlled VanderPol are governed by the following second-order nonlinear differential equations:
-$\dot{x}_{1}$ $=x_{2},$
-$\dot{x}_{2}$ $=\mu\left(1-x_{1}^{2}\right)x_{2}-x_{1}+u,$
-where $x_{1}$ and $x_{2}$ represent the oscillator displacement and velocity, respectively.
-The physical meanings of the variables are summarized in Table VI, and the specific parameter values and simulation settings used in our experiments are detailed in Table VII.
-TABLE VI: Variable Definitions for Controlled VanderPol
-Variable Physical Meaning
-$x_{1}$ Displacement of the oscillator (m)
-$x_{2}$ Velocity of the oscillator (m/s)
-$\dot{x}_{2}$ Acceleration of the oscillator (m/s2)
-$\mu$ Nonlinear damping strength parameter (dimensionless)
-$u$ Control input force applied to the oscillator (N)
-TABLE VII: Parameter Values for Controlled VanderPol
-Parameter Value
-$\mu$ (nonlinear strength) 1.0 (dimensionless)
-Simulation time step ($dt$) 0.01 s
-Control update interval 5 simulation steps
-Maximum episode steps 1000
--E3 State and Action Spaces
-For the reinforcement learning formulation, the system state vector is defined as $\mathbf{x}=[x_{1},x_{2}]^{\top}$.
-To reflect the physical limits of the oscillator and ensure numerical stability during training, the state space is constrained to $\mathcal{X}=[-10,10]\times[-10,10]$, where both displacement and velocity are bounded within the range of $\pm 10$.
-The control objective is to drive the system from a random initial configuration to the equilibrium state (goal state) defined as $\mathbf{x}_{g}=[0,0]^{\top}$, corresponding to the stable origin of the phase plane.
-The control action $u$ is a one-dimensional continuous input restricted to the action space $\mathcal{U}=[-5,5]$.
-These bounds are selected to provide sufficient authority for stabilization while preventing excessively large inputs that could violate physical actuation constraints or lead to divergent gradients.
-At the start of each episode, the initial state $\mathbf{x}_{0}$ is sampled from a uniform distribution $\mathbf{x}_{0}\sim\text{Uniform}([-5,5]\times[-5,5])$.
-This initialization range is intentionally chosen to place the system within a significant nonlinear regime, ensuring that the controller must actively counteract the self-sustaining oscillations to achieve stabilization.
--F Details of Pendulum
--F1 System Overview
-The Pendulum is a classic rigid-body system with a pivot point constrained to a vertical plane, where the primary control objective is to stabilize the system in an inverted (upright) position against gravity.
-Beyond its role as a fundamental pedagogical tool for illustrating nonlinear stability and feedback control [64], its underlying principles are essential for robotic locomotion in bipedal systems [65], aerospace attitude control, and industrial crane load stabilization [58].
--F2 Dynamics and Physical Parameters
-The continuous-time dynamics of the Pendulum are derived from Newtonian mechanics and are defined by the following equations of motion:
-$\dot{\theta}$ $=\dot{\theta},$
-$\ddot{\theta}$ $=\frac{mgL\sin(\theta)-b\dot{\theta}+u}{mL^{2}},$
-where $\theta$ and $\dot{\theta}$ denote the angular displacement and angular velocity, respectively.
-The physical variables and their definitions are summarized in Table VIII, while the specific parameter values and simulation settings used in this study are provided in Table IX.
-TABLE VIII: Variable Definitions for Pendulum
-Variable Physical Meaning
-$\theta$ Angular displacement of the pendulum (rad)
-$\dot{\theta}$ Angular velocity of the pendulum (rad/s)
-$\ddot{\theta}$ Angular acceleration of the pendulum (rad/s2)
-$m$ Mass of the pendulum bob (kg)
-$g$ Gravitational acceleration (m/s2)
-$L$ Length of the pendulum rod (m)
-[…]
-[…]
-TABLE IX: Physical Parameter Values for Pendulum
-Parameter Value
-$g$ 9.81 m/s2
-$L$ 0.5 m
-$m$ 0.15 kg
-[…]
-Simulation time step ($dt$) 0.01 s
-Control update interval 5 simulation steps
-Maximum episode steps 1000
--F3 State and Action Spaces
-The system state vector is defined as $\mathbf{x}=[\theta,\dot{\theta}]^{\top}$, with the state space constrained to $\mathcal{X}=[-\pi,\pi]\times[-10,10]$.
-This range covers the full angular displacement and accounts for physical angular velocity limits.
-The control objective is to drive the system to the unstable equilibrium point (upright position) defined as $\mathbf{x}_{g}=[0,0]^{\top}$.
-The control action $u$ is a one-dimensional continuous torque restricted to the action space $\mathcal{U}=[-5,5]$.
-These torque bounds are selected to ensure the system has sufficient authority for the swing-up maneuver while remaining within realistic actuation limits.
-At the start of each episode, the initial state $\mathbf{x}_{0}$ is uniformly sampled from the entire state space, $\mathbf{x}_{0}\sim\text{Uniform}([-\pi,\pi]\times[-10,10])$, requiring the controller to handle both stabilization from near-upright positions and complex swing-up maneuvers from high-energy initial states.
--F4 Comparison with OpenAI Gym Pendulum
-Our Pendulum implementation is significantly more rigorous than the standard OpenAI Gym version [40].
-Specifically, we expand the initial state space to the full range $\pi,\pi]\times[-10,10]$ and increase the actuation limit to [5,5][-5,5], whereas the Gym environment utilizes a narrower velocity initialization ([1,1][-1,1]) and limited torque range ([2,2][-2,2]).
-Furthermore, we impose strict boundary-based termination and extend the episode horizon to $T=1000$ steps.
-These more stringent constraints are intentionally designed to expose the limitations of standard RL baselines; for instance, while PPO converges readily in the Gym setting, it often fails to achieve robust stabilization in our high-energy regime.
--G Details of DuctedFan
--G1 System Overview
-The DuctedFan system models the coupled translational and rotational dynamics of a ducted-fan aircraft under the influence of gravitational and control forces [66].
-As a canonical benchmark for Vertical Takeoff and Landing (VTOL) research, it provides a rigorous platform for validating hover stabilization and attitude control algorithms [61, 67].
-The system is underactuated and exhibits strong nonlinear coupling between its lateral displacement and angular orientation.
--G2 Dynamics and Physical Parameters
-The continuous-time dynamics of the DuctedFan are derived from Newtonian mechanics and rigid-body dynamics, defined by the following state-space equations:
-$\dot{x}$ $=\dot{x},$
-$\dot{y}$ $=\dot{y},$
-$\dot{\theta}$ $=\dot{\theta},$
-$\ddot{x}$ $=\frac{-mg\sin(\theta)-d\dot{x}+u_{1}\cos(\theta)-u_{2}\sin(\theta)}{m},$
-$\ddot{y}$ $=\frac{mg(\cos(\theta)-1)-d\dot{y}+u_{1}\sin(\theta)+u_{2}\cos(\theta)}{m},$
-$\ddot{\theta}$ $=\frac{ru_{1}}{J},$
-where $u_{1}=f_{1}$ and $u_{2}=f_{2}-mg$, with $f_{1}$ and $f_{2}$ representing the control forces applied perpendicular and parallel to the fan axis, respectively.
-The variable definitions and physical parameters used for the simulation are summarized in Table X and Table XI.
-TABLE X: Variable Definitions for DuctedFan
-Variable Physical Meaning
-$x,y$ Horizontal and vertical displacement (m)
-$\theta$ Angular orientation (rad)
-$\dot{x},\dot{y}$ Horizontal and vertical velocity (m/s)
-$\dot{\theta}$ Angular velocity (rad/s)
-$m$ System mass (kg)
-$g$ Gravitational acceleration (m/s2)
-$r$ Moment arm of the control force (m)
-[…]
-[…]
-$u_{1},u_{2}$ Control force inputs (N)
-TABLE XI: Physical Parameter Values for DuctedFan
-Parameter Value
-$g$ 9.81 m/s2
-$m$ 8.5 kg
-$r$ 0.26 m
-[…]
-[…]
-Simulation time step ($dt$) 0.01 s
-Control update interval 5 simulation steps
-Maximum episode steps 1000
--G3 State and Action Spaces
-The system state vector is defined as $\mathbf{x}=[x,y,\theta,\dot{x},\dot{y},\dot{\theta}]^{\top}$.
-The admissible state space is $\mathcal{X}=[-5,5]^{2}\times[-\pi/2,\pi/2]\times[-5,5]^{3}$, which covers the translational workspace, the pitch angle range, and the respective velocity limits.
-The control objective is to stabilize the vehicle at the equilibrium state $\mathbf{x}_{g}=[0,0,0,0,0,0]^{\top}$, corresponding to a steady hover at the origin.
-The control vector $\mathbf{u}=[u_{1},u_{2}]^{\top}$ consists of two continuous force inputs, with the action space defined as $\mathcal{U}=[-5,5]\times[-5,5]$.
-These bounds reflect physical actuation limits while ensuring sufficient control authority to counteract gravity and damping.
-At the start of each episode, the initial state $\mathbf{x}_{0}$ is sampled from a uniform distribution $\mathbf{x}_{0}\sim\text{Uniform}([-0.5,0.5]^{6})$.
-This initialization provides a localized yet challenging set of starting conditions that require the controller to coordinate multi-axis forces to achieve precise stabilization.
--H Details of Two-link
--H1 System Overview
-The Two-link system captures the rigid-body rotational dynamics of a serial two-link planar structure subject to gravitational effects and joint torque inputs.
-This system serves as a representative benchmark for underactuated or fully actuated mechanical systems, where the nonlinear coupling between links presents a significant control challenge [59].
-Key applications of this model include robotic manipulator trajectory tracking, humanoid robot joint coordination, and the study of multi-degree-of-freedom Lagrangian systems [68].
--H2 Dynamics and Physical Parameters
-The continuous-time dynamics of the Two-link Planar Robot are derived from Lagrangian mechanics and are expressed by the standard Euler-Lagrange equation:
-$\mathbf{M}(\mathbf{q})\ddot{\mathbf{q}}+\mathbf{C}(\mathbf{q},\dot{\mathbf{q}})\dot{\mathbf{q}}+\mathbf{G}(\mathbf{q})=\mathbf{u},$
-where $\mathbf{q}=[\theta_{1},\theta_{2}]^{\top}$ denotes the vector of joint angles, $\mathbf{M}(\mathbf{q})$ is the $2\times 2$ symmetric positive-definite mass matrix, $\mathbf{C}(\mathbf{q},\dot{\mathbf{q}})$ is the Coriolis and centrifugal matrix, $\mathbf{G}(\mathbf{q})$ is the gravity vector, and $\mathbf{u}=[u_{1},u_{2}]^{\top}$ is the control vector of joint torques.
-The explicit forms of these matrices are given by:
-$\mathbf{M}(\mathbf{q})=\begin{bmatrix}M_{11}&M_{12}\\ M_{21}&M_{22}\in\mathbb{R}^{2\times 2}\end{bmatrix}$,
-$\mathbf{C}(\mathbf{q},\dot{\mathbf{q}})=\begin{bmatrix}-h\dot{\theta}_{2}&-h(\dot{\theta}_{1}+\dot{\theta}_{2})\\ h\dot{\theta}_{1}&0\end{bmatrix}$,
-$\mathbf{G}(\mathbf{q})=\begin{bmatrix}-(m_{1}l_{c1}+m_{2}l_{1})g\sin\theta_{1}-m_{2}l_{c2}g\sin(\theta_{1}+\theta_{2})\\ -m_{2}l_{c2}g\sin(\theta_{1}+\theta_{2})\end{bmatrix}$,
-where $h=m_{2}l_{1}l_{c2}\sin\theta_{2}$. The individual elements of the mass matrix are defined as:
-$M_{11}$ $=I_{1}+I_{2}+m_{1}l_{c1}^{2}+m_{2}(l_{1}^{2}+l_{c2}^{2}+2l_{1}l_{c2}\cos\theta_{2}),$
-$M_{12}$ $=M_{21}=I_{2}+m_{2}(l_{c2}^{2}+l_{1}l_{c2}\cos\theta_{2}),$
-$M_{22}$ $=I_{2}+m_{2}l_{c2}^{2}.$
-The variable definitions and the specific physical parameters used in our simulation are summarized in Table XII and Table XIII, respectively.
-TABLE XII: Variable Definitions for Two-link Planar Robot
-Variable Physical Meaning
-$\theta_{1},\theta_{2}$ Angular orientation of link 1 and link 2 (rad)
-$\dot{\theta}_{1},\dot{\theta}_{2}$ Angular velocity of link 1 and link 2 (rad/s)
-$\ddot{\theta}_{1},\ddot{\theta}_{2}$ Angular acceleration of link 1 and link 2 (rad/s2)
-$l_{1},l_{2}$ Length of link 1 and link 2 (m)
-$m_{1},m_{2}$ Mass of link 1 and link 2 (kg)
-$l_{c1},l_{c2}$ Distance from joint to center of mass (COM) (m)
-[…]
-[…]
-TABLE XIII: Physical Parameter Values for Two-link Planar Robot
-Parameter Value
-$g$ 9.81 m/s2
-$l_{1},l_{2}$ 1.0 m, 1.0 m
-$m_{1},m_{2}$ 1.0 kg, 1.0 kg
-$l_{c1},l_{c2}$ 0.5 m, 0.5 m
-[…]
-Simulation time step ($dt$) 0.01 s
-Control update interval 5 simulation steps
-Maximum episode steps 1000
--H3 State and Action Spaces
-The system state vector is defined as $\mathbf{x}=[\theta_{1},\theta_{2},\dot{\theta}_{1},\dot{\theta}_{2}]^{\top}$.
-To incorporate physical constraints, the state space is bounded by $\mathcal{X}=[-\pi/2,\pi/2]^{2}\times[-20,20]^{2}$, where the joint angles are limited to a $\pm\pi/2$ range and the angular velocities are capped at $\pm 20$ rad/s.
-The control objective is to achieve stabilization at the equilibrium state $\mathbf{x}_{g}=[0,0,0,0]^{\top}$, representing the system at rest in the downward vertical position.
-The control action vector $\mathbf{u}=[u_{1},u_{2}]^{\top}$ consists of continuous joint torques within the action space $\mathcal{U}=[-20,20]\times[-20,20]$.
-These torque limits are established to prevent excessively high control efforts while providing sufficient authority to manage the gravitational and inertial coupling effects.
-For each training episode, the initial state $\mathbf{x}_{0}$ is sampled from a uniform distribution $\mathbf{x}_{0}\sim\text{Uniform}([-0.5,0.5]^{4})$, ensuring that the controller begins each run in a neighborhood of the target equilibrium while requiring active correction to maintain stability.
--I Details of SingleCarTracking
--I1 System Overview
-The SingleCarTracking environment is designed for high-fidelity ground vehicle trajectory .
-It utilizes a single-track (bicycle) model that incorporates both lateral and longitudinal dynamics of a rear-wheel-drive vehicle, accounting for nonlinear tire-road friction [69] and steering constraints.
-This benchmark is crucial for validating control strategies in autonomous driving [60, 70], such as path following and high-precision maneuvers in mobile robotics [10].
--I2 Reference Trajectory and Tracking Error Formulation
-In this study, the reference trajectory focuses on the spatial path in the 2D plane.
-Let $\bm{p}_{t}^{\text{ref}}=[x_{t}^{\text{ref}},y_{t}^{\text{ref}}]^{\top}\in\mathbb{R}^{2}$ denote the target position on the horizontal plane.
-While the full reference state $\mathbf{x}_{t}^{\text{ref}}$ includes heading and velocity, the core objective is defined by the straight-line trajectory $\bm{p}_{t}^{\text{ref}}=[t,0]^{\top}$ for $t\in[0,50]$.
-The reference parameters are set to a longitudinal velocity $v_{\text{ref}}=1.0$ m/s, zero acceleration $a_{\text{ref}}=0.0$ m/s2, and a zero heading angle $\psi_{\text{ref}}=0.0$ with zero angular velocity $\omega_{\text{ref}}=0.0$ rad/s.
-A friction scaling factor $\mu_{\text{scale}}=0.1$ is applied to adjust tire nonlinearity.
-The physical state of the vehicle is described by the vector $\mathbf{x}=[x,y,\delta,v,\psi,\dot{\psi},\beta]^{\top}$, representing position, steering angle, longitudinal speed, heading angle, yaw rate, and side-slip angle.
-For control purposes, we define the tracking error vector as $\mathbf{e}=[s_{xe},s_{ye},\delta_{e},v_{e},\psi_{e},\dot{\psi}_{e},\beta_{e}]^{\top}$, where $s_{xe}=x-x_{\text{ref}}$ and $s_{ye}=y-y_{\text{ref}}$ are the longitudinal and lateral position errors (also denoted as $e_{x}$ and $e_{y}$ respectively in the main text).
-The admissible error space $\mathcal{E}$ is constrained to $1,1]^{2}\times[-1.066,1.066]\times[-1,1]\times[-\pi/2,\pi/2]^{2}\times[-\pi/3,\pi/3]$, reflecting realistic tracking tolerances and physical actuation limits.
-The control objective is to maintain the system at the equilibrium error state $\mathbf{e}_{g}=\mathbf{0}$.
-At the start of each episode, the error is initialized via $\mathbf{e}_{0}\sim\text{Uniform}([-0.5,0.5]^{7})$.
--I3 Dynamics and Physical Parameters
-The system dynamics are modeled as an affine nonlinear system $\dot{\mathbf{e}}=\mathbf{f}(\mathbf{e})+\mathbf{g}(\mathbf{e})\mathbf{u}$, where the control input $\mathbf{u}=[\dot{\delta},a_{\text{long}}]^{\top}$ consists of the steering angular velocity and longitudinal acceleration, bounded by $\mathcal{U}=[-5,5]^{2}$.
-To ensure numerical stability and physical accuracy across different speed regimes, the model transitions between a dynamic tire model and a kinematic bicycle model.
-Dynamic Model ($|v|\geq 0.1\mathrm{m/s}$)
-In the high-speed regime, the drift vector field is defined as $\mathbf{f}(\mathbf{e})=[v\cos(\psi_{e}+\beta_{e})-v_{\text{ref}}+\omega_{\text{ref}}s_{ye},v\sin(\psi_{e}+\beta_{e})-\omega_{\text{ref}}s_{xe},0,-a_{\text{ref}},\dot{\psi}_{e},f_{6}(\mathbf{e}),f_{7}(\mathbf{e})]^{\top}$.
-The nonlinear yaw and side-slip terms are:
-$f_{6}(\mathbf{e})=$ $\frac{\mu m}{I_{z}(l_{f}+l_{r})}\Bigg[-\frac{l_{f}^{2}C_{Sf}gl_{r}+l_{r}^{2}C_{Sr}gl_{f}}{v}\dot{\psi}$
-$+(l_{r}C_{Sr}gl_{f}-l_{f}C_{Sf}gl_{r})\beta_{e}+(l_{f}C_{Sf}gl_{r})\delta_{e}\Bigg],$
-$f_{7}(\mathbf{e})=$ $\left[\frac{\mu(C_{Sr}gl_{f}l_{r}-C_{Sf}gl_{r}l_{f})}{v^{2}(l_{f}+l_{r})}-1\right]\dot{\psi}$
-$-\frac{\mu(C_{Sr}gl_{f}+C_{Sf}gl_{r})}{v(l_{f}+l_{r})}\beta_{e}+\frac{\mu C_{Sf}gl_{r}}{v(l_{f}+l_{r})}\delta_{e}.$
-where
-$C_{Sf}=C_{Sr}=-\frac{\text{tire\_p\_ky1}}{\text{tire\_p\_dy1}}$
-The control gain matrix is $\mathbf{g}(\mathbf{e})=[\mathbf{0}_{2\times 2},[1,0],[0,1],\mathbf{0},[0,\mathbf{g}_{6,2}(\mathbf{e})],[0,\mathbf{g}_{7,2}(\mathbf{e})]]^{\top}$.
-The terms $g_{6,2}$ and $g_{7,2}$ account for CG height $h_{s}$ and vertical load shifts:
-$g_{6,2}(\mathbf{e})=$ $\frac{\mu m}{I_{z}(l_{f}+l_{r})}\Bigg[-\frac{-l_{f}^{2}C_{Sf}h_{s}+l_{r}^{2}C_{Sr}h_{s}}{v}\dot{\psi}$
-$+(l_{r}C_{Sr}h_{s}+l_{f}C_{Sf}h_{s})\beta_{e}-(l_{f}C_{Sf}h_{s})\delta_{e}\Bigg],$
-$g_{7,2}(\mathbf{e})=$ $\frac{\mu(C_{Sr}h_{s}l_{r}+C_{Sf}h_{s}l_{f})}{v^{2}(l_{f}+l_{r})}\dot{\psi}$
-$-\frac{\mu(C_{Sr}h_{s}-C_{Sf}h_{s})}{v(l_{f}+l_{r})}\beta_{e}-\frac{\mu C_{Sf}h_{s}}{v(l_{f}+l_{r})}\delta_{e}.$
-Kinematic Model ($|v|<0.1\mathrm{m/s}$)
-To avoid singularities at low speeds, the model simplifies to a kinematic representation where side-slip is negligible.
-Let $L=l_{f}+l_{r}$ denote the wheelbase.
-The drift field $\mathbf{f}(\mathbf{e})$ and input matrix $\mathbf{g}(\mathbf{e})$ are expressed as:
-$\mathbf{f}(\mathbf{e})=\Big[$ $v\cos(\psi_{e}+\beta_{e})-v_{\text{ref}}+\omega_{\text{ref}}s_{ye},$
-$v\sin(\psi_{e}+\beta_{e})-\omega_{\text{ref}}s_{xe},$
-$0,\qquad-a_{\text{ref}},\qquad\frac{v\cos(\beta_{e})}{L}\tan(\delta_{e})-\omega_{\text{ref}},\qquad 0,\qquad 0\Big]^{\top},$
-$\mathbf{g}(\mathbf{e})=\Big[$ $\mathbf{0}_{2\times 2},\qquad[1,0],\qquad[0,1],\qquad\mathbf{0},$
-$\frac{\cos(\beta_{e})\tan(\delta_{e})}{L},\qquad\beta_{e}^{\prime}\Big]^{\top},$
-where the side-slip rate auxiliary term is
-$\beta_{e}^{\prime}=\frac{l_{r}/(L\cos^{2}(\delta_{e}))}{1+\left(\tan(\delta_{e})\frac{l_{r}}{L}\right)^{2}}.$
-The system parameters are detailed in Tables XIV and XV.
-TABLE XIV: Variable Definitions for SingleCarTracking
-Category Variable Physical Meaning
-True State $\mathbf{x}\in\mathbb{R}^{7}$ $\mathbf{x}=[x,y,\delta,v,\psi,\dot{\psi},\beta]^{\top}$: Position ($x,y$), steering angle, speed, heading, yaw rate, side-slip.
-Tracking Error $\mathbf{e}\in\mathbb{R}^{7}$ $\mathbf{e}=[s_{xe},s_{ye},\delta_{e},v_{e},\psi_{e},\dot{\psi}_{e},\beta_{e}]^{\top}$: Tracking errors relative to reference trajectory.
-Control $\mathbf{u}\in\mathbb{R}^{2}$ $\mathbf{u}=[\dot{\delta},a_{\text{long}}]^{\top}$: Steering rate (rad/s) and longitudinal acceleration (m/s2).
-Parameters $m,I_{z},l_{f},l_{r},h_{s}$ Mass, yaw inertia, distance to front/rear axles, and CG height.
-$\mu,K_{\text{tire}}$ Tire-road friction coefficient and tire stiffness parameters.
-TABLE XV: Physical Parameter Values for SingleCarTracking
-Parameter Value Parameter Value
-[…]
-$l_{f},l_{r}$ 1.155 m, 1.423 m $h_{s}$ 0.614 m
-$\delta_{\text{max}}$ $\pm 1.066$ rad $a_{\text{long, max}}$ $\pm 5$ m/s2
-tire_p_dy1 1.0489 tire_p_ky1 -21.92
-dtdt 0.01 s Horizon $T$ 1000
--J Details of QuadrotorTracking
--J1 System Overview
-The QuadrotorTracking environment evaluates the 6-degree-of-freedom (6-DoF) rigid-body dynamics of a quadrotor UAV [71].
-This benchmark incorporates complex rotor thrust-torque coupling and attitude kinematics defined on the special Euclidean group $\text{SE}(3)$ [54].
-It serves as a standard testbed for geometric control and reinforcement learning algorithms in high-agility scenarios[72], such as autonomous aerial inspection and precise path following [73, 74].
--J2 Reference Trajectory and Tracking Error Formulation
-The control objective is to track a time-varying reference trajectory that specifies position, velocity, and heading.
-Let $\bm{p}_{t}^{\text{ref}}=[x_{t}^{\text{ref}},y_{t}^{\text{ref}},z_{t}^{\text{ref}}]^{\top}$ denote the desired 3D position at time step $t$.
-The benchmark utilizes a dynamic helical trajectory defined by:
-$\bm{p}_{t}^{\text{ref}}$ $=[0.4t,0.4\sin(t),0.6\cos(t)]^{\top},$
-$\bm{v}_{t}^{\text{ref}}$ $=[0.4,0.4\cos(t),-0.6\sin(t)]^{\top},$
-where $\bm{v}_{t}^{\text{ref}}$ is the reference velocity. The desired heading is determined by a rotating unit vector $\bm{b}_{1,t}^{\text{ref}}=[\cos(t),\sin(t),0]^{\top}$.
-The system state is $\mathbf{x}=[\mathbf{x}_{p}^{\top},\mathbf{v}^{\top},\mathbf{R},\bm{\Omega}^{\top}]^{\top}$, where $\mathbf{x}_{p}\in\mathbb{R}^{3}$ is the position, $\mathbf{v}\in\mathbb{R}^{3}$ is the translational velocity, $\mathbf{R}\in\text{SO}(3)$ is the rotation matrix, and $\bm{\Omega}\in\mathbb{R}^{3}$ is the body-frame angular velocity.
-Although $\mathbf{R}$ is a $3\times 3$ matrix, it fundamentally represents three degrees of freedom (3-DoF).
-Thus for the sake of notational consistency, with a slight abuse of notation, we treat $\mathbf{R}$ as a three-dimensional vector, thereby concatenating it with $\mathbf{x}_{p}^{\top}$, $\mathbf{v}^{\top}$, and $\bm{\Omega}^{\top}$ to constitute the system state $\mathbf{x}$.
-The tracking error $\mathbf{e}\in\mathbb{R}^{12}$ is defined by four components: position error $\mathbf{e}_{p}=\mathbf{x}_{p}-\bm{p}_{t}^{\text{ref}}$, velocity error $\mathbf{e}_{v}=\mathbf{v}-\bm{v}_{t}^{\text{ref}}$, attitude error $\mathbf{e}_{R}$, and angular velocity error $\mathbf{e}_{\Omega}$.
-Following [54], the attitude error is computed via the skew-symmetric matrix $\mathbf{S}=\mathbf{R}_{t}^{\text{ref}\top}\mathbf{R}-\mathbf{R}^{\top}\mathbf{R}_{t}^{\text{ref}}$, where $\mathbf{e}_{R}=0.5[\mathbf{S}_{2,1},\mathbf{S}_{0,2},\mathbf{S}_{1,0}]^{\top}$, and $\mathbf{S}_{i,j}$ denotes the $(i,j)$-th element of the $\mathbf{S}$ matrix.
-The angular velocity error is $\mathbf{e}_{\Omega}=\bm{\Omega}-\mathbf{R}^{\top}\mathbf{R}_{t}^{\text{ref}}\bm{\Omega}_{t}^{\text{ref}}$.
-The composite error vector $\mathbf{e}=[\mathbf{e}_{p}^{\top},\mathbf{e}_{v}^{\top},\mathbf{e}_{R}^{\top},\mathbf{e}_{\Omega}^{\top}]^{\top}$ is used as the input for the controller.
-At the start of each episode, the initial state $\mathbf{x}_{0}$ is uniformly randomly sampled to calculate $\mathbf{e}_{0}$, specifically
-$\left\{\begin{array}[]{l}\mathbf{x}_{p}^{\top}\sim\text{Uniform}\left([-0.01,0.01]^{3}\right),\\ \mathbf{v}^{\top}\sim\text{Uniform}\left([-0.01,0.01]^{3}\right),\\ \text{vec}(\mathbf{R}-\mathbf{I}_{3})\sim\text{Uniform}\left([-0.01,0.01]^{9}\right),\mathbf{R}\text{ is orthogonal},\\ \bm{\Omega}^{\top}\sim\text{Uniform}\left([-0.01,0.01]^{3}\right).\end{array}\right$.
--J3 Dynamics and Physical Parameters
-The continuous-time dynamics follow the Newton-Euler equations for rigid bodies:
-$\begin{cases}\dot{\mathbf{x}}_{p}=\mathbf{v},&\dot{\mathbf{R}}=\mathbf{R}[\bm{\Omega}]_{\times},\\ \dot{\mathbf{v}}=-\mathbf{g}+\frac{F}{m}\mathbf{b}_{3},&\dot{\bm{\Omega}}=\mathbf{J}^{-1}(\mathbf{M}-\bm{\Omega}\times\mathbf{J}\bm{\Omega}),\end{cases}$
-[…]
-$\begin{bmatrix}F\\ M_{x}\\ M_{y}\\ M_{z}\end{bmatrix}=\begin{bmatrix}1&1&1&1\\ 0&-d&0&d\\ d&0&-d&0\\ -c_{tf}&c_{tf}&-c_{tf}&c_{tf}\end{bmatrix}\begin{bmatrix}f_{1}\\ f_{2}\\ f_{3}\\ f_{4}\end{bmatrix}$
-where $d$ is the moment arm and $c_{tf}$ is the thrust-torque coefficient. Detailed parameter values are provided in Tables XVI and XVII.
-TABLE XVI: Variable Definitions for QuadrotorTracking
-Category Variable Physical Meaning
-True State $\mathbf{x}=[\mathbf{x}_{p}^{\top},\mathbf{v}^{\top},\mathbf{R},\bm{\Omega}^{\top}]^{\top}$ 3D position, velocity, rotation matrix elements, and angular velocity.
-Tracking Error $\mathbf{e}=[\mathbf{e}_{p}^{\top},\mathbf{e}_{v}^{\top},\mathbf{e}_{R}^{\top},\mathbf{e}_{\Omega}^{\top}]^{\top}$ Geometric tracking errors in position, velocity, attitude, and angular rates.
-[…]
-Parameters $m,\mathbf{J},d,c_{tf}$ Mass, inertia matrix, rotor distance, and torque coefficient.
-TABLE XVII: Physical Parameter Values for QuadrotorTracking
-Parameter Value Parameter Value
-$m$ 4.34 kg $\mathbf{g}$ $0,0,9.8]^{\top}$
-$\mathbf{J}$ $\text{diag}(0.08,0.09,0.14)$ $d$ 0.315 m
-[…]
-dtdt 0.01 s $T$ 1000
+This article presents MSACL, a novel Lyapunov-based RL algorithm that integrates exponential stability into MERL and provides a framework for developing verifiable, learning-based policies in safety-critical applications.
+By embedding Lyapunov theory within the SAC architecture, MSACL achieves stable control for model-free systems while preserving maximum entropy exploration benefits.
+Comprehensive evaluation demonstrates consistent superiority of MSACL over established baselines in both performance and stability metrics, with remarkable robustness and generalization capabilities.
+Future work will explore extensions to partially observable systems, adaptive Lyapunov architectures, and cross-domain transfer learning for complex dynamical environments.

### v3

@@
 Abstract
-For safety-critical applications, model-free reinforcement learning (RL) faces numerous challenges, particularly the difficulty of establishing verifiable stability guarantees while maintaining high exploration efficiency.
-To address these challenges, we present Multi-Step Actor-Critic Learning with Lyapunov Certificates (MSACL), a novel approach that seamlessly integrates exponential stability with maximum entropy reinforcement learning (MERL).
-In contrast to existing methods that rely on complex reward engineering and single-step constraints, MSACL utilizes intuitive rewards and multi-step data for actor-critic learning.
-Specifically, we first introduce Exponential Stability Labels (ESLs) to categorize samples and propose a $\lambda$-weighted aggregation mechanism to learn Lyapunov certificates.
-Leveraging these certificates, we then develop a stability-aware advantage function to guide policy optimization, thereby ensuring rapid Lyapunov descent and robust state convergence.
-We evaluate MSACL across six benchmarks, comprising four stabilization and two high-dimensional tracking tasks.
-Experimental results demonstrate its consistent superiority over both standard RL baselines and state-of-the-art Lyapunov-based RL algorithms.
-Beyond rapid convergence, MSACL exhibits significant robustness against environmental uncertainties and remarkable generalization to unseen reference signals.
+For stabilizing control tasks, model-free reinforcement learning (RL) approaches face numerous challenges, particularly regarding the issues of effectiveness and efficiency in complex high-dimensional environments with limited training data.
+To address these challenges, we propose Multi-Step Actor-Critic Learning with Lyapunov Certificates (MSACL), a novel approach that integrates exponential stability into off-policy maximum entropy reinforcement learning (MERL).
+In contrast to existing RL-based approaches that depend on elaborate reward engineering and single-step constraints, MSACL adopts intuitive reward design and exploits multi-step samples to enable exploratory actor-critic learning.
+Specifically, we first introduce Exponential Stability Labels (ESLs) to categorize training samples and propose a $\lambda$-weighted aggregation mechanism to learn Lyapunov certificates.
+Based on these certificates, we further design a stability-aware advantage function to guide policy optimization, thereby promoting rapid Lyapunov descent and robust state convergence.
+We evaluate MSACL across six benchmarks, comprising four stabilizing and two high-dimensional tracking tasks.
+Experimental results demonstrate its consistent performance improvements over both standard RL baselines and state-of-the-art Lyapunov-based RL algorithms.
+Beyond rapid convergence, MSACL exhibits robustness against environmental uncertainties and generalization to unseen reference signals.
 The source code and benchmarking environments are available at https://github.com/YuanZhe-Xing/MSACL.
 I Introduction
-Reinforcement learning (RL) has emerged as a transformative paradigm for the control of complex, nonlinear dynamical systems [1, 2], demonstrating remarkable empirical success across high-dimensional tasks without requiring explicit analytical models [3, 4].
-Despite these advancements, the deployment of RL in safety-critical applications, such as autonomous driving and human-robot interaction, remains fundamentally constrained by a lack of formal reliability guarantees [5, 6].
-While alternative frameworks like constrained Markov Decision Processes (cMDPs) [7, 8] and Hamilton-Jacobi reachability analysis [9, 10] offer potential solutions, certificate-based methods constitute a fundamental framework for verifying and stabilizing learning agents.
-By leveraging rigorous control-theoretic tools, such as Lyapunov certificates for establishing asymptotic stability [11, 12, 14, 13] and barrier certificates for ensuring set invariance [15, 16], these methods provide concise, data-driven proofs of system behavior.
-The synthesis of stability certificates has undergone a paradigm shift from computationally intensive analytical methods toward scalable neural architectures.
-Traditional approaches, such as sum-of-squares (SOS) optimization [17, 18, 19] and algebraic synthesis[20], are often restricted to polynomial dynamics.
-To bridge the gap between black-box learning and verifiable control, the field has evolved from probabilistic methods, notably the use of Gaussian Processes for safe exploration [21], to formal frameworks for synthesizing neural certificates that estimate regions of attraction (ROA) [22, 23].
-Building on this verification-centric perspective, recent research has advanced toward the joint synthesis of neural certificates and policies.
-For instance, Chang et al.
-[24, 25] and Quan et al.
-[26, 27, 28] demonstrated that neural certificates can function as critics during training, directly guiding policy optimization to satisfy formal stability constraints.
-In contrast to the stability-by-design paradigms presented above, standard model-free RL algorithms, such as DDPG [29], PPO [31], TD3 [32], SAC [33], and SAC’s distributional variant DSAC [34, 35], lack intrinsic stability guarantees, thus facing fundamental limitations in precision control tasks.
-These methods treat stability as an incidental byproduct of reward maximization rather than a structural property.
-Consequently, employing general RL algorithms for safety-critical tasks necessitates extensive reward engineering to achieve stable behaviors [1, 2].
+Stabilizing control is a central requirement in real-world control systems, especially in safety-critical applications such as autonomous driving [1], aerospace systems [2], and robotics [3].
+In these settings, a controller must not only accomplish the desired task but also guarantee stable system behavior under uncertainty and disturbances [4, 5].
+As control systems become increasingly complex and are deployed in dynamic, partially unknown environments, developing control strategies that can simultaneously ensure stability and scalability has become a fundamental challenge [6].
+Traditionally, control strategies have relied heavily on rigorous mathematical frameworks and precise analytical models[7, 8], such as constrained Markov Decision Processes (cMDPs) [9], Hamilton-Jacobi reachability analysis [10] and certificate-based methods utilizing strict control-theoretic tools (e.g., sum-of-squares optimization [11, 12] and algebraic synthesis [13]).
+However, all these traditional approaches face severe scalability challenges when applied to complex nonlinear systems with unknown dynamics.
+To address the complexities of real-world dynamical systems without requiring explicit analytical models, RL has emerged as a highly effective paradigm for optimal control [14], demonstrating considerable empirical success across various control tasks [15, 16].
+Nevertheless, standard model-free RL algorithms, such as DDPG [17], PPO [18], and SAC [19], lack intrinsic stability awareness.
+They treat stability as an incidental byproduct of reward maximization, necessitating extensive reward engineering to achieve stable or safe behaviors.
 To bridge this gap, Lyapunov-based RL algorithms have been developed to integrate stability constraints into the learning process.
-For instance, Policy Optimization with Lyapunov Certificates (POLYC) [25] incorporates Lyapunov certificates within the PPO framework, while Lyapunov-based Actor-Critic (LAC) [36] introduces an off-policy approach for learning stability-constrained critics.
-By augmenting standard RL architectures with learned Lyapunov certificates, these methods provide structural stability guarantees for nonlinear control systems.
-However, several critical limitations persist.
-First, these approaches suffer from low sample efficiency under on-policy training paradigms.
-Second, they often fail to account for control costs during policy optimization.
-More critically, most existing methods focus solely on asymptotic stability, which lacks the rapid stabilization characteristics necessary for high-performance response in dynamic control scenarios.
-To overcome these limitations, we propose Multi-Step Actor-Critic Learning with Lyapunov Certificates (MSACL), a Lyapunov-based off-policy RL algorithm that enhances sample efficiency [37, 38] while optimizing a comprehensive objective that balances state regulation and control cost.
-By integrating exponential stability constraints into the maximum entropy reinforcement learning (MERL) architecture, MSACL ensures rapid stabilization and enhanced robustness.
-A key feature of our approach is the multi-step learning mechanism, which captures long-horizon dynamics to establish connections between temporal difference learning and formal stability verification.
-By mitigating discretization errors and stochastic noise associated with single-step transitions, this mechanism enables the synthesis of policies that are both high-performing and mathematically certifiable.
-In summary, this article presents a Lyapunov-based RL approach for model-free exponentially stabilizing control. The main contributions are summarized as follows:
+For instance, Policy Optimization with Lyapunov Certificates (POLYC) [20] incorporates Lyapunov certificates within the PPO framework, while Lyapunov-based Actor-Critic (LAC) [21] introduces an approach for learning stability-constrained critics.
+By augmenting standard RL architectures with learned neural Lyapunov certificates, these methods empirically facilitate structural stability for nonlinear control systems.
+However, despite these advances, existing Lyapunov-based RL methods often struggle with both effectiveness and efficiency in complex, high-dimensional stabilizing control tasks (e.g., Quadrotor tracking).
+The primary bottleneck for effectiveness lies in their inability to learn sufficiently accurate Lyapunov certificates.
+First, existing methods rely on single-step constraints, providing limited long-horizon information and leading to myopic policy improvement.
+Second, on-policy methods like POLYC depend exclusively on newly collected data, restricting the diversity of behaviors needed to accurately shape the Lyapunov landscape.
+Consequently, the learned certificates fail to provide reliable structural guidance, causing performance degradation or divergence.
+To overcome the effectiveness issue, introducing a multi-step off-policy approach is essential, as it captures long-horizon trajectory evolution.
+Meanwhile, the multi-step off-policy approach naturally resolves the secondary issue: it boosts efficiency by deeply reusing historical experiences, overcoming the low sample efficiency inherent in previous on-policy methods.
+Inspired by the idea of using multi-step off-policy, we in this paper attempt to learn a high-fidelity Lyapunov certificate that can reliably guide policy updates.
+To this end, we propose Multi-Step Actor-Critic Learning with Lyapunov Certificates (MSACL).
+First, MSACL exploits multi-step samples to learn Lyapunov certificates, enabling the certificates to capture long-horizon system dynamics rather than relying only on local single-step transitions.
+Second, MSACL is built upon an off-policy maximum entropy reinforcement learning (MERL) framework, which not only improves sample efficiency through data reuse but also promotes broader exploration, thereby providing more diverse training data for characterizing system behaviors.
+While parameterizing Lyapunov certificates via neural networks inherently lacks the strict formal mathematical guarantees of traditional analytical methods, the proposed multi-step and off-policy learning paradigm yields a more reliable approximation empirically.
+This enhanced Lyapunov certificate, in turn, provides structurally grounded, stability-aware advantage estimates to guide policy optimization toward rapid and robust stabilization.
+The proposed MSACL approach makes the following contributions to safe model-free control:
   1. (i)
-    Exponential Stability via Multi-Step Learning.
-We propose the MSACL algorithm, which integrates exponential stability with Lyapunov certificate learning via a multi-step mechanism.
-This algorithm enforces exponential convergence rates, providing stability-guaranteed performance for complex nonlinear control, surpassing conventional asymptotic methods.
+    Multi-Step Learning with Lyapunov Guidance.
+We introduce a multi-step learning mechanism to learn high-precision Lyapunov certificates that better capture long-horizon system dynamics and provide more effective guidance for policy updates.
+Built upon this mechanism, MSACL further incorporates exponential stability through Exponential Stability Labels (ESLs) and a $\lambda$-weighted aggregation scheme, yielding stronger stability supervision and faster convergence than conventional single-step methods.
   2. (ii)
     Sample-Efficient Off-Policy Framework.
-We develop a sample-efficient off-policy framework that leverages the maximum entropy paradigm and importance sampling.
-This framework effectively manages distribution shifts and interaction data reuse, thus enhancing training stability compared to existing on-policy algorithms.
+We develop MSACL within an off-policy MERL framework, moving beyond the data inefficiency of existing on-policy Lyapunov-based methods.
+By improving data reuse and encouraging broader exploration, the proposed framework captures more diverse system behaviors and provides richer training samples for Lyapunov certificate learning.
+In addition, importance sampling is incorporated to mitigate distribution shifts, further enhancing sample efficiency and training stability.
   3. (iii)
-    Systematic Evaluation and Robustness Analysis.
-We establish an evaluation platform across six benchmarks comprising stabilization and high-dimensional tracking tasks.
-Experiment results demonstrate superior performance, robustness and generalization of MSACL, while sensitivity analyses elucidate the multi-step bias-variance trade-off to guide diverse control tasks.
+    Systematic Evaluation and Robustness Validation.
+We establish a comprehensive evaluation platform across six benchmarks, including high-dimensional stabilizing control tasks.
+Experimental results demonstrate MSACL’s consistent empirical superiority over both standard RL and state-of-the-art Lyapunov-based baselines.
+Furthermore, MSACL exhibits robustness against environmental noise and generalization to unseen reference signals.
 The remainder of this article is organized as follows.
@@
 This section establishes the theoretical foundations for MSACL.
-We first define the stabilization and tracking problems for model-free discrete-time systems, then introduce Lyapunov-based exponential stability criteria, and finally review the MERL framework.
+We first define the exponentially stabilizing control problems for model-free discrete-time systems, then introduce Lyapunov-based exponential stability criteria, and finally review the MERL framework.
 II-A Problem Statement
-In this article, we study the stabilization problem for general discrete-time deterministic control systems:
+In this article, we study the stabilizing problem for general discrete-time deterministic control systems:
 $\mathbf{x}_{t+1}=f(\mathbf{x}_{t},\mathbf{u}_{t}),$ (1)
-where $\mathbf{x}_{t}\in\mathcal{X}\subseteq\mathbb{R}^{d}$ denotes the state at time $t$, $\mathbf{u}_{t}\in\mathcal{U}\subseteq\mathbb{R}^{m}$ denotes the control input, $\mathcal{X}$ and $\mathcal{U}$ represent the sets of admissible states and control inputs, respectively, and $f:\mathcal{X}\times\mathcal{U}\to\mathcal{X}$ is the flow map.
+where $\mathbf{x}_{t}\in\mathcal{X}\subseteq\mathbb{R}^{d}$ denotes the state at time $t$, $\mathbf{u}_{t}\in\mathcal{U}\subseteq\mathbb{R}^{m}$ denotes the control input, $\mathcal{X}$ and $\mathcal{U}$ represent the sets of admissible states and control inputs respectively, and $f:\mathcal{X}\times\mathcal{U}\to\mathcal{X}$ is the flow map.
 Given that $f$ is unknown, we adopt a model-free control paradigm.
@@
 It is worth noting that the discrete-time case presented in this article can be similarly extended to the continuous-time case; interested readers may refer to [22] and [20] for details.
-In the context of control systems, our objective is to find a state feedback policy $\pi:\mathcal{X}\to\mathcal{U}$ such that the control input $\mathbf{u}_{t}=\pi(\mathbf{x}_{t})$ renders the closed-loop system $\mathbf{x}_{t+1}=f(\mathbf{x}_{t},\pi(\mathbf{x}_{t}))$ to satisfy exponential stability.
+For the control system (1), our objective is to find a state feedback policy $\pi:\mathcal{X}\to\mathcal{U}$ such that the control input $\mathbf{u}_{t}=\pi(\mathbf{x}_{t})$ renders the closed-loop system $\mathbf{x}_{t+1}=f(\mathbf{x}_{t},\pi(\mathbf{x}_{t}))$ to satisfy exponential stability.
 For clarity, we consider the scenario where the state $\mathbf{x}_{t}$ is fully observable, and the policy $\pi$ is stationary, i.e., it depends only on the state and not explicitly on time.
@@
 Problem 1
-(Exponential Stabilization) For the control system (1), design a stationary policy $\pi$ such that for any initial state $\mathbf{x}_{0}\in\mathcal{X}_{0}$, the resulting closed-loop trajectory $\mathbf{x}(t,\mathbf{x}_{0})$ exponentially converges to the equilibrium state $\mathbf{x}_{g}$.
+(Exponential stabilizing control) For the control system (1), design a policy $\pi$ such that for any initial state $\mathbf{x}_{0}\in\mathcal{X}_{0}$, the resulting closed-loop trajectory $\mathbf{x}(t,\mathbf{x}_{0})$ exponentially converges to the equilibrium state $\mathbf{x}_{g}$.
 Without loss of generality, let $\mathbf{x}_{g}$ denote the origin, i.e., $\mathbf{x}_{g}=\mathbf{0}$.
-Furthermore, the tracking problem is inherently encompassed within the stabilization framework, as tracking a time-varying reference signal $\mathbf{x}_{t}^{\text{ref}}$ reduces to stabilizing the error $\mathbf{e}_{t}=\mathbf{x}_{t}-\mathbf{x}_{t}^{\text{ref}}$ at zero.
+Furthermore, the tracking problem is inherently encompassed within the stabilizing framework, as tracking a time-varying reference signal $\mathbf{x}_{t}^{\text{ref}}$ reduces to stabilizing the error $\mathbf{e}_{t}=\mathbf{x}_{t}-\mathbf{x}_{t}^{\text{ref}}$ at zero.
 II-B Exponential Stability and Lyapunov Certificates
@@
 $\|\mathbf{x}_{t}-\mathbf{x}_{g}\|\leq C\eta^{t}\|\mathbf{x}_{0}-\mathbf{x}_{g}\|,\quad\forall\mathbf{x}_{0}\in\mathcal{X}_{0},\forall t\geq 0.$ (2)
-To guarantee that the closed-loop system is exponentially stable with respect to $\mathbf{x}_{g}$, we introduce the following lemma:
+To facilitate that the closed-loop system is exponentially stable with respect to $\mathbf{x}_{g}$, we introduce the following lemma:
 Lemma 1
@@
 This reward is formulated to penalize state deviations from the target $\mathbf{x}_{g}$ (or reference signal $\mathbf{x}_{t}^{\text{ref}}$) while accounting for the associated control effort.
-Furthermore, we employ stochastic policies in this article to facilitate exploration within the maximum entropy RL (MERL) framework.
+Furthermore, we employ stochastic policies in this article to facilitate exploration within the MERL framework.
 A stochastic policy is defined as a mapping $\pi:\mathcal{X}\to\mathcal{P}(\mathcal{U})$, where $\mathcal{P}(\mathcal{U})$ denotes the set of probability distributions over $\mathcal{U}$, and $\mathbf{u}_{t}\sim\pi(\cdot|\mathbf{x}_{t})$ is sampled from the probability distribution.
@@
 III-A Data Collection
-In contrast to conventional Lyapunov-based RL algorithms [25, 36], which primarily target asymptotic stability through single-step transitions, our proposed MSACL enforces the stricter exponential stability requirement.
+In contrast to conventional Lyapunov-based RL algorithms [20, 21], which primarily target asymptotic stability through single-step transitions, our proposed MSACL enforces the stricter exponential stability requirements.
 Guided by the multi-step relationship (4), establishing exponential stability necessitates leveraging multi-step sequence samples from state trajectories.
-This sequential data structure provides a sufficient temporal horizon to validate the contraction properties of Lyapunov certificates.
+This sequential data structure provides a sufficient temporal horizon to effectively learn the contraction properties of Lyapunov certificates.
 To operationalize this, we define the fundamental transition tuple as $\mathbf{d}_{t}=(\mathbf{x}_{t},\mathbf{u}_{t},r_{t},\pi_{\phi}(\mathbf{u}_{t}|\mathbf{x}_{t}),\mathbf{x}_{t+1})$.
@@
 On the other hand, learning Lyapunov certificates directly from off-policy data may not be accurate, as empirical sequences in the early stages of training often fail to satisfy the exponential decay condition (5), leading to unreliable certificate estimation.
-To establish a rigorous correspondence between empirical data and theoretical stability requirements, we propose the Exponential Stability Label (ESL) mechanism.
+To establish a strong empirical alignment between empirical data and theoretical stability requirements, we propose the Exponential Stability Label (ESL) mechanism.
 Specifically, we introduce positive and negative samples to bridge the gap between theory and practice:
@@
 IV-B Loss Function Design of Policy Network
-This subsection details the policy network optimization, which seamlessly integrates the maximum entropy objective with a novel Lyapunov-based stability guidance mechanism.
+This subsection details the policy network optimization, which integrates the maximum entropy objective with a novel Lyapunov-based stability guidance mechanism.
 (1) Soft-Q Value Maximization.
@@
 Second, we implement delayed updates for both policy $\pi_{\phi}$ and temperature $\alpha$, performing $d$ consecutive updates within each delayed cycle to ensure sufficient policy improvement.
-These techniques, adapted from the cleanrl framework [42], significantly improve training stability and performance.
+These techniques, adapted from the cleanrl framework [31], improve training stability and performance.
 V Experiments
-This section evaluates the MSACL algorithm across six benchmarking environments, covering stabilization and high-dimensional tracking tasks.
+This section evaluates the MSACL algorithm across six benchmarking environments, covering stabilizing and high-dimensional tracking tasks.
 We compare our approach against model-free and Lyapunov-based RL baselines using performance and stability metrics.
@@
 This subsection presents the benchmarking environments for evaluating MSACL and baseline algorithms.
-We categorize environments by control objective: stabilization tasks include Controlled VanderPol (VanderPol)[43], Pendulum[44], Planar DuctedFan (DuctedFan)[45], and Two-link planar robot (Two-link)[46]; tracking tasks comprise SingleCarTracking[47] and QuadrotorTracking[48], as illustrated in Figure 3.
-For stabilization tasks, the goal is to drive the state to the origin $\mathbf{x}_{g}=\mathbf{0}\in\mathbb{R}^{d}$; for tracking tasks, the system follows a time-varying reference signal $\mathbf{x}_{t}^{\text{ref}}$.
+We categorize environments by control objective: stabilizing tasks include Controlled VanderPol (VanderPol) [32], Pendulum [33], Planar DuctedFan (DuctedFan) [34], and Two-link planar robot (Two-link) [35]; tracking tasks comprise SingleCarTracking [36] and QuadrotorTracking [37], as illustrated in Figure 3.
+For stabilizing tasks, the goal is to drive the state to the origin $\mathbf{x}_{g}=\mathbf{0}\in\mathbb{R}^{d}$; for tracking tasks, the system follows a time-varying reference signal $\mathbf{x}_{t}^{\text{ref}}$.
 All tasks use a unified reward structure:
@@
 We evaluate MSACL against classic model-free RL baselines (SAC [19], PPO [18]) and state-of-the-art Lyapunov-based algorithms (LAC [21], POLYC [20]).
-SAC and PPO serve as performance baselines for stabilization and tracking tasks, while LAC and POLYC provide stability guarantees through Lyapunov constraints.
+SAC and PPO serve as performance baselines for stabilizing and tracking tasks, while LAC and POLYC facilitate stability through Lyapunov constraints.
 For LAC, the standard reward is replaced with a quadratic cost function defined as:
@@
 Note that in contrast to this state-only penalty formulation, our reward $r_{t}$ accounts for both state deviations and control effort.
-By comparing against these baselines, we aim to demonstrate the advantages of MSACL in terms of performance and stability metrics.
+By comparing against these baselines, we aim to demonstrate the advantages of MSACL in terms of performance and stability metrics introduced in the next subsection.
 For fair comparison, all algorithms are implemented in the GOPS framework [40].
@@
 All actors employ stochastic diagonal Gaussian policies.
-Critic architectures vary: LAC uses only a Lyapunov network, while POLYC incorporates both value and Lyapunov functions.
+Critic architectures vary: LAC uses only a Lyapunov network, while POLYC incorporates both value and Lyapunov networks.
 SAC and LAC also implement clipped double Q-learning and delayed policy updates according to Remark 3.
@@
 Specifically, MSACL consistently achieves the highest AMCR across all six benchmarks.
-Regarding AMCC, MSACL attains the absolute minimum in the Pendulum, SingleCarTracking, and QuadrotorTracking tasks, while remaining competitive in others.
+Regarding AMCC, MSACL attains the minimum cost among all evaluated methods in the Pendulum, SingleCarTracking, and QuadrotorTracking tasks, while remaining competitive in others.
 Furthermore, MSACL maintains the highest RR across all radii.
@@
 As shown in Fig.
-4 and Table II, the integration of the Lyapunov certificate significantly enhances both training performance and stability metrics compared to the standard SAC.
-This validates that embedding Lyapunov-based constraints into MERL effectively yields neural policies with superior stability guarantees for model-free control systems.
+4 and Table II, the integration of the Lyapunov certificate enhances both training performance and stability metrics compared to the standard SAC.
+This validates that embedding Lyapunov-based constraints into MERL effectively yields neural policies with stability properties for model-free control systems.
 Figure 5: Visualization of learned Lyapunov certificates and their corresponding contour plots across six benchmarks: (a) VanderPol; (b) Pendulum; (c)–(e) DuctedFan (projections in the $(x,y)$-plane with $\theta=0$, $(y,\theta)$-plane with $x=0$, and $(x,\theta)$-plane with $y=0$); (f) Two-link; (g) SingleCarTracking; (h)–(j) QuadrotorTracking (projections in the $(e_{x},e_{y})$-plane with $e_{z}=0$, $(e_{y},e_{z})$-plane with $e_{x}=0$, and $(e_{x},e_{z})$-plane with $e_{y}=0$).
@@
 This subsection evaluates the robustness and generalization of MSACL-trained policies, with the aim of demonstrating their real-world applicability under inherent uncertainties.
-For stabilization tasks (VanderPol, Pendulum, DuctedFan, Two-link), we introduce process noise $\bm{\epsilon}_{t}\sim\mathcal{N}(\mathbf{0},\sigma_{\text{env}}\cdot\mathbf{1}_{d})$ to the system dynamics, where $\mathbf{1}_{d}$ denotes the all-ones vector of dimension $d$, yielding the perturbed form:
+For stabilizing tasks (VanderPol, Pendulum, DuctedFan, Two-link), we introduce process noise $\bm{\epsilon}_{t}\sim\mathcal{N}(\mathbf{0},\sigma_{\text{env}}\cdot\mathbf{1}_{d})$ to the system dynamics, where $\mathbf{1}_{d}$ denotes the all-ones vector of dimension $d$, yielding the perturbed form:
 $\mathbf{x}_{t+1}=f(\mathbf{x}_{t},\mathbf{u}_{t})+\bm{\epsilon}_{t}.$
@@
 QuadrotorTracking $0.01,0.02$
-Robustness results for stabilization tasks in Fig.
+Robustness results for stabilizing tasks in Fig.
 6 show that all trajectories converge to the equilibrium $\mathbf{x}_{g}=\mathbf{0}$ despite parametric perturbations and process noise.
@@
 8e), demonstrating superior generalization capability.
-Figure 6: System trajectories under the MSACL-trained policies in stabilization tasks subjected to environmental parametric uncertainties and process noise: (a) VanderPol with state variables $\mathbf{x}=[x,\dot{x}]^{\top}$; (b) Pendulum with angular position $\theta$ and angular velocity $\dot{\theta}$; (c) DuctedFan with coordinates $x,y$ and orientation $\theta$; (d) Two-link with joint angles $q_{1}$ and $q_{2}$.
+Figure 6: System trajectories under the MSACL-trained policies in stabilizing tasks subjected to environmental parametric uncertainties and process noise: (a) VanderPol with state variables $\mathbf{x}=[x,\dot{x}]^{\top}$; (b) Pendulum with angular position $\theta$ and angular velocity $\dot{\theta}$; (c) DuctedFan with coordinates $x,y$ and orientation $\theta$; (d) Two-link with joint angles $q_{1}$ and $q_{2}$.
 All trajectories exhibit robust convergence to the neighborhood of the equilibrium $\mathbf{x}_{g}=\mathbf{0}$.
@@
 As summarized in Table IV, performance generally improves with larger $n$, as a longer horizon enables the policy to better identify the stability direction, i.e., the descent of $V_{\psi}$.
-This reduces training bias and reinforces stability guarantees.
+This reduces training bias and reinforces stability properties.
 Empirically, $n=20$ delivers satisfactory performance across all benchmarks, serving as a versatile default.
@@
 VI-A Discussion
-This subsection elucidates the functional roles of MSACL components.
-As highlighted in Remark 4, MSACL augments the SAC framework with a stability-oriented critic, departing from standard RL that relies on exhaustive exploration to discover stabilizing behaviors.
-Instead, MSACL leverages a learned Lyapunov certificate to provide explicit, structured guidance for policy updates.
-Theoretically, MSACL is grounded in Lemma 1, which establishes the equivalence between Lyapunov stability and exponential convergence.
-To operationalize this exponential stability requirement, the ESL mechanism (Definition 2) categorizes trajectory segments based on their contraction properties, transforming the stability condition into a supervised learning task.
-This ensures the synthesis of a high-fidelity Lyapunov certificate $V_{\psi}$ that accurately captures the underlying system dynamics.
-Such fidelity is crucial for providing the actor with theoretically sound, stability-aware gradients.
-Consequently, the stability advantage in (18) prioritizes actions that facilitate Lyapunov descent, effectively guiding the policy toward the region of stability.
-While developed for stabilization and tracking, MSACL is extensible to safety-critical applications including autonomous vehicle trajectory planning [52], spacecraft proximity operations [53], and safety-guaranteed robotic manipulation [54, 55].
-The integration of Lyapunov guarantees within MERL represents a step toward verifiable learning-based control.
+This subsection elucidates the functional roles of MSACL components and discusses the fundamental limitations of our approach.
+Functionally, MSACL augments the SAC framework with a stability-oriented certificate, departing from standard RL that relies on exhaustive exploration to discover stabilizing behaviors.
+Instead, MSACL leverages a learned Lyapunov certificate to provide guidance for policy updates.
+By categorizing trajectory segments through proposed sampling mechanism, the exponential stability constraints are coupled with the training samples, resulting in an empirically accurate Lyapunov certificate.
+Consequently, the stability advantage explicitly prioritizes actions that facilitate Lyapunov descent, effectively guiding the policy toward the region of stability.
+Despite its favorable empirical performance, particularly its strong effectiviness and high efficiency, it is crucial to acknowledge that MSACL, fundamentally rooted in model-free RL, lacks formal theoretical guarantees for system stability.
+In contrast, traditional frameworks such as cMDPs [9, 42] and Hamilton-Jacobi reachability analysis [10, 43] offer rigorous mathematical formulations that provide strict theoretical guarantees for system behavior.
+Nevertheless, these formal methods often struggle with the curse of dimensionality and high computational costs in unknown environments.
+Bridging the gap between the empirical scalability of deep RL methods like MSACL and the rigorous theoretical guarantees of formal mathematical frameworks remains an essential challenge for future research.
 VI-B Conclusion
-This article presents MSACL, a novel Lyapunov-based RL algorithm that integrates exponential stability into MERL and provides a framework for developing verifiable, learning-based policies in safety-critical applications.
-By embedding Lyapunov theory within the SAC architecture, MSACL achieves stable control for model-free systems while preserving maximum entropy exploration benefits.
-Comprehensive evaluation demonstrates consistent superiority of MSACL over established baselines in both performance and stability metrics, with remarkable robustness and generalization capabilities.
-Future work will explore extensions to partially observable systems, adaptive Lyapunov architectures, and cross-domain transfer learning for complex dynamical environments.
+We in this article presented MSACL, a multi-step Lyapunov-guided reinforcement learning algorithm that provides an efficient framework for stabilizing control tasks.
+By integrating a Lyapunov-oriented critic into an off-policy MERL architecture, MSACL enables rapid and robust stabilization for model-free systems while improving sample efficiency through data reuse and broader exploration.
+Extensive experiments across multiple benchmarks demonstrated the consistent empirical superiority of MSACL over established baselines in terms of performance, robustness, and generalization capabilities.
+Despite these promising results, the learned Lyapunov certificates in MSACL remain empirical approximations and do not yet provide strict formal guarantees comparable to those of classical analytical methods.
+Therefore, an important direction for future work is to integrate MSACL with formal verification tools.
+In particular, we plan to combine the proposed multi-step RL framework with structured analytical Lyapunov certificate templates, with the goal of enhancing both theoretical rigor and practical reliability.
+By bridging the scalability and exploration efficiency of MSACL with the mathematical rigor of formal methods, we hope to develop learning-based control strategies that are not only highly efficient, but also equipped with provable stability and safety guarantees.